### v1

# High-Dimensional Sparse Clustering via Iterative Semidefinite Programming Relaxed K-Means

Abstract
We propose an iterative algorithm for clustering high-dimensional data, where the true signal lies in a much lower-dimensional space. Our method alternates between feature selection and clustering, without requiring precise estimation of sparse model parameters. Feature selection is performed by thresholding a rough estimate of the discriminative direction, while clustering is carried out via a semidefinite programming (SDP) relaxation of K-means. In the isotropic case, the algorithm is motivated by the minimax separation bound for exact recovery of cluster labels using varying sparse subsets of features. This bound highlights the critical role of variable selection in achieving exact recovery. We further extend the algorithm to settings with unknown sparse precision matrices, avoiding full model parameter estimation by computing only the minimally required quantities. Across a range of simulation settings, we find that the proposed iterative approach outperforms several state-of-the-art methods, especially in higher dimensions.
Keywords: exact recovery; Gaussian mixture models; high-dimensional data; K-means; precision matrix; semidefinite relaxation; sparsity; unsupervised learning; variable selection
1 Introduction
High-dimensional clustering algorithms group unlabeled observations when the number of features exceeds or matches the sample size. It is widely applied to data from advanced recording technologies, such as neurophysiological recordings (Cunningham and Yu,, 2014; Kadir et al.,, 2014; Park et al.,, 2024), gene expression (Bittner et al.,, 2000; Curtis et al.,, 2012; He et al.,, 2021), and neuroimaging (Eshaghi et al.,, 2021; Heller et al.,, 2006; Poulakis et al.,, 2022), which produce high-dimensional measurements but limited samples due to acquisition costs. In these settings, typically only a small subset of features is relevant for clustering. Then, a key challenge lies in the interdependence between the two unsupervised tasks of variable selection and clustering. On one hand, with unknown sparsity, including too many noise features in the clustering algorithm hinders the results. On the other hand, it is challenging to pinpoint the signal features when the true cluster structure is unknown. This motivates an iterative algorithm that alternates between clustering and variable selection until convergence leading to the identification of the true clusters.
Sparse high-dimensional clustering model.
To effectively study the role of sparsity, we assume a Gaussian mixture model with sparse discriminative directions. Let $\mathbf{X}\in\mathbb{R}^{p\times n}$ denote the data matrix, where $p$ may exceed $n$. Its columns $\mathbf{X}_{1},\ldots,\mathbf{X}_{n}$ are independently generated as $\mathbf{X}_{i}\sim\mathcal{N}(\boldsymbol{\mu}^{\ast}_{k},\boldsymbol{\Sigma}^{\ast})$ if $i\in G_{k}^{\ast}$, where the non-random sets $G_{1}^{\ast},\ldots,G_{K}^{\ast}$ form a partition of $\{1,\ldots,n\}$. The number of clusters $K$ is fixed and known. The unknown cluster centers are denoted by $\boldsymbol{\mu}^{\ast}_{1},\ldots,\boldsymbol{\mu}^{\ast}_{K}\in\mathbb{R}^{p}$, and $\boldsymbol{\Sigma}^{\ast}\in\mathbb{R}^{p\times p}$ is the common covariance matrix. Let us denote the true precision matrix as $\boldsymbol{\Omega}^{\ast}:=(\boldsymbol{\Sigma}^{\ast})^{-1}$. We consider both cases where $\boldsymbol{\Omega}^{\ast}$ is known and where it is unknown but have a sparse structure. Our objective is to recover $G_{1}^{\ast},\dots,G_{K}^{\ast}$ while treating $\boldsymbol{\mu}^{\ast}_{k}$’s and $\boldsymbol{\Sigma}^{\ast}$ as nuisance parameters. We assume that only a small, unknown subset of features, indexed by $S_{0}$, is relevant for clustering. Formally, $S_{0}:=\bigcup_{1\leq k\neq l\leq K}\mathrm{supp}\bigl{(}\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})\bigr{)}\subset\{1,\ldots,p\}$ with $|S_{0}|\ll p$. This assumption aligns with results from classification of two-class Gaussian mixtures, where the optimal decision boundary is proportional to $\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$, maximizing the signal-to-noise ratio (Fan et al.,, 2013).
Previous methods.
High-dimensional sparse clustering methods roughly fall into two types: two-step and iterative. We start with the two-step type. Since clustering after conventional dimension reduction, such as PCA (Ghosh and Chinnaiyan,, 2002; Liu et al.,, 2003; Tamayo et al.,, 2007), can miss a sparse clustering signal (Chang,, 1983), many methods adopt (Gaussian mixture) models to select features based on the clustering signal. They select features using multiple testing based on goodness-of-fit (Jin and Wang,, 2016) or moments (Azizyan et al.,, 2013; Verzelen and Arias-Castro,, 2017), model selection (Maugis et al.,, 2009; Raftery and Dean,, 2006), sparse PCA (Löffler et al.,, 2022), or an estimated sparse discriminating direction (denoted $\boldsymbol{\beta}^{\ast}$; see Section 3.1) (Azizyan et al.,, 2015). More examples include Bouveyron and Brunet-Saumard, (2014); Brodinová et al., (2019); Kadir et al., (2014); Liu et al., (2023). Since it is difficult to identify the signal features without knowing the cluster structure, two-step methods often require a stronger signal compared to iterative methods (compare the conditions in Jin and Wang, (2016) with Cai et al., (2019)). On the other hand, iterative methods repeatedly refine cluster assignments and model parameters estimates. Most methods alternate between clustering and regularized estimation of model parameters, such as feature weights (Chakraborty et al.,, 2023; Huo and Tseng,, 2017; Li et al.,, 2018; Tsai and Chiu,, 2008; Witten and Tibshirani,, 2010; Yuan et al.,, 2024), cluster centers (Sun et al.,, 2012), and feature rankings (Arias-Castro and Pu,, 2017; Zhang et al.,, 2020). Regularized EM algorithms follow the same principle: the E-step updates cluster assignments, while the M-step performs regularized estimation of sparse model parameters (Bouveyron et al.,, 2007; Dong et al.,, 2024; Fraley and Raftery,, 2007; Fu et al.,, 2021; Guo et al.,, 2010; Pan and Shen,, 2007; Wang and Zhu,, 2008; Xie et al.,, 2008; Zhou et al.,, 2009). Notably, Cai et al., (2019) establish a minimax rate of order $(s\log p)/n$ for the excess clustering risk on new samples, based on estimating $\boldsymbol{\beta}^{\ast}$ from the training data using $\ell_{1}$-regularized EM. Sparse Bayesian hierarchical models (Tadesse et al.,, 2005; Yao et al.,, 2025) and sparse convex clustering (Chakraborty and and Xu,, 2023; Wang et al.,, 2018) also follow the iterative paradigm, as their fitting procedures, such as MCMC or ADMM , progressively refine parameter estimates and feature selection.
Our contributions.
In the existing literature of iterative methods, clustering heavily depends on accuracy of estimated nuisance parameters such as cluster centers. Additionally, their estimation requires tuning sparsity-inducing regularization parameters or hyperpriors, which are challenging to tune in unsupervised problems. In contrast, our method alternates between clustering and feature selection, where our feature selection step does not require precise estimation of nuisance parameters. Features are selected by thresholding a rough estimate of the discriminating direction $\boldsymbol{\beta}^{\ast}$; this estimate is used solely to guide feature selection and has no direct impact on the clustering step. Based on the selected features, clustering is carried out using an SDP-relaxed K-means on a transformed dataset, avoiding accurate estimation of cluster centers or the covariance matrix. To obtain the transformed dataset, we adapt a high-dimensional precision matrix estimation tool (Fan and Lv,, 2016), modifying it to estimate only the minimally required quantities without estimating high dimensonal nuisance parameters. In the case of a known isotropic covariance matrix, we motivate our algorithm by deriving a minimax separation bound for exact recovery under varying feature subsets by SDP $K$-means, which scales as $\log n+s\log p/n+\sqrt{s\log p/n}$, while the minimax lower bound scales as $\log n$. In the common high-dimensional regime where $s\log p/n\leq\log n$, the upper and lower bounds match.
Notation.
We denote vectors and matrices by boldface letters, and their individual elements by the corresponding non-boldface letters. For a matrix $\mathbf{A}$, single and double subscripts denote columns and (row, column) elements, respectively. A subscript may be a single index or an index set. For vectors not derived from a matrix, a subscript denotes an element or a sub-vector. If the vector or matrix already contains a subscript, we use parentheses to avoid ambiguity, for example, $(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{S}$. We write $\mathbf{A}\geq 0$, $\mathbf{A}>0$, and $\mathbf{A}\succeq 0$ to denote elementwise nonnegativity, elementwise positivity, and positive semidefiniteness of a matrix, respectively. For a set $S$, let $|S|$ denotes its cardinality. For a positive integer $a$, let $\mathbf{1}_{a}\in\mathbb{R}^{a}$ denote the all-one vector, and $[a]$ denote $\{1,\cdots,a\}$. For a set $G_{k}\subset[n]$, let $\mathbf{1}_{G_{k}}\in\{0,1\}^{n}$ denote the indicator vector that takes the value 1 at indices in $G_{k}$ and 0 elsewhere.
2 SDP relaxation of K-means and theoretical motivation
Section 2.1 presents the SDP relaxation of $K$-means (Peng and Wei,, 2007; Zhuang et al.,, 2023) and its limitation in high dimensions due to its inability to adapt to sparsity, motivating our sparsity-aware extension. Section 2.2 provides theoretical motivation for the proposed approach, supporting iterative methods that solve multiple SDP relaxations over varying subsets of variables and highlighting the importance of sparsity.
2.1 SDP relexation of K-means problem and its limitation in high dimensions
It has been derived in (Peng and Wei,, 2007; Zhuang et al.,, 2023) that finding the partition $G_{1},\ldots,G_{K}$ that maximizes the K-means objective can be formulated as the following NP-hard mixed-integer program (Mahajan et al.,, 2012):
$\max_{\mathbf{H}\in\{0,1\}^{n\times K}}\;\langle\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X},\mathbf{H}\mathbf{B}\mathbf{H}^{\top}\rangle,\quad\text{s.t.}\quad\mathbf{H}\mathbf{1}_{K}=\mathbf{1}_{n},$ (1)
where the inner product denotes the Frobenius inner product, $\mathbf{B}:=\mathrm{diag}(|G_{1}|,\ldots,|G_{K}|)$, and each row of $\mathbf{H}$ is a one-hot cluster indicator, formally $H_{i,k}:=\mathbbm{1}(\mathbf{X}_{i}\in G_{k})$ for $i\in[n]$ and $k\in[K]$. This objective is connected to the log likelihood function in the Gaussian mixture model (see Appendix E.1 for details). By observing $\mathbf{Z}=\mathbf{H}\mathbf{B}\mathbf{H}^{\top}$ satisfies $\mathbf{Z}^{\top}=\mathbf{Z},\mathbf{Z}\succeq 0,\mathrm{tr}(\mathbf{Z})=K,\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n}$, and $\mathbf{Z}\geq 0$, we lift $\mathbf{Z}$ into the convex space and solve the following semidefinite problem:
$\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\langle\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X},\mathbf{Z}\rangle,\quad\text{s.t.}\quad\mathbf{Z}^{\top}=\mathbf{Z},~{}\mathbf{Z}\succeq 0,~{}\mathrm{tr}(\mathbf{Z})=K,~{}\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n},~{}\text{and}~{}\mathbf{Z}\geq 0.$ (2)
This relaxation, derived in (Zhuang et al.,, 2023), can be solved using off-the-shelf SDP solvers (for example, (Sun et al.,, 2020)). Moreover, faster approximation methods for solving this relaxation have been proposed (Zhuang et al.,, 2022, 2024).
When $\boldsymbol{\Sigma}^{\ast}$ is known to be $\sigma^{2}\mathbf{I}_{p}$ with fixed dimensionality $p$, Chen and Yang, (2021) proves that this SDP relaxation is a minimax optimal clustering algorithm under the exact recovery loss, with the minimax separation rate for clustering scaling as $\sigma^{2}\big{(}\log n+\sqrt{(\log n)^{2}+Kp/n}\big{)}$. This rate increases with $p$ at a square root rate. It also highlights that, the SDP relaxation may fail in high-dimensional settings where $p$ is comparable or even larger than $n$, unless the separation between clusters also scales with $p$ quickly at a square root rate – an assumption that can be very stringent for large $p$.
On the other hand, sparsity assumption has been commonly imposed in the literature to address the challenges related to high dimensionality, and has been verified to hold true in broad application scenarios such as medical imaging (Mishra et al.,, 2017), medical signal processing (Zhang et al.,, 2021), financial modeling (Nystrup et al.,, 2021), and genetics (Lee et al.,, 2015). This motivates us to investigate how to modify the SDP $K$-means algorithm to incorporate sparsity in high dimensions and explore the corresponding separation rate. Before presenting our algorithm, we first theoretically investigate the fundamental difficulty of the problem and understand the role of sparsity in high-dimensional clustering using SDP relaxation.
2.2 Minimax rate for separation bound in isotropic case
For theoretical study, this section assumes that $\boldsymbol{\Sigma}^{\ast}=\sigma^{2}\mathbf{I}_{p}$ is known. We study how the separation bound for exact recovery scales with dimension $p$, sample size $n$, and sparsity $s$. We adopt a minimax framework to assess worst-case performance over a distribution class characterized by $\ell_{2}$ separation between sparse cluster centers. Given a variable subset $S$, the clustering step of our algorithms presented in Section 3 solves a reduced version of SDP problem (2) by restricting to variables in $S$:
$\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\mathbf{Z}\rangle\quad\text{s.t.}\quad\mathbf{Z}^{\top}=\mathbf{Z},~{}\mathbf{Z}\succeq 0,~{}\mathrm{tr}(\mathbf{Z})=K,~{}\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n},~{}\text{and}~{}\mathbf{Z}\geq 0,$ (3)
The analysis proceeds in two steps: First, we show that if there exists a feature subset whose $\ell_{2}$ separation (i.e., $\Delta$ in (4) below) exceeds a threshold (explicitly determined by the sample size $n$, dimension $p$, and sparsity level $s$) then exact cluster label recovery is possible. Second, we construct a worst-case distribution where the separation is no more than a constant multiple of this threshold, under which no method can achieve exact recovery. This establishes that no clustering algorithm can succeed under weaker separation conditions. Together, these results underscore the role of sparsity in problem difficulty and serves as a motivation of our iterative algorithm that incorporates variable selection for addressing sparsity.
We begin by specifying the minimax framework. For each parameter $\boldsymbol{\theta}=(\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{K},\mathbf{H})$, where $\boldsymbol{\mu}_{k}\in\mathbb{R}^{p}$ are cluster centers and $\mathbf{H}\in\{0,1\}^{n\times K}$ denotes the cluster assignment matrix, as explained in (1), let $\mathbb{P}_{\boldsymbol{\theta}}$ be the corresponding joint distribution of $n$ independent Gaussian vectors. Each vector has mean given by the assigned cluster center and covariance $\sigma^{2}\mathbf{I}_{p}$. For all $1\leq k\neq l\leq K$, we impose two restrictions on $\boldsymbol{\mu}_{k}-\boldsymbol{\mu}_{l}$. First, it has at most $s$ nonzero entries. Second, its squared $\ell_{2}$ norm exceeds a specified threshold $\Delta^{2}$. These conditions define the following Gaussian mixture distribution class:
$\Theta(s,\Delta,K):=\mathcal{M}(s,\Delta,K)\times\mathcal{H}(K),$ (4)
$\text{where }\mathcal{M}(s,\Delta,K):=\bigl{\{}(\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{K}):$ $\boldsymbol{\mu}_{k}\in\mathbb{R}^{p},\|\boldsymbol{\mu}_{k}-\boldsymbol{\mu}_{l}\|_{2}^{2}\geq\Delta^{2},~{}\text{for all}~{}1\leq k\neq l\leq K,$
$~{}\cup_{k,l}\mathrm{supp}(\boldsymbol{\mu}_{k}-\boldsymbol{\mu}_{l})=S_{0}\subset[p],|S_{0}|\leq s\bigr{\}},$
and $\mathcal{H}(K):=\bigl{\{}\mathbf{H}=(H_{i,k})\in\{0,1\}^{n\times K}:\sum_{k=1}^{n}H_{i,k}=1~{}\text{for all }i=1,\ldots,n\bigr{\}}$. For a given clustering algorithm $\hat{\mathbf{H}}:\mathbb{R}^{n}\to\mathcal{H}(K)$, we assess its performance using the worst-case probability of exact cluster recovery over the distribution class $\Theta(s,\Delta,K)$:
$\sup_{\boldsymbol{\theta}\in\Theta(s,\Delta,K)}\min_{\pi\in S_{K}}\mathbb{P}_{\boldsymbol{\theta}}\bigl{(}\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H}\bigr{)},$ (5)
where $S_{K}$ is the set of label permutations of $[K]$. For brevity, we omit $\min_{\pi\in S_{K}}$ henceforth.
Given a problem instance $\boldsymbol{\theta}^{\ast}=(\boldsymbol{\mu}_{1}^{\ast},\ldots,\boldsymbol{\mu}_{k}^{\ast},\mathbf{H}^{\ast})\in\Theta(s,\Delta,K)$, the true cluster assignment $\mathbf{H}^{\ast}$ has one-to-one correspondence to the block-diagonal matrix $\mathbf{Z}^{\ast}:=\mathbf{H}^{\ast}\mathbf{B}^{\ast}(\mathbf{H}^{\ast})^{\top}=\sum_{k=1}^{K}|G_{k}^{\ast}|^{-1}\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top},$ up to label permutations. Let us define:
$m$ $:=2\min_{1\leq k\neq l\leq K}\left\{\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\right\},\quad\Delta_{S\cap S_{0}}^{2}:=\min_{1\leq k\neq l\leq K}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2},~{}\text{and}$
$\mathcal{S}$ :=$:=\left\{S:S\subset[p],|S|\leq\sqrt{p},\Delta_{S\cap S_{0}}^{2}\gtrsim\sigma^{2}\left(\log n+\frac{|S|\log p}{m}+\sqrt{\frac{|S|\log p}{m}}\right)\right\}.$ (6)
The following theorem presents a separation condition for exact recovery by the SDP K-means.
Theorem 1 (Separation condition).
Assume that there exists a universal constant $C_{1}$ such that $m\geq C_{1}n/\log n.$ Also assume $|G_{k}^{\ast}|\geq 2$ for all $k\in[K]$. Let $\hat{\mathbf{Z}}(S)$ be the minimizer of (3) corresponding to the subset $S\in\mathcal{S}$. Then $\mathbb{P}(\hat{\mathbf{Z}}(S)=\mathbf{Z}^{*},\forall S\in\mathcal{S})\geq 1-C_{3}K/n,$ where $C_{3}>0$ is some constant.
The proof of Theorem 1 is presented in Appendix C. The stated probability in Theorem 1 holds uniformly over all support sets in $\mathcal{S}$, ensuring the validity of iterative algorithms that solve multiple SDP problems over varying subsets $S$ across iterations. The assumption on $m$ prevents overly imbalanced clusters. It is essential for theoretical guarantees for clustering and is either explicitly stated (Azizyan et al.,, 2013; Cai et al.,, 2019; Chen and Yang,, 2021; Chen and Zhang,, 2024; Jin and Wang,, 2016; Löffler et al.,, 2021; Verzelen and Arias-Castro,, 2017) or implicitly acknowledged (Giraud and Verzelen,, 2019; Royer,, 2017) in much of the literature.
Next, we investigate the fundamental limits of sparse Gaussian mixture clustering. Specifically, we establish that for any clustering method to guarantee worst-case exact recovery over the $\ell_{2}$-separated class in (4), its required separation threshold must be at least $\sigma^{2}\log n.$
Theorem 2.
Consider the regime $(s\log p)/n=o(1)$. For $n\geq 2$, no clustering rule can guarantee exact recovery in the worst case over the class $\overline{\Theta}:=\Theta\bigl{(}s,C_{4}\sigma^{2}\log n,K\bigr{)},$ with $C_{4}$ some small positive constant. In other words, we have $\inf_{\hat{\mathbf{H}}}\sup_{\boldsymbol{\theta}\in\overline{\Theta}}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H})\geq C_{5},$ with $C_{5}$ a constant.
The proof of Theorem2 is presented in Appendix D. Theorem 2 implies that, under the regime of $|S|\log p/m\lesssim\log n$, the separation condition in Theorem 1 (i.e., lower bound for $\Delta^{2}_{S\cap S_{0}}$) is dominated by $\log n$ and thus matches the bound in Theorem 2 up to a constant.
The above theoretical results also indicate that, in the sparse regime where $|S_{0}|\ll p$, variable selection can be critical for clustering accuracy. To understand that, consider the simple scenario where $K=2$, $|G_{1}^{\ast}|=|G_{2}^{\ast}|$ and $(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})_{S_{0}}=\mu_{0}\mathbf{1}_{|S_{0}|}$. Then Theorems 1 and 2 prove that for any set $S$ satisfying $|S|\lesssim\sqrt{p}$, and $|S|\lesssim(n\log n)/\log p$ , exact recovery is possible if and only if
$|S\cap S_{0}|\mu_{0}^{2}\gtrsim\log n+\frac{|S|\log p}{n}+\sqrt{\frac{|S|\log p}{n}}\asymp\log n.$
This condition becomes very stringent if $S$ is a very poor estimate of $S_{0}$ (i.e., missing too many signal variables). It also suggests that a slight overestimate of $S_{0}$ may not be very damaging. These motivate us to design an algorithm that incorporates some conservative variable selection method into clustering.
3 Proposed iterative algorithms
This section presents our proposed algorithms under two scenarios. Section 3.1 considers the case where the covariance matrix is known, which includes the setting in Section 2.2 as a special case with $\boldsymbol{\Sigma}^{\ast}=\sigma^{2}\mathbf{I}_{p}$. Section 3.2 then extends the methodology to scenarios with unknown covariance. For simplicity, the algorithms are presented for $K=2$, though they naturally extend to any $K$.
3.1 Iterative algorithm under known covariance using the innovated transformation
The success of feature selection critically depends on the signal strength. In the related but different problem of classification, it has been explored in the literature Fan et al., (2013, 2015) that the innovated data transformation of $\tilde{\mathbf{X}}:=\boldsymbol{\Omega}^{\ast}\mathbf{X}$ boosts the signal strength for important features the most for a general precision matrix that may not be identity. See Fan et al., (2013) for detailed arguments and optimality study for the innovated transformation in Gaussian classification.
Our proposed procedure works with the transformed data $\tilde{\mathbf{X}}$ (or its estimate when $\boldsymbol{\Omega}^{\ast}$ unknown, in Section 3.2). It alternates between a variable selection step and a clustering step, detailed below, with the complete procedure summarized in Algorithm 2.
Variable selection step.
As argued in Cai et al., (2019), clustering is closely related to classification where the optimal rule is the Bayes rule with linear classification boundary $\boldsymbol{x}^{T}\boldsymbol{\beta}^{\ast}$ for any given test point $\boldsymbol{x}\in\mathbb{R}^{p}$. Here, $\boldsymbol{\beta}^{\ast}:=\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$. It is seen that the magnitudes of entries in $\boldsymbol{\beta}^{\ast}$ reflect the feature importance. Motivated by this, we select variables by thresholding a rough estimate of the sparse discriminating direction $\boldsymbol{\beta}^{\ast}$, used solely to guide variable selection without aiming for precise support recovery. Note that if the true clusters $G_{1}^{*}$ and $G_{2}^{*}$ were known, $\boldsymbol{\beta}^{\ast}$ can be accurately estimated by the sample mean difference $\bar{\tilde{\mathbf{X}}}_{G_{1}^{*}}-\bar{\tilde{\mathbf{X}}}_{G_{2}^{*}}$, where $\bar{\tilde{\mathbf{X}}}_{G_{k}^{*}}\in\mathbb{R}^{p}$ stands for the sample mean vector formed by columns of $\tilde{\mathbf{X}}$ in cluster $k=1,2$.
However, in the clustering problems $G_{1}^{*}$ and $G_{2}^{*}$ are unknown. We propose to iteratively construct an estimate of $\boldsymbol{\beta}^{\ast}$ based on the previous step clustering result. Specifically, at iteration $t$, let the estimated cluster assignments from the previous step be $\hat{G}_{1}^{t-1}$ and $\hat{G}_{2}^{t-1}$. We estimate $\boldsymbol{\beta}^{\ast}$ by the difference of empirical cluster centers: $\hat{\boldsymbol{\beta}}^{t}=\bar{\tilde{\mathbf{X}}}_{\hat{G}^{t-1}_{1}}-\bar{\tilde{\mathbf{X}}}_{\hat{G}^{t-1}_{2}},$ and form the set of selected features in the current step as
$\hat{S}^{t}:=\biggl{\{}j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\omega_{jj}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\biggr{\}}.$ (7)
We provide some intuition on the above proposed variable selection method. Note that if $j\not\in S_{0}$, then $\tilde{X}_{ji}\overset{i.i.d.}{\sim}\mathcal{N}(\tilde{\mu}_{j},\omega_{jj})$, where $\omega_{jj}$ is the $j$th diagonal entry of $\boldsymbol{\Omega}^{\ast}$ and $\tilde{\mu}_{j}$ is the mean value that is invariant to class label. For any data partition $G_{1}\cup G_{2}=[n]$ and $j\not\in S_{0}$, the $j$th entry of the mean difference vector $\bar{\tilde{\mathbf{X}}}_{G_{1}}-\bar{\tilde{\mathbf{X}}}_{G_{2}}$ follows the distribution $\mathcal{N}\bigl{(}0,(w_{jj}n)/(|G_{1}|\cdot|G_{2}|)\bigr{)}$. Then we can derive that for any $\delta>0$,
$\mathbb{P}\biggl{(}\max_{G_{1},G_{2}}\max_{j\in S_{0}^{c}}\left|\sqrt{\frac{|G_{1}|\cdot|G_{2}|}{\omega_{jj}}}(\bar{\tilde{\mathbf{X}}}_{G_{1}}-\bar{\tilde{\mathbf{X}}}_{G_{2}})_{j}\right|\leq\sqrt{2n\log\bigl{(}2(p-|S_{0}|)/\delta\bigr{)}}\biggr{)}\geq 1-\delta,$ (8)
by the Gaussian maximal inequality (see Appendix F.1 for details). We choose a slightly more conservative threshold in (7), following the intuition discussed in the last section that some mild over selection does not hurt clustering.
Clustering step.
At iteration $t$, the clustering step solves the reduced version of SDP-relaxed K-means problem (2), formulated with the selected variables $\hat{S}^{t}$ and using the identity $\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}=\tilde{\mathbf{X}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{X}}$. Since this step is reused in unknown covariance scenario, we present it as Algorithm 1. In line 1 of Algorithm 1, $\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}^{\top}\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}$ is different from $(\tilde{\mathbf{X}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{X}})_{\hat{S}^{t},\hat{S}^{t}}$, and captures the key sparse clustering signal in $\boldsymbol{\beta}^{\ast}$ (see Appendix E.2 for details).
Algorithm 1 SDP K-means subroutine for $K=2$ : SDPkmeans
1:Transformed and truncated data matrix $\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}$, covariance sub-matrix $\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}$
2:$\hat{\mathbf{Z}}^{t}\leftarrow\arg\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\langle\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}^{\top}\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot},\mathbf{Z}\rangle\newline \hskip 27.0pt\text{s.t.}\quad\mathbf{Z}^{\top}=~{}\mathbf{Z},~{}\mathbf{Z}\succeq~{}0,~{}\mathrm{tr}(\mathbf{Z})=~{}2,\mathbf{Z}\boldsymbol{1}_{n}=~{}\boldsymbol{1}_{n},~{}\mathbf{Z}\geq 0$
3:$\hat{G}^{t}_{1},\hat{G}^{t}_{2}\leftarrow$ spectral clustering result (for example, Han et al.,, 2023) on matrix $\hat{\mathbf{Z}}^{t}$ as an affinity matrix
4:Output $\hat{G}^{t}_{1},\hat{G}^{t}_{2}$
Algorithm 2 Iterative SDP K-means with known covariance
1:Data matrix $\mathbf{X}$, population covariance matrix $\boldsymbol{\Sigma}$, maximum number of iteration $T$
2:$\hat{G}_{1}^{0},\hat{G}_{2}^{0}\leftarrow$ cluster estimates obtained by running initial clustering (i.e., spectral clustering) on $\mathbf{X}$
3:$\tilde{\mathbf{X}}\leftarrow\boldsymbol{\Sigma}^{-1}\mathbf{X}$triangleright Transform
4:$(\omega_{11},\ldots,\omega_{pp})\leftarrow\operatorname{diag}(\boldsymbol{\Sigma}^{-1})$
5:for  t = 1, 2, …$T$ do
6:     $\hat{\boldsymbol{\beta}}^{t}\leftarrow\bigl{(}|\hat{G}^{t-1}_{1}|^{-1}\sum_{i\in\hat{G}^{t-1}_{1}}\tilde{\mathbf{X}}_{i}\bigr{)}-\bigl{(}|\hat{G}^{t-1}_{2}|^{-1}\sum_{i\in\hat{G}^{t-1}_{2}}\tilde{\mathbf{X}}_{i}\bigr{)}$
7:     $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\omega_{jj}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\right\}$triangleright Variable selection
8:     $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{SDPkmeans}(\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot},\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}})$triangleright Clustering (Algorithm 1)
9:Output $\hat{G}^{T}_{1},\hat{G}^{T}_{2}$
3.2 Iterative algorithm under unknown covariance
We now extend Algorithm 2 to the unknown covariance scenario by adapting the estimation idea proposed in Fan and Lv, (2016). The major advantage of our estimation method is that it does not require estimating the full precision matrix $\boldsymbol{\Omega}^{*}$. The main idea is that, at the $t$th iteration, we form an estimate of $\tilde{\mathbf{X}}$ directly based on the current cluster assignments without estimating $\boldsymbol{\Omega}^{*}$, and then calculate the estimate $\hat{S}^{t}$ in (7) using the estimated $\tilde{\mathbf{X}}$.
The direct estimation of $\tilde{\mathbf{X}}$ is done via the Innovated Scalable Efficient Estimation (ISEE) procedure (Fan and Lv,, 2016), originally developed for high-dimensional sparse precision matrix estimation. We modify it to selectively estimate only the required quantities for our algorithm, avoiding full precision matrix recovery, and use its theory to set the threshold for variable selection. Section 3.2.1 summarizes the ISEE procedure adapted as a subroutine in our method, followed by the full algorithm in Section 3.2.2.
3.2.1 Innovated Scalable Efficient Estimation (ISEE) subroutine
The ISEE procedure is motivated by two key observations. First, the transformed vector $\tilde{\mathbf{X}}_{i}=\boldsymbol{\Omega}^{\ast}\mathbf{X}_{i}\in\mathbb{R}^{p}$ remains Gaussian, with $\operatorname{Cov}(\tilde{\mathbf{X}}_{i})=\boldsymbol{\Omega}^{\ast}\boldsymbol{\Sigma}^{\ast}\boldsymbol{\Omega}^{\ast}=\boldsymbol{\Omega}^{\ast}.$ Without loss of generality, assume $p$ is an even integer. Consider a subset $A\subset\{1,\ldots,p\}$ of size 2. Assume $i\in G_{1}^{\ast}$, and consider the following decomposition of $\tilde{\mathbf{X}}_{A,i}$ as the sum of deterministic mean part and mean-zero random noise vector:
$\tilde{\mathbf{X}}_{A,i}=(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A}+\tilde{\mathcal{E}}_{A,i},$ (9)
Using the the conditional distribution property of multivariate Gaussian and Shur complement, we derive the following regression relationship between sub-vectors of $\mathbf{X}_{i}\sim\mathcal{N}(\boldsymbol{\mu}_{1}^{\ast},\boldsymbol{\Sigma}^{\ast})$:
[…]
where $\mathbf{E}_{A,i}\sim\mathcal{N}(0,(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1})$. Leveraging this relationship, the quantities in (9) can be further expressed as follows (Fan and Lv,, 2016):
$(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A}=\boldsymbol{\Omega}^{\ast}_{A,A}(\boldsymbol{\alpha}_{1})_{A}\quad\text{and}\quad\tilde{\mathcal{E}}_{A,i}=\boldsymbol{\Omega}^{\ast}_{A,A}\mathbf{E}_{A,i}.$ (11)
Appendix F.2 provides detailed derivations for (10) and (11). Since $p-|A|$ can be comparable or larger than $n$, one may want to leverage the existing regularization method for fitting the regression model (10) to obtain sample estimates of $\boldsymbol{\Omega}^{\ast}_{A,A}$, $\mathbf{E}_{A,i}$ and $(\boldsymbol{\alpha}_{1})_{A}$, which further yields plug-in estimate of $\tilde{\mathbf{X}}_{A,i}$. This process can be repeated for all $A\in\mathcal{A}$ with $\mathcal{A}$ a partition set of $\{1,\cdots,p\}$ consisting of sets of size 2. By combining estimates for $\tilde{\mathbf{X}}_{A,i}$ for all $A\in\mathcal{A}$, we can form an estimate of the full matrix $\tilde{\mathbf{X}}_{G_{1}^{*}}\in\mathbb{R}^{|G_{1}^{*}|\times p}$.
A similar expression holds for $i\in G_{2}^{\ast}$, with $\boldsymbol{\mu}_{1}^{\ast}$ and $\boldsymbol{\alpha}_{1}$ replaced by $\boldsymbol{\mu}_{2}^{\ast}$ and $\boldsymbol{\alpha}_{2}$, respectively. Following the same steps as described above, we can obtain the estimate of the submatrix $\tilde{\mathbf{X}}_{G_{2}^{*}}\in\mathbb{R}^{|G_{2}^{*}|\times p}$.
Two issues remain to be addressed. First, the true cluster labels are unknown. Second, we need an implementable ISEE subroutine. Since our algorithm is iterative by nature, we use the cluster assignments in the current iteration as the proxy for true cluster labels to implement the ISEE subroutine, and then iteratively update the estimates for $\tilde{\mathbf{X}}$ and cluster assignments. We next explain the ISEE subroutine for a given pair of cluster assignment $(G_{1},G_{2})$.
ISEE subroutine.
[…]
3.2.2 Full algorithm with iterative variable selection via ISEE
Given the ISEE subroutine, we extend Algorithm 2 to the unknown covariance scenario. In the variable selection step, we threshold $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t}$, instead of $\bar{\tilde{\mathbf{X}}}_{\hat{G}_{1}^{t-1}}-\bar{\tilde{\mathbf{X}}}_{\hat{G}_{2}^{t-1}}$. In the clustering step, we replace $\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}$ with the estimated transformed data matrix $\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot}$. We replace $\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}$ with the pooled sample covariance, combining the sample covariances of $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{1}}$ and $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{2}}$. We leverage the convergence rate of the ISEE procedure (Fan and Lv,, 2016) to guide the choice of threshold for $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t}$. Specifically, for each $k\in\{1,2\}$ and for each block $A_{\ell}$ with $\ell\in\{1,\ldots,L\}$, under some regularity conditions (Fan and Lv,, 2016), we have:
$\bigl{\|}\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-$ $(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\bigr{)}-\bigl{(}({\tilde{\boldsymbol{\mu}}}_{1}^{\ast})_{A_{\ell}}-({\tilde{\boldsymbol{\mu}}}_{2}^{\ast})_{A_{\ell}}\bigr{)}\bigr{\|}_{2}\lesssim C_{6}\max\{J\log p/n,\sqrt{\log p/n}\},$ (12)
where $J$ is the maximum number of nonzero off-diagonal entries per row of $\boldsymbol{\Omega}^{\ast}$, and $C_{6}=\|\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}\|_{2}$. This bound comes from the convergence rates of $\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}$ and $\|(\hat{\boldsymbol{\alpha}}_{k})_{A_{\ell}}-(\boldsymbol{\alpha}_{k})_{A_{\ell}}\|_{2}$ proved in Fan and Lv, (2016). The detailed derivation is provided in Appendix F.2. In the standard sparse high-dimensional regime of $J\sqrt{\log p/n}=o(1)$, the term $\sqrt{\log p/n}$ dominates the bound. Since ISEE only involves small submatrices of size $2\times 2$ or $3\times 3$, the terms in $C_{6}$ are bounded and contribute only constant factors. To account for these constants and suppress false discoveries, we include a slowly diverging factor $\sqrt{\log n}$ in our threshold. Combining all the pieces, we set the variable selection threshold as $\sqrt{\log n\log p/n}$. The final algorithm incorporating these thresholding principles is presented in Algorithm 3.
Algorithm 3 Iterative SDP K-means with unknown covariance
1:Data matrix $\mathbf{X}$, maximum number of iteration $T$
2:$\hat{G}_{1}^{0},\hat{G}_{2}^{0}\leftarrow$ cluster estimates obtained by running spectral clustering on $\mathbf{X}$triangleright Initialization
3:for  t = 1, 2, …$T$ do
4:     $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t},\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t},\hat{\tilde{\mathbf{X}}}^{t},\leftarrow\texttt{ISEE}(\mathbf{X},\hat{G}_{1}^{t-1},\hat{G}_{2}^{t-1})$$\triangleright$ Estimated transformation (Algorithm 4)
5:     $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|(\widehat{\tilde{\boldsymbol{\mu}}}_{1}^{t}-\widehat{\tilde{\boldsymbol{\mu}}}_{2}^{t})_{j}|>\sqrt{\log n\log p/n}\right\}$triangleright Variable selection
6:     $\hat{\boldsymbol{\Sigma}}_{\hat{S}^{t},\hat{S}^{t}}\leftarrow$ pooled covariance of covariances computed from $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{1}}$ and $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{2}}$
7:     $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{SDPkmeans}(\hat{\tilde{\mathbf{X}}}_{\hat{S}^{t},\cdot},\hat{\boldsymbol{\Sigma}}_{\hat{S}^{t},\hat{S}^{t}})$triangleright Clustering (Algorithm 1)
8:Output $\hat{G}^{T}_{1},\hat{G}^{T}_{2}$
4 Numerical results
For all simulations, we consider a symmetric bicluster Gaussian mixture model with equal cluster sizes and a common covariance. For data dimension $p$ and sample size $n$, the data are generated as
$\mathbf{X}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\boldsymbol{\mu}^{\ast},\boldsymbol{\Sigma}^{\ast}),~{}i=1,\dots,n/2,~{}\text{and}~{}\mathbf{X}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(-\boldsymbol{\mu}^{\ast},\boldsymbol{\Sigma}^{\ast}),i=n/2+1,\dots,n.$
We define $\boldsymbol{\beta}^{\ast}:=\boldsymbol{\Omega}^{\ast}\boldsymbol{\mu}^{\ast}$ here and set its support to $S_{0}=[10]$. All non-zero entries of $\boldsymbol{\beta}^{\ast}$ have the same magnitude, chosen to keep the signal strength, measured by the Mahalanobis distance $\Delta_{\boldsymbol{\Sigma}^{\ast}}=2\sqrt{(\boldsymbol{\mu}^{\ast})^{\top}\boldsymbol{\Omega}^{\ast}\boldsymbol{\mu}^{\ast}}=\sqrt{(\boldsymbol{\beta}^{\ast})^{\top}\boldsymbol{\Sigma}^{\ast}\boldsymbol{\beta}^{\ast}},$ as constant, while increasing the noise level through higher ambient dimension $p$. We evaluate performance using the mis-clustering rate, averaged over 200 independent replications. For initialization, we use spectral clustering (Han et al.,, 2023). We set the number of iteration as $T=100$, but used early stopping criterion. The details on this criterion, code implementation, and details required for replicating our results are provided in Appendix G.1.
Table 1: Summary of Simulation Settings. $i,j=1,\ldots,p$.
Methods Evaluated Covariance structure $p$ $n$ Separation
Algorithm 2 (for known $\boldsymbol{\Sigma}^{\ast}$) known, $\boldsymbol{{\Sigma}}^{\ast}=\mathbf{I}_{p}$ 50-5000 200 $\Delta^{2}=4^{2},5^{2}$
Algorithm 3 (for unknown $\boldsymbol{\Sigma}^{\ast}$) unknown, $\Omega^{*}_{i,i+1}=\Omega^{*}_{i+1,i}=\rho$, $\Omega^{*}_{i,i}=1$, $\Omega^{*}_{i,j}=0$ o.w., $\rho\in\{0.2,0.45\}$ 50-400 500 $\Delta_{\boldsymbol{\Sigma}^{\ast}}^{2}=3^{2},4^{2}$
Scenario 1: known isotropic covariance
We evaluate Algorithm 2 under the setting of line 1 of Table 1. We first compare non-sparse clustering methods, including spectral clustering (Han et al.,, 2023), hierarchical clustering, and SDP-relaxed K-means with all features (Chen and Yang,, 2021), with Algorithm 2 initialized by each of them. Here, initialization refers to a rough cluster assignment $\hat{G}_{1}^{0},\hat{G}_{2}^{0}$ used to guide the first variable selection step. Results in Figures 1-(a) and 1-(b) show that while the non-sparse methods fail to adapt to sparsity, Algorithm 2, when properly initialized, is robust to high-dimensional noise and effectively adapts to sparsity. Even moderately good initializations are sufficient, as evidenced by the improved performance of Algorithm 2 when initialized with spectral or SDP K-means clustering. In contrast, poor initializations can hinder recovery, as reflected in the weaker performance of Algorithm 2 initialized by hierarchical clustering. Next, we benchmark our method against existing sparse clustering approaches, including a two-step method (IF-PCA; Jin and Wang,, 2016) and iterative methods such as SAS (Arias-Castro and Pu,, 2017), Sparse K-means (Witten and Tibshirani,, 2010), and CHIME (Cai et al.,, 2019). The results in Figures 1-(c) and 1-(d) highlight the competitiveness of our approach. Details of the baseline methods are provided in Appendix G.2.
Figure 1: Mis-clustering rates of Algorithm 2 compared to sparse and non-sparse baselines. The dotted lines (non-sparse methods) can also be interpreted as the mis-clustering rates of Algorithm 2 immediately after initialization, before any iteration. The settings are provided in line 1 of Table 1.
Scenario 2: Unknown covariance.
We evaluate the performance of Algorithm 3 in the setting with a chain-graph sparse precision matrix with conditional correlation $\rho\in\{0.2,0.4\}$, following the setup described in line 2 of Table 1. Figures 2-(a) to 2-(d) compare Algorithm 3 with two alternative variants for the variable selection step: Algorithm 5 estimates the full precision matrix using graphical lasso (Friedman et al.,, 2008), while Algorithm 6 applies thresholding to the empirical cluster means of the transformed data, using a cutoff derived from the Gaussian maximal inequality. Details of these algorithms are provided in Appendix G.2. The results show that Algorithm 3 outperforms both alternatives, highlighting the benefits of bypassing full precision matrix estimation and employing convergence-rate-guided thresholding. We also compare our method against other sparsity-aware clustering approaches. The result presented in Figures 2-(e) to 2-(h) show that our method outperforms the others.
Figure 2: Mis-clustering rates of Algorithm 3, compared to its variants (Algorithms 5 and 6; plots a–d) and other sparsity-aware baseline methods (plots e-h), under unknown covariance scenario with varying conditional correlations. The settings are provided in line 2 of Table 1.
5 Discussion
We investigated the performance of the SDP relaxation of the K-means problem in the context of high-dimensional sparse clustering. Assuming a known isotropic covariance matrix, we derived the minimax separation bound for exact recovery of cluster labels with sparse subsets of features using SDP K-means. Our results demonstrate the critical role of variable selection in achieving exact recovery of cluster labels. Building on this insight, we proposed an iterative algorithm that integrates variable selection into the SDP relaxation framework. Our method avoids the precise estimation of nuisance parameters such as cluster centers and the covariance matrix (which may not be the identity matrix). There are two potential directions for future research: i) establish theoretical guarantees that our algorithm can identify a subset with sufficient separation, and ii) extend our minimax framework to the general case of an unknown covariance matrix.
Appendix A Overview of Appendix
This supplementary material provides the technical proofs deferred in the main text, along with some additional results of interest. The content is organized as follows:
  • Appendix B contains the technical lemmas and definitions used in the proof of Theorem 1.
  • Appendix C provides the proof of the separation bound for exact recovery by semidefinite relaxation of K-means, as presented in Theorem 1.
  • Appendix D proves the minimax lower bound for exact recovery, as presented in Theorem 2.
  • Appendix E provides additional details on the SDP objectives deferred from Sections 2.1 and 3.1, including their derivation and how they capture the sparse clustering signal.
  • Appendix F provides an omitted details for the iterative algorithms, which were deferred from Sections 3.1 and 3.2.1.
  • Appendix G provides additional information on the numerical studies presented in Section 4.
Appendix B Technical lemmas and definitions
This section presents technical lemmas and definitions that are used throughout the main proofs, beginning the tail bounds for the chi-square distribution.
Lemma 3 (Lemma 1 of Laurent and Massart, 2000).
If $Z$ has a chi-square distribution with $p$ degrees of freedom, then for all $t>0$, we have
$\mathbb{P}(Z\geq p+2\sqrt{pt}+2t)\leq e^{-t},\quad\text{and}\quad\mathbb{P}(Z\leq p-2\sqrt{pt})\leq e^{-t}.$
Before presenting the next two lemmas, we formally define the sub-exponential norm.
Definition 4.
Let $\psi_{1}(x):=e^{x}-1$. The sub-exponential norm of a random variable $X$, denoted by $\|X\|_{\psi_{1}}$, is defined as
$\|X\|_{\psi_{1}}:=\inf\left\{\lambda>0:\mathbb{E}\left[\psi_{1}\left(\frac{|X|}{\lambda}\right)\right]\leq 1\right\}.$
Using this definition, Lemma 5 provides an upper bound on the subexponential norm of a chi-square random variable.
Lemma 5 (Lemma VIII.2 of Chen and Yang, 2021).
Let $\mathbf{X}_{1},\mathbf{X}_{2}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,I_{q})$. Then there exists a constant $C>0$ such that
$\|\|\mathbf{X}_{1}\|_{2}^{2}-q\|_{\psi_{1}}+\|\langle\mathbf{X}_{1},\mathbf{X}_{2}\rangle\|_{\psi_{1}}\leq C\sqrt{q}.$
Next, Lemma 6 provides a maximal inequality for sub-exponential norm.
Lemma 6 (Lemma 2.2.2 of Van Der Vaart and Wellner, 2023).
For random variables $X_{1},\ldots,X_{m}$, there exists a constant $C_{\psi_{1}}$ such that
$\left\|\max_{1\leq i\leq m}X_{i}\right\|_{\psi_{1}}\leq C_{\psi_{1}}\,\log(m+1)\max_{1\leq i\leq m}\|X_{i}\|_{\psi_{1}}.$
Next, Lemma 7 provides an upper bound on the $\ell_{2}$ operator norm of an $m\times n$ random matrix with i.i.d. Gaussian entries.
Lemma 7 (Corollary 5.35 of Vershynin, 2012).
Let $\mathbf{A}\in\mathbb{R}^{m\times n}$ have i.i.d. $\mathcal{N}(0,\sigma^{2})$ entries. Then, for any $t>0$, we have
$\mathbb{P}\bigl{(}\|\mathbf{A}\|_{\mathrm{op}}\geq\sigma(\sqrt{m}+\sqrt{n}+\sqrt{2t})\bigr{)}\leq e^{-t}.$
Next, Lemma 8 provides a concentration inequality for empirical processes indexed by an unbounded function class.
Lemma 8 (Theorem 4 of Adamczak, 2008).
Let $X_{1},\dots,X_{n}$ be independent random variables taking values in a measurable space $(S,\mathcal{B})$, and let $\mathcal{F}$ be a countable class of measurable functions $f:S\to\mathbb{R}$. Assume that for all $f\in\mathcal{F}$ and all $i=1,\ldots,n$, it holds that $\mathbb{E}[f(X_{i})]=0$ and
$\bigl{\|}\sup_{f\in\mathcal{F}}|f(X_{i})|\bigr{\|}_{\psi_{1}}<\infty,$
Define
$Z:=\sup_{f\in\mathcal{F}}\bigl{|}\sum_{i=1}^{n}f(X_{i})\bigr{|},~{}\gamma^{2}:=\sup_{f\in\mathcal{F}}\sum_{i=1}^{n}\mathrm{Var}[f(X_{i})],~{}\text{and}~{}M_{\psi_{1}}:=\|\max_{i\in[n]}\sup_{f\in\mathcal{F}}|f(X_{i})|\|_{\psi_{1}}$
Then there exists a constant $C>0$ such that for all $t>0$, we have
$\mathbb{P}\left(Z\geq 2\mathbb{E}[Z]+t\right)\leq\exp\left(-\frac{t^{2}}{3\gamma^{2}}\right)+3\exp\left(-\frac{t}{CM_{\psi_{1}}}\right).$
Next, Lemma 9 presents a concentration inequality for the suprema of the Gaussian process.
Lemma 9 (Theorem 5.8 of Boucheron et al., 2016).
Let $(X_{t})_{t\in\mathcal{T}}$ be an almost surely continuous, centered Gaussian process indexed by a totally bounded set $\mathcal{T}$. Define
$\gamma^{2}:=\sup_{t\in\mathcal{T}}\operatorname{Var}(X_{t}^{2}),\quad\text{and}\quad Z:=\sup_{t\in\mathcal{T}}X_{t}.$
Then, we have $\operatorname{Var}(Z)\leq\gamma^{2}$. Moreover, for any $t>0$,
$\mathbb{P}(Z-\mathbb{E}[Z]\geq t)\leq\exp\left(-\frac{t^{2}}{2\gamma^{2}}\right).$
The next lemma, Dudley’s integral inequality, bounds the expected supremum of a Gaussian process.
Lemma 10 (Theorem 8.1.3 of Vershynin, 2018).
Let $(T,d)$ be a metric space, and let $N(T,d,\varepsilon)$ denote the $\varepsilon$-covering number of $T$ with respect to $d$. Let $(X_{t})_{t\in T}$ be a mean-zero random process on a metric space with sub-Gaussian increments. Then
[…]
where $C$ is an absolute constant, and $K_{\psi_{2}}$ is the sub-Gaussian parameter.
Finally, Lemma 11 presents the uniform Hanson-Wright inequality for Gaussian quadratic forms.
Lemma 11 (Lemma VIII.4 of Chen and Yang, 2021).
Let $\mathbf{X}\sim\mathcal{N}(0,I_{q})$ and $\mathcal{A}$ be a bounded class of $q\times q$ matrices. Define the random variable
$Z=\sup_{\mathbf{A}\in\mathcal{A}}\biggl{(}\mathbf{X}^{\top}\mathbf{A}\mathbf{X}-\mathbb{E}[\mathbf{X}^{\top}\mathbf{A}\mathbf{X}]\biggr{)},$
which corresponds to the supremum of a centered random process. Then, there exists a constant $C>0$ such that for any $t>0$,
$\mathbb{P}(|Z-\mathbb{E}[Z]|\geq t)\leq 2\exp\left[-C\min\left(\frac{t^{2}}{\|\mathbf{X}\|_{\mathcal{A}}^{2}},\frac{t}{\sup_{\mathbf{A}\in\mathcal{A}}\|\mathbf{A}\|_{\mathrm{op}}}\right)\right],$
where $\|\mathbf{X}\|_{\mathcal{A}}:=\mathbb{E}\left[\sup_{\mathbf{A}\in\mathcal{A}}\|(\mathbf{A}+\mathbf{A}^{T})\mathbf{X}\|_{2}\right]$.
Appendix C Proof of Theorem 1
The proof follows the strategy of Theorem II.1 of Chen and Yang, (2021), with modifications to account for the sparsity assumption. The proof proceeds in three key steps.
  1. 1.
    Reformulation into high-probability bound problem (Appendix C.1): For each $S$, we construct three statistics that bound a dual certificate $\lambda^{S}$, which verifies exact cluster recovery by the SDP problem (3) corresponding to $S$. A nontrivial gap between these bounds guarantees the existence of such a certificate (Corollary 18).
  2. 2.
    First bounds and construction of the dual variable (Appendix C.2). Under the separation and set size conditions imposed by $\mathcal{S}$, we derive uniform high-probability bounds on the two statistics defined in the first step and determine the value of $\lambda^{S}$ accordingly (Lemma 21).
  3. 3.
    Second bounds and conclusion (Appendix C.3). We establish a uniform high-probability bound on the third statistic and verify that the previously chosen value of $\lambda^{S}$ exceeds this bound, thereby confirming the existence of a dual certificate and completing the proof.
[…]
Definition 12 (Matrix-based notations).
We view the data as a matrix $\mathbf{X}\in\mathbb{R}^{p\times n}$ and write $\mathbf{X}=\mathbf{M}^{\ast}+\mathcal{E}$, where $\mathbf{M}^{\ast}$ contains the cluster means column-wise, and $\mathcal{E}$ collects the noise vectors. For any support set $S\subset[p]$, we denote the $i$th observation and its noise restricted to $S$ by
$\mathbf{X}_{S,i}\quad\text{and}\quad\mathcal{E}_{S,i},$
respectively. Given true clusters $G_{1}^{\ast},\ldots,G_{K}^{\ast}$, we define the empirical mean and average noise of cluster $k$ on support $S$ as
$\overline{\mathbf{X}}^{\ast}_{S,k}:=\frac{1}{|G_{k}^{\ast}|}\sum_{i\in G_{k}^{\ast}}\mathbf{X}_{S,i},\quad\overline{\mathcal{E}}^{\ast}_{S,k}:=\frac{1}{|G_{k}^{\ast}|}\sum_{i\in G_{k}^{\ast}}\mathcal{E}_{S,i}.$
Before we begin, we summarize the new notations used in the proof of Theorem 1 in Table 2 for the reader’s reference.
Table 2: Notations used in the proof of Theorem 1.
Symbol Description Type
$\Gamma_{K}$ Orthogonal complement of cluster span Subspace of $\mathbb{R}^{n}$
$\dot{\lambda}^{S}$ Constructed dual variable for trace constraint Scalar
The following statistics depend on the value of $\lambda^{S}$ and assymetric with respect to their indices:
$\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})\in\mathbb{R}^{n}$ Constructed dual variable for symmetry constraint Random vector
$\dot{\mathbf{B}}^{S}(\lambda^{S})\in\mathbb{R}^{n\times n}$ Constructed dual variable for nonnegativity Random matrix
$\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\in\mathbb{R}^{|G_{k}^{\ast}|\times|G_{l}^{\ast}|}$ $(k,l)$th off-diagonal block of $\dot{\mathbf{B}}^{S}$ Random matrix
$\dot{\mathbf{W}}^{S}(\lambda^{S})\in\mathbb{R}^{n\times n}$ A function of $\lambda^{S}$, $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$, and $\dot{\mathbf{B}}^{S}(\lambda^{S})$ Random matrix
$\mathbf{r}^{(k,l,S)}\in\mathbb{R}^{|G_{k}^{\ast}|}$ Row-sum of $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}$;   $\mathbf{r}^{(k,l,S)}\neq\mathbf{r}^{(l,k,S)}$. Random vector
$t^{S}_{lk}(\lambda^{S})$ Sum of $\mathbf{r}^{(l,k,S)}$ (not $\mathbf{r}^{(k,l,S)}$); Total sum of $\dot{\mathbf{B}}^{S}_{G_{l}^{\ast},G_{k}^{\ast}}$ Random variable
$U^{S}(\mathbf{X})$ Upper bound on $\lambda^{S}$ Random variable
$L_{1}^{S}(\mathbf{X})$ Lower bound on $\lambda^{S}$; independent of $\lambda^{S}$ Random variable
$L_{2}^{S}(\mathbf{X},\lambda^{S})$ Lower bound on $\lambda^{S}$; depends on $\lambda^{S}$ Random variable
The following statistics are random variables and asymmetric with respect to their indices:
$D^{S}_{kli}(\mathbf{X})$ Cluster-separation statistic that defines $U^{S}(\mathbf{X})$;
$I^{S}_{kl}(\mathbf{v},\mathbf{X})$ Signal–noise interaction term that defines $T^{S}_{1,kl}$ and further $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$N^{S}_{kl}(\mathbf{v},\mathbf{X})$ Noise–noise interaction term that defines $T^{S}_{2,kl}$ and further $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$T^{S}_{1,kl}(\mathbf{v})$ $I_{kl}(\mathbf{v},\mathbf{X})\cdot I_{lk}(\mathbf{v},\mathbf{X})$; Defines $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$T^{S}_{2,kl}(\mathbf{v})$ $N_{kl}(\mathbf{v},\mathbf{X})\cdot N_{lk}(\mathbf{v},\mathbf{X})$; Defines $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$T^{S}_{3,kl}(\mathbf{v})$ Mixture of $T^{S}_{1,kl}(\mathbf{v})$ and $T^{S}_{2,kl}(\mathbf{v})$; Defines $L_{2}^{S}(\mathbf{X},\lambda^{S})$
$G^{S}_{1,kl}(\mathbf{v})$ Sum of noise inner products that comprises $N^{S}_{kl}(\mathbf{v},\mathbf{X})$
$G^{S}_{2,kl}(\mathbf{v})$ Sum of mean noise inner products that comprises $N^{S}_{kl}(\mathbf{v},\mathbf{X})$
C.1 Reformulation into high-probability bound problem
The primary goal of this section is to derive Corollary 18, which converts the optimality condtion into high-probability bounds. Let $S\subset[p]$ be a support set, and consider the primal SDP problem (13) corresponding to $S$, recalled below:
$\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\mathbf{Z}\rangle\quad\text{s.t.}\quad\mathbf{Z}^{\top}=\mathbf{Z},~{}\mathbf{Z}\succeq 0,~{}\mathrm{tr}(\mathbf{Z})=K,~{}\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n},~{}\mathbf{Z}\geq 0.$ (13)
Let $\mathcal{L}(\mathbf{Z},\mathbf{Q}^{S},\lambda^{S},\boldsymbol{\alpha}^{S},\mathbf{B}^{S})$ denote the associated Lagrangian function. Here, the Lagrange multipliers
$\mathbf{Q}^{S}\in\mathbb{R}^{n\times n},\quad\lambda^{S}\in\mathbb{R},\quad\boldsymbol{\alpha}^{S}\in\mathbb{R}^{n},\quad\text{and}\quad\mathbf{B}^{S}\in\mathbb{R}^{n\times n}$
correspond to the primal constraints
$\mathbf{Z}\succeq 0,\quad\mathrm{tr}(\mathbf{Z})=K,\quad\mathbf{Z}^{\top}=\mathbf{Z},\quad\text{and}\quad\mathbf{Z}\geq 0,$
respectively. Note that $\mathbf{Q}^{S}\succeq 0$ and $\mathbf{B}^{S}\geq 0$. Based on this formulation, this section proceeds in three main steps to ultimately establish Corollary 18:
  1. 1.
    Necessary conditions for dual certificates (Appendix C.1.1): We formalize sufficient conditions for the existence of dual certificates (Lemma 13) and leverage them to derive necessary conditions (Lemma 14) that guide the construction of dual variables.
  2. 2.
    Reduction to a single dual variable construction (Appendix C.1.2): Leveraging the necessary conditions, we parameterize the dual variables in terms of $\lambda^{S}$ , reducing the construction to a one-dimensional problem (Lemma 15).
  3. 3.
    Reformulation into high-probability bound problem (Appendix C.1.3): We derive statistics that bound the dual certificate $\lambda^{S}$ from above (Lemma 16) and below (Lemma 17). This shows that a nontrivial gap between them guarantees the existence of dual variables satisfying the conditions of Lemma 13 (Corollary 18).
C.1.1 Necessary conditions for dual certificates
Let $S\subset[p]$ be a support set, and consider the corresponding primal SDP problem (13). Lemma 13 provides sufficient conditions under which the dual variables $\lambda^{S}$, $\boldsymbol{\alpha}^{S}$, and $\mathbf{B}^{S}$ form a valid dual certificate (with $\mathbf{Q}^{S}$ implicitly captured by the conditions). In particular, the lemma guarantees that the duality gap vanishes at the true cluster membership matrix
$\mathbf{Z}^{\ast}=\sum_{k=1}^{K}\frac{1}{|G_{k}^{\ast}|}\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top}.$
Lemma 13 (Optimality conditions).
Let $S\subset[p]$ be a support set, and consider the corresponding primal SDP problem (13). The matrix $\mathbf{Z}^{\ast}$ is the unique solution to this problem if there exist dual variables $\lambda^{S}\in\mathbb{R}$, $\boldsymbol{\alpha}^{S}\in\mathbb{R}^{n}$, and $\mathbf{B}^{S}\in\mathbb{R}^{n\times n}$ such that the following conditions are satisfied:
[…]
In Lemma 13, conditions (C1) and (C2) ensure dual feasibility, (C3) and (C4) guarantee a zero duality gap, and (C5) ensures the uniqueness of the solution. The proof is provided in Appendix C.4.1. Building on these conditions, we now establish necessary conditions for dual certificates. Leveraging this notation, we state the necessary conditions in the following lemma.
Lemma 14 (Necessary conditions for optimality).
Let $S\subset[p]$ be a support set, and consider the corresponding primal SDP problem (13). Suppose dual variables $\lambda^{S}\in\mathbb{R}$, $\boldsymbol{\alpha}^{S}\in\mathbb{R}^{n}$, and $\mathbf{B}^{S}\in\mathbb{R}^{n\times n}$ satisfy conditions (C2), (C3), and (C4) of Lemma 13. Then, for all $1\leq k\neq l\leq K$ and $i\in G_{k}^{\ast}$, the following statements hold:
$\boldsymbol{\alpha}^{S}_{i}$ $=2\mathbf{X}_{S,i}^{\top}\overline{\mathbf{X}}^{\ast}_{S,k}-\frac{\lambda^{S}}{|G_{k}^{\ast}|}-\|\overline{\mathbf{X}}^{\ast}_{S,k}\|_{2}^{2},~{}\text{and}~{}$ (14)
$\bigl{(}\mathbf{B}^{S}_{G_{k}^{\ast}G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}\bigr{)}_{i}$ $=-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{k}^{\ast}|}\lambda^{S}+\frac{|G_{l}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)},~{}$ (15)
where $\bigl{(}\mathbf{B}^{S}_{G_{k}^{\ast}G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}\bigr{)}_{i}$ denotes the $i$th entry of the row-wise sum of $\mathbf{B}^{S}_{G_{k}^{\ast}G_{l}^{\ast}}\in\mathbb{R}^{|G_{k}^{\ast}|\times|G_{l}^{\ast}|}$, which is the $(k,l)$th off-diagonal block of $\mathbf{B}^{S}$.
The proof of Lemma 14 is provided in Appendix C.4.2.
C.1.2 Reduction to a single dual variable construction
For any given $\lambda^{S}\in\mathbb{R}$, condition (14) uniquely determines the corresponding value of $\boldsymbol{\alpha}^{S}$, which we denote by $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$. We then turn to the dual variable $\mathbf{B}^{S}$ and construct $\dot{\mathbf{B}}^{S}(\lambda^{S})$, designed to satisfy condition (15). The matrix $\dot{\mathbf{B}}^{S}(\lambda^{S})$ is structured as a block matrix, with symmetric rank-one matrices as its off-diagonal blocks and zero matrices as its diagonal blocks. These constructions follow the proof of Theorem II.1 in Chen and Yang, (2021), with modifications to account for varying support sets.
We now describe the construction of $\dot{\mathbf{B}}^{S}(\lambda^{S})$. Fix any value of $\lambda^{S}\in\mathbb{R}$, and for notational simplicity, we suppress the dependence on $\lambda^{S}$ in what follows. For each $k=1,\ldots,K$, define the diagonal block $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}\in\mathbb{R}^{|G_{k}^{\ast}|\times|G_{k}^{\ast}|}$ to be the zero matrix. Next, consider the off-diagonal blocks. Given $\lambda^{S}$, the data, and the true cluster structure, condition (15) uniquely determines the row-wise sum of the $(k,l)$ off-diagonal block of the dual certificate. Denote this unique vector by $\mathbf{r}^{(k,l,S)}\in\mathbb{R}^{|G_{k}^{\ast}|}$. Then, for each $1\leq k\neq l\leq K$, we define the off-diagonal block $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\in\mathbb{R}^{|G_{k}^{\ast}|\times|G_{l}^{\ast}|}$ as the symmetric rank-one matrix:
$\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}:=\frac{1}{\mathbf{1}_{|G_{l}^{\ast}|}^{\top}\mathbf{r}^{(l,k,S)}}\mathbf{r}^{(k,l,S)}(\mathbf{r}^{(l,k,S)})^{\top},$ (16)
This construction ensures that the row-sum condition in (15) is satisfied:
$\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}$ $=\frac{1}{\mathbf{1}_{|G_{l}^{\ast}|}^{\top}\mathbf{r}^{(l,k,S)}}\bigl{\{}\mathbf{r}^{(k,l,S)}(\mathbf{r}^{(l,k,S)})^{\top}\bigr{\}}\mathbf{1}_{|G_{l}^{\ast}|}$
$=\frac{1}{\mathbf{1}_{|G_{l}^{\ast}|}^{\top}\mathbf{r}^{(l,k,S)}}\mathbf{r}^{(k,l,S)}\bigr{\{}(\mathbf{r}^{(l,k,S)})^{\top}\mathbf{1}_{|G_{l}^{\ast}|}\bigr{\}}$
$=\mathbf{r}^{(k,l,S)}.$
We summarize the construction of $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$ and $\dot{\mathbf{B}}^{S}(\lambda^{S})$ in the following lemma.
Lemma 15 (Parametrized dual certificate candidates).
Fix the data matrix $\mathbf{X}$ and the true cluster structure $G_{1}^{\ast},\ldots,G_{K}^{\ast}$. Let $S\subset[p]$ be a support set, and consider the corresponding primal SDP problem (13). Fix a value of the dual variable $\lambda^{S}\in\mathbb{R}$. Define $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})\in\mathbb{R}^{n}$ by setting, for each $k=1,\ldots,K$,
$\dot{\boldsymbol{\alpha}}^{S}_{G_{k}^{\ast}}(\lambda^{S}):=2\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}^{\ast}_{S,k}-\frac{\lambda^{S}}{|G_{k}^{\ast}|}\mathbf{1}_{|G_{k}^{\ast}|}-\|\overline{\mathbf{X}}^{\ast}_{S,k}\|_{2}^{2}\mathbf{1}_{|G_{k}^{\ast}|},$
Next, define $\dot{\mathbf{B}}^{S}(\lambda^{S})\in\mathbb{R}^{n\times n}$ as the block matrix:
$\dot{\mathbf{B}}^{S}(\lambda^{S}):=\begin{bmatrix}\mathbf{0}&\dot{\mathbf{B}}^{S}_{G_{1}^{\ast},G_{2}^{\ast}}(\lambda^{S})&\cdots&\dot{\mathbf{B}}^{S}_{G_{1}^{\ast},G_{K}^{\ast}}(\lambda^{S})\\ \dot{\mathbf{B}}^{S}_{G_{2}^{\ast},G_{1}^{\ast}}(\lambda^{S})&\mathbf{0}&\cdots&\dot{\mathbf{B}}^{S}_{G_{2}^{\ast},G_{K}^{\ast}}(\lambda^{S})\\ \vdots&\vdots&\ddots&\vdots\\ \dot{\mathbf{B}}^{S}_{G_{K}^{\ast},G_{1}^{\ast}}(\lambda^{S})&\dot{\mathbf{B}}^{S}_{G_{K}^{\ast},G_{2}^{\ast}}(\lambda^{S})&\cdots&\mathbf{0}\end{bmatrix},$
where for each $1\leq k\neq l\leq K$, the off-diagonal block $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}(\lambda^{S})\in\mathbb{R}^{|G_{k}^{\ast}|\times|G_{l}^{\ast}|}$ is defined as
$\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}(\lambda^{S}):=\frac{1}{\mathbf{1}_{|G_{l}^{\ast}|}^{\top}\mathbf{r}^{(l,k,S)}}\,\mathbf{r}^{(k,l,S)}\,(\mathbf{r}^{(l,k,S)})^{\top},$ (17)
with $\mathbf{r}^{(k,l,S)}\in\mathbb{R}^{|G_{k}^{\ast}|}$ defined componentwise by
$r_{i}^{(k,l,S)}:=-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{k}^{\ast}|}\lambda^{S}+\frac{|G_{l}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)}.$ (18)
Finally, we construct
$\dot{\mathbf{W}}^{S}(\lambda^{S}):=\lambda^{S}\mathbf{I}_{n}-\dot{\mathbf{B}}^{S}(\lambda^{S})+\frac{1}{2}\mathbf{1}_{n}\bigl{(}\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})\bigr{)}^{\top}+\frac{1}{2}\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})\mathbf{1}_{n}^{\top}-\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}.$
Constructed $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$ and $\dot{\mathbf{B}}^{S}(\lambda^{S})$ satisfy the conditions in Lemma 14.
C.1.3 Reformulation into high-probability bound problem
Next, we identify conditions of $\lambda^{S}$, that ensures $\dot{\mathbf{B}}^{S}(\lambda^{S})$ and $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$ satisfy the conditions in Lemma 13. We begin with conditions (C1), (C3), (C4) and (C5). To this end, we introduce the following data-dependent quantities:
$D^{S}_{kli}(\mathbf{X}):=\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2},$
and
$U^{S}(\mathbf{X}):=\min_{1\leq k\neq l\leq K}\biggl{\{}\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\min_{i\in G_{k}^{\ast}}D^{S}_{kli}(\mathbf{X})\biggr{\}}.$ (19)
Based on these definitions, we now state an upper bound on $\lambda^{S}$:
Lemma 16 (Upper bound on $\lambda^{S}$).
Let $S\subset[p]$ be a support set, and consider the corresponding primal SDP problem (13). For any value of $\lambda^{S}$, corresponding $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$ , $\dot{\mathbf{B}}^{S}(\lambda^{S})$, and $\dot{\mathbf{W}}^{S}(\lambda^{S})$ satisfy conditions (C3) and (C4) of Lemma 13. Moreover, if
$\lambda^{S}<U^{S}(\mathbf{X}),$
then $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$ and $\dot{\mathbf{B}}^{S}(\lambda^{S})$ also satisfy conditions (C1) and (C5) of Lemma 13.
Proof of Lemma 16 is provided in Appendix C.4.3.
Next we move onto condition (C2) of Lemma 13. To this end, we define some notation. We first define a linear subspace of $\mathbb{R}^{n}$ as:
$\Gamma_{K}:=\operatorname{span}\left(\mathbf{1}_{G_{k}^{*}}:k\in[K]\right)^{\perp},$
where $\perp$ means orthogonal complement. For each $\mathbf{v}\in\Gamma_{K}$ and support set $S\subset[p]$, we define a collection of data-dependent random variables. We begin with the following supremum of noise quadratic form process:
$L_{1}^{S}(\mathbf{X}):=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2},$ (20)
Next, we define two fundamental components representing signal–noise and noise–noise interactions:
$I^{S}_{kl}(\mathbf{v},\mathbf{X})$ $:=|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\;\mathcal{E}_{S\cap S_{0},i}\rangle,~{}\text{and}$ (21)
$N^{S}_{kl}(\mathbf{v},\mathbf{X})$ $:=|G_{l}^{\ast}|\sum_{i\in G_{k}^{*}}v_{i}\bigl{\langle}\overline{\mathcal{E}}_{S,k}^{\ast}-\overline{\mathcal{E}}_{S,l}^{\ast},\;\mathcal{E}_{S,i}\bigr{\rangle}.$ (22)
Note that both definitions are asymmetric with respect to the first two indices. Using these definitions, we introduce three composite random quantities:
$T^{S}_{1,kl}(\mathbf{v})~{}:=~{}$ $I^{S}_{kl}(\mathbf{v},\mathbf{X})\cdot I^{S}_{lk}(\mathbf{v},\mathbf{X}),$ (23)
$T^{S}_{2,kl}(\mathbf{v})~{}:=~{}$ $N^{S}_{kl}(\mathbf{v},\mathbf{X})\cdot N^{S}_{lk}(\mathbf{v},\mathbf{X}),~{}\text{and}$ (24)
$T^{S}_{3,kl}(\mathbf{v})~{}:=~{}$ $N^{S}_{kl}(\mathbf{v},\mathbf{X})\cdot I^{S}_{lk}(\mathbf{v},\mathbf{X})+I^{S}_{kl}(\mathbf{v},\mathbf{X})\cdot N^{S}_{lk}(\mathbf{v},\mathbf{X}).$ (25)
Next, the following quantity represents the sum of all elements of $\dot{\mathbf{B}}^{S}_{G_{l}^{\ast},G_{k}^{\ast}}$, $(l,k)$th off-diagonal block of $\dot{\mathbf{B}}^{S}(\lambda^{S})$:
$t_{lk}^{S}(\lambda^{S}):=\sum_{j\in G_{l}^{\ast}}r_{j}^{(l,k,S)}=\sum_{j\in G_{l}^{\ast}}\biggl{\{}\frac{|G_{k}^{\ast}|}{2}D_{lkj}(\mathbf{X},S)-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\lambda^{S}\biggr{\}}.$ (26)
Leveraging all these notations, we finally define:
$L_{2}^{S}(\mathbf{X},\lambda^{S}):=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\biggl{\{}\sum_{1\leq k\neq l\leq K}\frac{1}{t^{S}_{lk}(\lambda^{S})}\biggl{(}T^{S}_{1,kl}(\mathbf{v})+T^{S}_{2,kl}(\mathbf{v})+T^{S}_{3,kl}(\mathbf{v})\biggr{)}\biggr{\}},$ (27)
Using this quantity, Lemma 17 states the lower bound condition on $\lambda_{S}$ that guarantees condition (C2) of Lemma 13 is satisfied.
Lemma 17 (Lower bound on $\lambda_{S}$).
Let $S\subset[p]$ be a support set, and consider the corresponding primal SDP problem (13). If
$\lambda^{S}>L_{1}^{S}(\mathbf{X})+L_{2}^{S}(\mathbf{X},\lambda^{S}),$
then the constructed dual variables $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$ and $\dot{\mathbf{B}}^{S}(\lambda^{S})$ satisfy condition (C2) of Lemma 13.
Proof of Lemma 17 is provided in Appendix C.4.4. Combining Lemmas 16 and 17, we finally arrive at the following corollary:
Corollary 18 (Problem reformulation).
Let $S\subset[p]$ be a support set, and consider the corresponding primal SDP problem (13). If there exists $\lambda^{S}$ such that
$L_{1}^{S}(\mathbf{X})+L_{2}^{S}(\mathbf{X},\lambda^{S})<\lambda^{S}<U^{S}(\mathbf{X}),$
then the dual variables constructed with this $\lambda^{S}$ via Lemma 15 yield a valid certificate that satisfies all conditions of Lemma 13, verifying that $\mathbf{Z}^{\ast}$ is the unique optimal point of the primal problem.
C.2 First Bounds and Construction of the Dual Variable
This section is devoted to the derivation of Lemma 21, which establishes that, under the separation condition, there exists a $\lambda^{S}$ that partially satisfies the condition stated in Corollary 18. We proceed in two main steps:
  1. 1.
    High-probability bounds (Appendix C.2.1) : We derive a lower bound for $U^{S}(\mathbf{X})$ (Lemma 19) and an upper bound for $L_{1}^{S}(\mathbf{X})$ (Lemma 20), under the separation and set size condition of $\mathcal{S}$.
  2. 2.
    Specifying the value of the dual variable (Appendix C.2.2): We set the value of $\lambda^{S}$ to lie between the bounds established in the previous step (Lemma 21).
C.2.1 High-probability Bounds
Recall from (19) that $U^{S}(\mathbf{X})$ is defined as
$U^{S}(\mathbf{X})=\min_{1\leq k\neq l\leq K}\biggl{\{}\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\min_{i\in G_{k}^{\ast}}D^{S}_{kli}(\mathbf{X})\biggr{\}},$
where
$D^{S}_{kli}(\mathbf{X})=\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}.$
Recall from (6) that we define $\mathcal{S}$ as
$\mathcal{S}:=\left\{S:S\subset[p],S\cap S_{0}\neq\emptyset,|S|\leq\sqrt{p},\Delta_{S\cap S_{0}}^{2}\gtrsim\sigma^{2}\left(\log n+\frac{|S|\log p}{m}+\sqrt{\frac{|S|\log p}{m}}\right)\right\},$
where
$\Delta_{S\cap S_{0}}^{2}=\min_{1\leq k\neq l\leq K}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2},$
and
$m=\min_{1\leq k\neq l\leq K}\biggl{(}\frac{2|G_{l}^{\ast}||G_{k}^{\ast}|}{|G_{l}^{\ast}|+|G_{k}^{\ast}|}\biggr{)}\geq C_{1}\frac{n}{\log n}.$
The following lemma establishes a high-probability lower bound for $D^{S}_{kli}(\mathbf{X})$.
Lemma 19.
Uniformly across all index $(k,l,i)$ ranging over all pairs $1\leq k\neq l\leq K$ and $i\in G^{*}_{k}$, and all $S\in\mathcal{S}$, with probability at least $1-C_{7}/n$, it holds that
$D^{S}_{kli}(\mathbf{X})\gtrsim\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}|S|-R^{S}_{kl}.$
where $C_{7}>0$ is a constant and
$R^{S}_{kl}:=\sqrt{\frac{\xi^{S}}{|G_{l}^{\ast}|}}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}+\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}\sqrt{|S|\xi^{S}}+\frac{1}{|G_{k}^{\ast}|}\zeta^{S},$
and $\zeta^{S}:=\log(n^{2}|S|^{2}{\binom{p}{|S|}})$.
The proof of Lemma 19 is presented in Appendix C.4.5.
Next, Lemma 20 establishes a high-probability upper bound for $L_{1}^{S}(\mathbf{X})$, where we recall from (20):
$L_{1}^{S}(\mathbf{X})=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2},$
and $\mathcal{E}_{S,\cdot}$ is a submatrix of $\mathcal{E}\in\mathbb{R}^{p\times n}$, whose entries are i.i.d. Gaussian random variables.
Lemma 20.
Uniformly accross all $S\subset[p]$ such that $|S|\leq\sqrt{p}$, with probability at least $1-C_{8}/n$, where $C_{8}>0$ is a constant, it holds that
$L_{1}^{S}(\mathbf{X})\lesssim\sigma^{2}(n+\log n+|S|\log p).$
The proof of Lemma 20 is provided in Appendix C.4.6.
C.2.2 Specifying the Value of the Dual Variable
Since $L_{2}^{S}(\mathbf{X},\lambda^{S})$ in (27) depends on $\lambda^{S}$, we first specify a value which lies between $L_{1}^{S}(\mathbf{X})$ and $U^{S}(\mathbf{X})$ with high probability, uniformly over $S\in\mathcal{S}$.
Lemma 21.
Let $\dot{\lambda}^{S}:=\sigma^{2}|S|+m\Delta_{S\cap S_{0}}^{2}/4$. Then Uniformly over $S\in\mathcal{S}$, with probability at least $1-C_{9}/n$, where $C_{9}>0$ is a constant, it holds that
$L_{1}^{S}(\mathbf{X})\lesssim\dot{\lambda}^{S}\lesssim U^{S}(\mathbf{X}).$
Proof of Lemma 21 is provided in Appendix C.4.7.
C.3 Second high-probability bounds and conclusion
This section completes the proof by showing that, with high probability, $L_{2}^{S}(\mathbf{X},\dot{\lambda}^{S})\lesssim\dot{\lambda}^{S}$. As defined in equation (27), $L_{2}^{S}(\mathbf{X},\dot{\lambda}^{S})$ is the supremum of an empirical process involving statistics $t^{S}_{lk}$ (26), $I^{S}_{kl}(\mathbf{v},\mathbf{X})$ $I^{S}_{lk}(\mathbf{v},\mathbf{X})$ (21), $N^{S}_{kl}(\mathbf{v},\mathbf{X})$, and $N^{S}_{lk}(\mathbf{v},\mathbf{X})$ (22). Therefore, the bound on $L_{2}^{S}(\mathbf{X},\dot{\lambda}^{S})$ follows directly from bounds on the $t^{S}_{lk}$, $I^{S}_{kl}(\mathbf{v},\mathbf{X})$ and $N^{S}_{kl}(\mathbf{v},\mathbf{X})$. Accordingly, the remainder of this section proceeds as follows:
  1. 1.
    Bounding the off-diagonal block sum of the dual variable matrix (Appendix C.3.1): We derive a high-probabilty lower bound for $t^{S}_{lk}$ (Lemma 22).
  2. 2.
    Bounding the signal–noise interaction term (Appendix C.3.2): We derive a high-probability upper bound for $I^{S}_{kl}(\mathbf{v},\mathbf{X})$ (Lemma 23) and show that the corresponding component of $L_{2}^{S}(\mathbf{X},\dot{\lambda}^{S})$ is dominated by $\dot{\lambda}^{S}$ (Lemma 24).
  3. 3.
    Bounding the noise–noise interaction term (Appendix C.3.3): We derive a high-probability upper bound for $N^{S}_{kl}(\mathbf{v},\mathbf{X})$ (Lemmas 25 and 26), and show that the corresponding component of $L_{2}^{S}(\mathbf{X},\dot{\lambda}^{S})$ is dominated by $\dot{\lambda}^{S}$ (Lemma 27).
  4. 4.
    Conclusion (Appendix C.3.4).
C.3.1 Bounding the off-diagonal block sum of the dual variable matrix
With the value of $\dot{\lambda}^{S}$ set in Lemma 21, we provide a high-probability lower bound of $t^{S}_{lk}(\dot{\lambda}^{S})$ defined in (26).
Lemma 22.
For any $1\leq l\neq k\leq K$, if $|G_{l}^{\ast}|\geq 2$ and $|G_{k}^{\ast}|\geq 2$, uniformly over $S\in\mathcal{S}$, with probability at least $1-C_{7}/n$, where $C_{0}$ is the constant used in Lemma 19, it holds that
$t^{S}_{lk}(\dot{\lambda}^{S})\gtrsim|G_{l}^{\ast}|\;|G_{k}^{\ast}|\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}.$
The proof of Lemma 22 is provided in Appendix C.4.8.
C.3.2 Bounding the signal-noise interaction term
We first provide a high-probability upper bound of
$I^{S}_{kl}(\mathbf{v},\mathbf{X})=|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\;\mathcal{E}_{S\cap S_{0},i}\rangle,$
defined in (21).
Lemma 23.
Uniformly across $S\in\mathcal{S}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{10}/n$, where $C_{10}>0$ is a constant, it holds that
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}I^{S}_{kl}(\mathbf{v},\mathbf{X})$
$\quad\quad\lesssim|G_{l}^{\ast}|\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\biggl{(}|G_{k}^{\ast}|+\sqrt{|G_{k}^{\ast}|\log n}+\sqrt{|G_{k}^{\ast}||S|\log p}+\log n+|S|\log p\biggr{)}^{1/2}.$
The proof of Lemma 23 is provided in Appendix C.4.9. Building on this result, we establish the first partial upper bound for $L_{2}^{S}(\mathbf{X},\dot{\lambda}^{S})$, corresponding to the term $T^{S}_{1,kl}(\mathbf{v})=I^{S}_{kl}(\mathbf{v},\mathbf{X})\cdot I^{S}_{lk}(\mathbf{v},\mathbf{X})$, defined in (23).
Lemma 24.
Uniformly across $S\in\mathcal{S}$, with probability at least $1-C_{11}/n$, where $C_{11}>0$ is a constant, it holds that
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\sum_{1\leq k\neq l\leq K}\frac{T^{S}_{1,kl}(\mathbf{v})}{t^{S}_{lk}(\dot{\lambda}^{S})}\lesssim\dot{\lambda}^{S}.$
Proof of Lemma 24 is provided in Appendix C.4.10.
C.3.3 Bounding the noise-noise interaction term
Since $N^{S}_{kl}(\mathbf{v},\mathbf{X})$, defined in (22), is an inner product of dependent Gaussian variables, we decompose it into a sum of inner products between independent Gaussian variables.
$N^{S}_{kl}(\mathbf{v},\mathbf{X})$ $=|G_{l}^{\ast}|\sum_{i\in G_{k}^{*}}v_{i}\bigl{\langle}\overline{\mathcal{E}}_{S,k}^{\ast}-\overline{\mathcal{E}}_{S,l}^{\ast},\;\mathcal{E}_{S,i}\bigr{\rangle}$
$=|G_{l}^{\ast}|\sum_{i\in G_{k}^{*}}v_{i}\bigl{\langle}\overline{\mathcal{E}}_{S,k}^{\ast},\;\mathcal{E}_{S,i}\bigr{\rangle}-|G_{l}^{\ast}|\sum_{i\in G_{k}^{*}}v_{i}\bigl{\langle}\overline{\mathcal{E}}_{S,l}^{\ast},\;\mathcal{E}_{S,i}\bigr{\rangle}$
$=\underbrace{\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}v_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle}_{:=G^{S}_{1,kl}(\mathbf{v})}\underbrace{-|G_{l}^{\ast}|\sum_{i\in G_{k}^{*}}v_{i}\bigl{\langle}\overline{\mathcal{E}}_{S,l}^{\ast},\;\mathcal{E}_{S,i}\bigr{\rangle},}_{:=G^{S}_{2,kl}(\mathbf{v})}$
We now present high-probability upper bounds for $G^{S}_{1,kl}(\mathbf{v})$ and $G^{S}_{2,kl}(\mathbf{v})$ in order. Since all terms consist purely of noise random variables, no separation condition is needed in this case.
Lemma 25.
Uniformly across $S\in[p]$ such that $|S|\leq\sqrt{p}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{11}/n$, where $C_{11}>0$ is a constant, it holds that
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}G^{S}_{1,kl}(\mathbf{v})\lesssim|G_{l}^{\ast}|\sqrt{|S|(\log n+|S|\log p)}.$
The proof of Lemma 25 is provided in Appendix C.4.11.
Lemma 26.
Uniformly across $S\in[p]$ such that $|S|\leq\sqrt{p}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{12}/n$, where $C_{12}>0$ is a constant, it holds that
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}|G_{2,kl}^{S}(\mathbf{v})|\lesssim\sqrt{|G_{l}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}\sqrt{|G_{k}^{\ast}|+\sqrt{\xi^{S}}},$
where $\xi^{S}=\log(n|S|^{2}{\binom{p}{|S|}})$.
The proof of Lemma 26 is provided in Appendix C.4.12.
Lemma 27.
Uniformly across $S\in\mathcal{S}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{13}/n$, where $C_{13}>0$ is a constant, it holds that
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\sum_{1\leq k\neq l\leq K}\frac{T^{S}_{2,kl}(\mathbf{v})}{t^{S}_{lk}(\dot{\lambda}^{S})}\lesssim\dot{\lambda}^{S}.$
Proof of Lemma 27 is provided in Appendix C.4.13.
C.3.4 Conclusion
The inequality regarding $T^{S}_{3,kl}(\mathbf{v})$ is a direct consequence of Lemmas 24, and 27, by the Cauchy-Schwarz inequality. Therefore, by combining Lemmas 21, 24, and 27, we conclude that, uniformly over $S\in\mathcal{S}$, with probability at least $1-C_{14}/n$, where $C_{14}>0$ is a constant, the quantity $\dot{\lambda}^{S}$ constructed in Lemma 20 satisfies
$L_{1}^{S}(\mathbf{X})+L_{2}^{S}(\mathbf{X},\dot{\lambda}^{S})\lesssim\dot{\lambda}^{S}\lesssim U^{S}(\mathbf{X}).$
Therefore, by Corollary 18, the dual variables constructed in Lemma 15 serve as dual certificates verifying that $\mathbf{Z}^{\ast}$ is the unique optimal solution to the primal SDP problem corresponding to $S$, stated in (13). This completes the proof of Theorem 1.
C.4 Proof of Supporting Lemmas
This section presents omitted proof of lemmas presented in the proof of Theorem 1.
C.4.1 Proof of Lemma 13
The proof proceeds in three steps:
  1. 1.
    Conditions (C1) and (C2) from dual feasibility,
  2. 2.
    Conditions (C3) and (C4) from zero dual gap,
  3. 3.
    Condition (C5) from the block-diagonal structure of $\mathbf{Z}^{\ast}=\sum_{k=1}^{K}|G_{k}^{\ast}|^{-1}\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top}$.
Proof.
We proceed by following the steps outlined above in sequence.
1. Conditions (C1) and (C2) from dual feasibility.
After simplification, the Lagrangian function becomes an affine function of $\mathbf{Z}$:
$\mathcal{L}(\mathbf{Z},\mathbf{Q}^{S},\lambda^{S},\boldsymbol{\alpha}^{S},\mathbf{B}^{S})$
$\quad\quad:=\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\;\mathbf{Z}\rangle+\langle\mathbf{Q}^{S},\mathbf{Z}\rangle+\lambda^{S}(K-\operatorname{tr}(\mathbf{Z}))+(\boldsymbol{\alpha}^{S})^{\top}\biggl{(}\mathbf{1}_{n}-\frac{1}{2}(\mathbf{Z}+\mathbf{Z}^{\top})\mathbf{1}_{n}\biggr{)}+\langle\mathbf{B}^{S},\mathbf{Z}\rangle$
$\quad\quad=\lambda^{S}K+(\boldsymbol{\alpha}^{S})^{\top}\mathbf{1}_{n}+\bigl{\langle}\mathbf{Q}^{S}-\mathbf{W}^{S},\;\mathbf{Z}\bigr{\rangle},$
where we use $\mathrm{tr}(\mathbf{Z})=\langle\mathbf{I}_{n},\mathbf{Z}\rangle$ and recall from Lemma 13 that
$\mathbf{W}^{S}=\lambda^{S}\mathbf{I}_{n}-\mathbf{B}^{S}+\frac{1}{2}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}+\frac{1}{2}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}-\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}.$
Therefore, the Lagrange dual function, shown below, can be derived analytically, as a linear function is bounded only when the slope part is zero.
$g(\mathbf{Q}^{S},\lambda^{S},\boldsymbol{\alpha}^{S},\mathbf{B}^{S})$ $:=\sup_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\mathcal{L}(\mathbf{Z},\mathbf{Q}^{S},\lambda^{S},\boldsymbol{\alpha}^{S},\mathbf{B}^{S})$
$=\lambda^{S}K+(\boldsymbol{\alpha}^{S})^{\top}\mathbf{1}_{n}+\sup_{\mathbf{Z}\in\mathbb{R}^{n\times n}}\bigl{\langle}\mathbf{Q}^{S}-\mathbf{W}^{S},\;\mathbf{Z}\bigr{\rangle},$
As a result, we have:
$g(\mathbf{Q}^{S},\lambda^{S},\boldsymbol{\alpha}^{S},\mathbf{B}^{S})=\begin{cases}\lambda^{S}K+(\boldsymbol{\alpha}^{S})^{\top}\mathbf{1}_{n},\quad&\mathbf{Q}^{S}=\mathbf{W}^{S}~{}(\text{slope is }\mathbf{0})\\ \infty,\quad&\text{otherwise}.\end{cases}$
Making the following domain of the dual function as an explicit linear equality constraint, and combining it with $\mathbf{Q}^{S}\succeq 0$ and $\mathbf{B}^{S}\geq 0$, with some abuse of terminology, we refer to the following problem as the Lagrange dual problem associated with the primal SDP problem
$\min_{\lambda^{S}\in\mathbb{R},\boldsymbol{\alpha}^{S}\in\mathbb{R}^{n},\mathbf{B}^{S}\in\mathbb{R}^{n\times n}}\hskip-20.00003pt\lambda^{S}K+(\boldsymbol{\alpha}^{S})^{\top}\mathbf{1}_{n},\quad\text{s.t.}\quad\mathbf{B}^{S}\geq 0,\mathbf{W}^{S}\succeq 0.$
The constraints in this optimization problem correspond to conditions (C1) and (C2).
2. Conditions (C3) and (C4) from zero dual gap.
The duality gap is simply the difference between the dual and primal objective function:
$\lambda^{S}K+(\boldsymbol{\alpha}^{S})^{\top}\mathbf{1}_{n}-\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\;\mathbf{Z}\rangle$
=$\quad\quad\stackrel{{\scriptstyle(i)}}{{=}}\lambda^{S}\mathrm{tr}(\mathbf{Z})+\frac{1}{2}(\boldsymbol{\alpha}^{S})^{\top}(\mathbf{Z}+\mathbf{Z}^{\top})\boldsymbol{1}_{n}-\langle\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot},\;\mathbf{Z}\rangle$
=$\quad\quad\stackrel{{\scriptstyle(ii)}}{{=}}\bigl{\langle}\lambda^{S}\mathbf{I}_{n}+\frac{1}{2}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}+\frac{1}{2}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}-\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}-\mathbf{B}^{S},\;\mathbf{Z}\bigr{\rangle}+\langle\mathbf{B}^{S},\mathbf{Z}\rangle$
$\quad\quad=\langle\mathbf{W}^{S},\mathbf{Z}\rangle+\langle\mathbf{B}^{S},\mathbf{Z}\rangle$
where step $(i)$ uses primal constraints $\mathrm{tr}(\mathbf{Z})=K$, $\mathbf{Z}^{\top}=\mathbf{Z}$ and $\mathbf{Z}\boldsymbol{1}_{n}=\boldsymbol{1}_{n}$, step $(ii)$ uses $\mathrm{tr}(\mathbf{Z})=\langle\mathbf{I}_{n},\mathbf{Z}\rangle$, and added and subtracted $\langle\mathbf{B}^{S},\mathbf{Z}\rangle$. This duality gap is zero if $\langle\mathbf{W}^{S},\mathbf{Z}\rangle=0$ (C3) and $\langle\mathbf{B}^{S},\mathbf{Z}\rangle=0$ (C4).
3. Condition (C5) from the block-diagonal structure.
Finally, we derive condition (C5), which ensures that $\mathbf{Z}^{\ast}$ is the unique solution to the primal SDP. To this end, we first show that it is the only primal feasible point with the following block-diagonal structure: where $\mathbf{Z}_{G_{k}^{\ast},G_{l}^{\ast}}=0$ for all $1\leq k\neq l\leq K$:
$\mathbf{Z}=\begin{bmatrix}\mathbf{Z}^{(1)}&\mathbf{0}&\cdots&\mathbf{0}\\ \mathbf{0}&\mathbf{Z}^{(2)}&\cdots&\mathbf{0}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&\mathbf{Z}^{(K)}\end{bmatrix}.$
We use proof by contradiction. Suppose there exists a primal feasible $\mathbf{Z}\neq\mathbf{Z}^{\ast}$ with the block diagonal structure described above. By the primal constraint $\mathbf{Z}\mathbf{1}_{n}=\mathbf{1}_{n}$, each block $\mathbf{Z}^{(k)}$ satisfies $\mathbf{Z}^{(k)}\mathbf{1}_{|G_{k}^{\ast}|}=\mathbf{1}_{|G_{k}^{\ast}|}$, which implies that $(1,n_{k}^{-1/2}\mathbf{1}_{|G_{k}^{\ast}|})$ is an eigenvalue-eigenvector pair of $\mathbf{Z}^{(k)}$. Since the trace of a matrix equals the sum of its eigenvalues, this implies $\mathrm{tr}(\mathbf{Z}^{(k)})\geq 1$ for every $k$. Combining this with the overall trace constraint $\mathrm{tr}(\mathbf{Z})=\sum_{k=1}^{K}\mathrm{tr}(\mathbf{Z}^{(k)})=K$, we deduce that $\mathrm{tr}(\mathbf{Z}^{(k)})=1$ for every $k$. Additionally, each $\mathbf{Z}^{(k)}$ is positive semidefinite, due to the primal constraint $\mathbf{Z}\succeq 0$, along with the Schur complement condition for the positive semidefiniteness of block matrices. Then $\mathbf{Z}^{(k)}$ must take the form $n_{k}^{-1}\mathbf{1}_{|G_{k}^{\ast}|}\mathbf{1}_{|G_{k}^{\ast}|}^{\top}$, since $1$ is the only nonzero eigenvalue of $\mathbf{Z}^{(k)}$, corresponding to the eigenvector $n_{k}^{-1/2}\mathbf{1}_{|G_{k}^{\ast}|}$. Consequently, $\mathbf{Z}=\mathbf{Z}^{\ast}$, a contradiction.
Given this block diagonal structure and the condition $\mathrm{tr}(\mathbf{B}^{S}\mathbf{Z}^{\ast})=0$, we conclude that $\mathbf{Z}^{\ast}$ is the unique solution to the primal SDP problem if (C1)-(C4) is satisfied and $\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}>0$ for all $1\leq k\neq l\leq K$. This corresponds to condition (C5) and completes the proof of Lemma 13. ∎
C.4.2 Proof of Lemma 14
The proof proceeds in three steps:
  1. 1.
    Preliminary computations,
  2. 2.
    Necessary condition for $\boldsymbol{\alpha}^{S}$,
  3. 3.
    Necessary condition for $\mathbf{B}^{S}$.
Proof.
Throughout the proof, we assume that conditions (C1)–(C5) from 13 are satisfied. Recall that true cluster membership matrix is defined as $\mathbf{Z}^{\ast}=\sum_{k=1}^{K}{|G_{k}^{\ast}|}^{-1}\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top}$, where $\mathbf{1}_{G_{k}^{\ast}}$, as defined in Section 1, is the indicator vector for membership in $G_{k}^{\ast}$.
1. Preliminary computations.
We first summarize the results for reference. The derivations will follow.
$\bigl{(}(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})\mathbf{1}_{G_{l}^{\ast}}\bigr{)}_{G_{k}^{\ast}}$ $=|G_{l}^{\ast}|\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}^{\ast}_{S,l},~{}\text{for }(k,l)\in[K]^{2}~{}(\text{no condition required}),$ (28)
[…]
$(\mathbf{B}^{S}\mathbf{1}_{G_{l}^{\ast}})_{G_{k}^{\ast}}$ $=\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}~{}\text{for }1\leq k\neq l\leq K~{}(\text{no condition required}).$ (33)
We start with (28). For $1\leq k\leq K$ and $1\leq l\leq K$, using only the definition of $\mathbf{1}_{G_{k}^{\ast}}$ and $\mathbf{1}_{G_{l}^{\ast}}$ given in Section 1, we have
$\bigl{(}(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})\mathbf{1}_{G_{l}^{\ast}}\bigr{)}_{G_{k}^{\ast}}$ $=(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})_{G_{k}^{\ast},G_{k}^{\ast}}(\mathbf{1}_{G_{l}^{\ast}})_{G_{k}^{\ast}}+(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})_{G_{k}^{\ast},(G_{k}^{\ast})^{C}}(\mathbf{1}_{G_{l}^{\ast}})_{(G_{k}^{\ast})^{C}}$
=$\overset{(i)}{=}(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})_{G_{k}^{\ast},(G_{k}^{\ast})^{C}}(\mathbf{1}_{G_{l}^{\ast}})_{(G_{k}^{\ast})^{C}}$
$=(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}$
$=(\mathbf{X}_{S,G_{k}^{\ast}})^{\top}\mathbf{X}_{S,G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}$
$=|G_{l}^{\ast}|\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}^{\ast}_{S,l}.$
where step $(i)$ uses $(\mathbf{1}_{G_{l}^{\ast}})_{G_{k}^{\ast}}=0$.
Using the definition of $\mathbf{Z}^{\ast}$ and condition (C3), which states that $\langle\mathbf{W}^{S},\mathbf{Z}^{\ast}\rangle=0$, we obtain:
$\langle\mathbf{W}^{S},\mathbf{Z}^{\ast}\rangle=\sum_{k=1}^{K}\frac{1}{|G_{k}^{\ast}|}\langle\mathbf{W}^{S},\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top}\rangle=\sum_{k=1}^{K}\frac{1}{|G_{k}^{\ast}|}\mathbf{1}_{G_{k}^{\ast}}^{\top}\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}=0.$
By the condition (C2), which states that $\mathbf{W}^{S}\succeq 0$ , we have $\mathbf{1}_{G_{k}^{\ast}}^{\top}\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}\geq 0$ for all $k=1,\ldots,K$. This along with the equality above implies (29):
$\mathbf{1}_{G_{k}^{\ast}}^{\top}\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}=0~{}\text{for all}~{}k=1,\ldots,K.$
Given that $\mathbf{W}^{S}\succeq 0$, the function $\mathbf{v}\mapsto\mathbf{v}^{\top}\mathbf{W}^{S}\mathbf{v}$ is convex and bounded below by 0. Therefore, $\mathbf{1}_{G_{k}^{\ast}}^{\top}\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}=0$ implies that $\mathbf{1}_{G_{k}^{\ast}}$ minimizes $\mathbf{v}^{\top}\mathbf{W}^{S}\mathbf{v}$. By the first order condition, we obtain (30)
$\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}=\mathbf{0}~{}\text{for all}~{}k=1,\ldots,K.$
By condition (C4), which states that $\langle\mathbf{B}^{S},\mathbf{Z}^{\ast}\rangle=0$, and the block-diagonal structure of $\mathbf{Z}^{\ast}$, we obtain (31):
$\mathbf{B}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}=\mathbf{0},$
which implies (32):
$(\mathbf{B}^{S}\mathbf{1}_{G_{k}^{\ast}})_{G_{k}^{\ast}}=\mathbf{B}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}(\mathbf{1}_{G_{k}^{\ast}})_{G_{k}^{\ast}}+\mathbf{B}^{S}_{G_{k}^{\ast},{G_{k}^{\ast}}^{C}}(\mathbf{1}_{G_{k}^{\ast}})_{{G_{k}^{\ast}}^{C}}=0.$
On the other hand, since the off-diagonal block $\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}$ may be nonzero for $k\neq l$, we can only conclude as (33):
$(\mathbf{B}^{S}\mathbf{1}_{G_{l}^{\ast}})_{G_{k}^{\ast}}=\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}.$
2. Necessary condition for $\boldsymbol{\alpha}^{S}$.
Our goal is to derive condition (14). Only using the definition of $\mathbf{W}^{S}$ given in Lemma 13, we compute:
$\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}$ $=\lambda^{S}\mathbf{1}_{G_{k}^{\ast}}+\frac{1}{2}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}\mathbf{1}_{G_{k}^{\ast}}+\frac{1}{2}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}\mathbf{1}_{G_{k}^{\ast}}-(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})\mathbf{1}_{G_{k}^{\ast}}-\mathbf{B}^{S}\mathbf{1}_{G_{k}^{\ast}}$
$=\lambda^{S}\mathbf{1}_{G_{k}^{\ast}}+\frac{1}{2}\bigl{(}\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr{)}\mathbf{1}_{n}+\frac{|G_{k}^{\ast}|}{2}\boldsymbol{\alpha}^{S}-(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})\mathbf{1}_{G_{k}^{\ast}}-\mathbf{B}^{S}\mathbf{1}_{G_{k}^{\ast}}.$ (34)
Then the restriction of $\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}$ to the indices in $G_{k}^{\ast}$ is given by:
$(\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}})_{G_{k}^{\ast}}\overset{(i)}{=}\lambda^{S}\mathbf{1}_{|G_{k}^{\ast}|}+\frac{1}{2}\bigl{(}\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr{)}\mathbf{1}_{|G_{k}^{\ast}|}+\frac{|G_{k}^{\ast}|}{2}\boldsymbol{\alpha}^{S}_{G_{k}^{\ast}}-|G_{k}^{\ast}|\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}_{S,k}^{\ast}\overset{(ii)}{=}0,$ (35)
where step $(i)$ uses (28) and (32), which require (C4), and step $(ii)$ uses (30), which requires (C2) and (C3). Rearranging the terms, this computation implies:
$\boldsymbol{\alpha}^{S}_{G_{k}^{\ast}}=2\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}_{S,k}^{\ast}-\biggl{\{}\frac{2\lambda^{S}}{|G_{k}^{\ast}|}+\frac{1}{|G_{k}^{\ast}|}\bigl{(}\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr{)}\biggr{\}}\mathbf{1}_{|G_{k}^{\ast}|}.$ (36)
It remains to express $\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}$ on the right-hand side solely in terms of the data and $\lambda^{S}$. To this end, we compute:
$\mathbf{1}_{G_{k}^{\ast}}^{\top}\mathbf{W}^{S}\mathbf{1}_{G_{k}^{\ast}}$ =$\overset{(i)}{=}\mathbf{1}_{G_{k}^{\ast}}^{\top}\biggl{\{}\lambda^{S}\mathbf{1}_{G_{k}^{\ast}}+\frac{1}{2}\bigl{(}\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr{)}\mathbf{1}_{n}+\frac{|G_{k}^{\ast}|}{2}\boldsymbol{\alpha}^{S}-(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})\mathbf{1}_{G_{k}^{\ast}}-\mathbf{B}^{S}\mathbf{1}_{G_{k}^{\ast}}\biggr{\}}$
=$\overset{(ii)}{=}\lambda^{S}|G_{k}^{\ast}|+|G_{k}^{\ast}|\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}-|G_{k}^{\ast}|^{2}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}$ (37)
=$\overset{(iii)}{=}0,$
where step $(i)$ uses (34) step $(ii)$ uses (31), which require (C4), and step $(iii)$ uses (29), which require (C2) and (C3). Rearranging the terms, this computation implies:
$\sum_{i\in G_{k}^{\ast}}\boldsymbol{\alpha}^{S}_{i}=|G_{k}^{\ast}|\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}-\lambda^{S}.$ (38)
This with (36) implies:
$\boldsymbol{\alpha}^{S}_{G_{k}^{\ast}}=2\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}_{S,k}^{\ast}-\biggl{\{}\frac{\lambda^{S}}{|G_{k}^{\ast}|}+\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}\biggr{\}}\mathbf{1}_{|G_{k}^{\ast}|}.$
This corresponds precisely to condition 14.
3. Necessary condition for $\mathbf{B}^{S}$.
Our goal is to derive condition (15). Fix $1\leq k\neq l\leq K$. Similar to (34), only using the definition of $\mathbf{W}^{S}$ given in Lemma 13, we compute:
$\mathbf{W}^{S}\mathbf{1}_{G_{l}^{\ast}}=\lambda^{S}\mathbf{1}_{G_{l}^{\ast}}+\frac{1}{2}\bigl{(}\sum_{i\in G_{l}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr{)}\mathbf{1}_{n}+\frac{|G_{l}^{\ast}|}{2}\boldsymbol{\alpha}^{S}-(\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot})\mathbf{1}_{G_{l}^{\ast}}-\mathbf{B}^{S}\mathbf{1}_{G_{l}^{\ast}}.$
Then, the restriction of $\mathbf{W}^{S}\mathbf{1}_{G_{l}^{\ast}}$ to the indices in $G_{k}^{\ast}$ is given by:
$(\mathbf{W}^{S}\mathbf{1}_{G_{l}^{\ast}})_{G_{k}^{\ast}}\overset{(i)}{=}\frac{1}{2}\bigl{(}\sum_{i\in G_{l}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr{)}\mathbf{1}_{|G_{k}^{\ast}|}+\frac{|G_{l}^{\ast}|}{2}\boldsymbol{\alpha}^{S}_{G_{k}^{\ast}}-|G_{l}^{\ast}|\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}\overset{(ii)}{=}0,$ (39)
where step $(i)$ uses $(\mathbf{1}_{G_{l}^{\ast}})_{G_{k}^{\ast}}=0$, (28), and (33), and step $(ii)$ uses (30), which requires (C2) and (C3). Rearranging the terms, this computation implies:
$\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}=\frac{1}{2}\bigl{(}\sum_{i\in G_{l}^{\ast}}\boldsymbol{\alpha}^{S}_{i}\bigr{)}\mathbf{1}_{|G_{k}^{\ast}|}+\frac{|G_{l}^{\ast}|}{2}\boldsymbol{\alpha}^{S}_{G_{k}^{\ast}}-|G_{l}^{\ast}|\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}^{\ast}_{S,l}.$ (40)
Using (38) with $k$ replaced by $l$, and (36), which require (C2) and (C3), this computation implies:
$\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|}=$$\biggl{\{}\frac{|G_{l}^{\ast}|}{2}\|\overline{\mathbf{X}}^{\ast}_{S,l}\|_{2}^{2}-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{k}^{\ast}|}\lambda^{S}-\frac{|G_{l}^{\ast}|}{2}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}\biggr{\}}\mathbf{1}_{|G_{k}^{\ast}|}$
$+|G_{l}^{\ast}|\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}_{S,k}^{\ast}-|G_{l}^{\ast}|\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}^{\ast}_{S,l}.$
This implies that, for $i\in G_{k}^{\ast}$, the $i$th row sum of $\mathbf{B}^{S}_{G_{k}^{\ast}G_{l}^{\ast}}$ is expressed as:
$(\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{1}_{|G_{l}^{\ast}|})_{i}\overset{(i)}{=}$ $-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{k}^{\ast}|}\lambda^{S}+\frac{|G_{l}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}\|_{2}^{2}-2\mathbf{X}_{S,i}^{\top}\overline{\mathbf{X}}^{\ast}_{S,l}+\|\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)}$
$-\frac{|G_{l}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}-2\mathbf{X}_{S,i}^{\top}\overline{\mathbf{X}}_{S,k}^{\ast}+\|\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)},$
$=$ $-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{k}^{\ast}|}\lambda^{S}+\frac{|G_{l}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)},$
where step $(i)$ adds and subtracts $\frac{|G_{l}^{\ast}|}{2}\|\mathbf{X}_{S,i}\|_{2}^{2}$. This concludes the derivation of condition (15) and completes the proof of Lemma 14. ∎
Remark 28.
If we construct $\boldsymbol{\alpha}^{S}$ and $\mathbf{B}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}$ according to conditions (14) and (15), then step (ii) in (35) and (39) hold without requiring conditions (C2) and (C3), as the construction is explicitly designed to satisfy these equalities. This observation is useful in the proof of Lemma 17, presented in Appendix C.4.4.
C.4.3 Proof of Lemma 16
Let $S\subset[p]$ be a support set, and consider the corresponding primal SDP problem (13). We show that constructing $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$, $\dot{\mathbf{B}}^{S}(\lambda^{S})$ and $\dot{\mathbf{W}}^{S}(\lambda^{S})$ as in Lemma 15 guarantees the following:
  1. 1.
    Conditions (C3) and (C4) in Lemma 13 are satisfied for any $\lambda^{S}$,
  2. 2.
    Conditions (C1) and (C5) are satisfied for $\lambda^{S}<U^{S}(\mathbf{X})$.
Proof.
For notational simplicity, we omit the notation $\lambda^{S}$ from $\dot{\boldsymbol{\alpha}}^{S}(\lambda^{S})$, $\dot{\mathbf{B}}^{S}(\lambda^{S})$, and $\dot{\mathbf{W}}^{S}(\lambda^{S})$ and denote them $\dot{\boldsymbol{\alpha}}^{S}$, $\dot{\mathbf{B}}^{S}$, and $\dot{\mathbf{W}}^{S}$, respectively.
1. Conditions (C3) and (C4).
Condition (C4) follows directly from the block-diagonal structure of both $\mathbf{Z}^{\ast}=\sum_{k=1}^{K}|G_{k}^{\ast}|^{-1}\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top}$ and $\dot{\mathbf{B}}^{S}$. To verify condition (C3), which requires that $\langle\dot{\mathbf{W}}^{S},\mathbf{Z}^{\ast}\rangle=0$, we use the definition $\mathbf{Z}^{\ast}=\sum_{k=1}^{K}|G_{k}^{\ast}|^{-1}\mathbf{1}_{G_{k}^{\ast}}\mathbf{1}_{G_{k}^{\ast}}^{\top}$ to expand the inner product as
$\langle\dot{\mathbf{W}}^{S},\mathbf{Z}^{\ast}\rangle=\frac{1}{|G_{k}^{\ast}|}\sum_{k=1}^{K}\mathbf{1}_{G_{k}^{\ast}}^{\top}\dot{\mathbf{W}}^{S}\mathbf{1}_{G_{k}^{\ast}}.$
Thus, it suffices to verify that $\mathbf{1}_{G_{k}^{\ast}}^{\top}\dot{\mathbf{W}}^{S}\mathbf{1}_{G_{k}^{\ast}}=0$ for all $k=1,\ldots,K$. Since our construction sets $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}=\mathbf{0}$, we can invoke (37) to state that
$\mathbf{1}_{G_{k}^{\ast}}^{\top}\dot{\mathbf{W}}^{S}\mathbf{1}_{G_{k}^{\ast}}=\lambda^{S}|G_{k}^{\ast}|+|G_{k}^{\ast}|\sum_{i\in G_{k}^{\ast}}\bigl{(}\dot{\boldsymbol{\alpha}}^{S}(\lambda_{S})\bigr{)}_{i}-|G_{k}^{\ast}|^{2}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}.$
Now using our construction of $\dot{\boldsymbol{\alpha}}^{S}(\lambda_{S})$, recalled below:
$\dot{\boldsymbol{\alpha}}^{S}_{G_{k}^{\ast}}=2\mathbf{X}_{S,G_{k}^{\ast}}^{\top}\overline{\mathbf{X}}^{\ast}_{S,k}-\frac{\lambda^{S}}{|G_{k}^{\ast}|}\mathbf{1}_{|G_{k}^{\ast}|}-\|\overline{\mathbf{X}}^{\ast}_{S,k}\|_{2}^{2}\mathbf{1}_{|G_{k}^{\ast}|},$
we can conclude:
$\mathbf{1}_{G_{k}^{\ast}}^{\top}\dot{\mathbf{W}}^{S}\mathbf{1}_{G_{k}^{\ast}}=\lambda^{S}|G_{k}^{\ast}|+2|G_{k}^{\ast}|^{2}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}-\lambda^{S}|G_{k}^{\ast}|-|G_{k}^{\ast}|^{2}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}-|G_{k}^{\ast}|^{2}\|\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}=0.$
This concludes the verification of condition (C3).
2. Conditions (C1) and (C5).
We now show that the condition $\lambda^{S}<U^{S}(\mathbf{X})$ ensures conditions (C1) and (C5). Since our construction sets the diagonal block $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}$ as a zero matrix, it suffices to show that if $\lambda^{S}<U^{S}(\mathbf{X})$, then $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}>0$ for all $1\leq k\neq l\leq K$. By our construction in (17), for $i\in G_{k}^{\ast}$ and $j\in G_{l}^{\ast}$, the $(i,j)$th element of $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}$ is
$\frac{r_{i}^{(k,l,S)}\,r_{j}^{(l,k,S)}}{\sum_{j\in G_{l}^{\ast}}r_{j}^{(l,k,S)}.}$ (41)
where
$r_{i}^{(k,l,S)}=-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{k}^{\ast}|}\lambda^{S}+\frac{|G_{l}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)},$
and
$r_{j}^{(l,k,S)}=-\frac{|G_{l}^{\ast}|+|G_{k}^{\ast}|}{2|G_{l}^{\ast}|}\lambda^{S}+\frac{|G_{k}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,j}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,j}\|_{2}^{2}\bigr{)}.$
Therefore, all elements of $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}$ for all $1\leq k\neq l\leq K$ are strictly positive if, for all $1\leq k\neq l\leq K$ and $i\in G_{k}^{\ast}$, the inequality $r_{i}^{(k,l,S)}>0$ holds, or equivalently,
$\lambda^{S}<\frac{|G_{l}^{\ast}||G_{k}^{\ast}|}{|G_{k}^{\ast}|+|G_{l}^{\ast}|}\bigl{(}\|\overline{\mathbf{X}}_{S,l}^{\ast}-\mathbf{X}_{i}\|_{2}^{2}-\|\overline{\mathbf{X}}_{S,k}^{\ast}-\mathbf{X}_{i}\|_{2}^{2}\bigr{)}.$
This condition is precisely $\lambda^{S}<U^{S}(\mathbf{X})$, completing the proof of Lemma 16. ∎
C.4.4 Proof of Lemma 17
The proof proceeds in four steps:
  1. 1.
    Reduction to uniform control of a quadratic form over a subspace (equation (42)),
  2. 2.
    Derivation of $L^{S}_{1}(\mathbf{X})$ (equation (44)),
  3. 3.
    Derivation of $t_{lk}^{S}(\lambda^{S})$ (equation (45)),
  4. 4.
    Derivation of $T^{S}_{1,kl}$, $T^{S}_{2,kl}$, and $T^{S}_{3,kl}$ (equation 46).
Proof.
Given $\lambda_{S}$, we construct $\dot{\mathbf{B}}(\lambda_{S})$, $\dot{\boldsymbol{\alpha}}^{S}(\lambda_{S})$, and $\dot{\mathbf{W}}^{S}(\lambda_{S})$ according to Lemma 15 and denote them by $\dot{\boldsymbol{\alpha}}^{S}$, $\dot{\mathbf{B}}^{S}$, and $\dot{\mathbf{W}}^{S}$, respectively, for notational simplicity.
1. Reduction to uniform control of a quadratic form over a subspace.
By Lemma 16, these constructed dual variables satisfy condition (C4) of Lemma 13. Therefore, by Remark 28, for any $k=1,\ldots,K$, we have $\bigl{(}\dot{\mathbf{W}}^{S}\mathbf{1}_{G_{k}^{\ast}}\bigr{)}_{G_{k}^{\ast}}=\mathbf{0},$ from (14), and for all $l\neq k$, we have $(\dot{\mathbf{W}}^{S}\mathbf{1}_{G_{k}^{\ast}})_{G_{l}^{\ast}}=\mathbf{0}$ by (15). Therefore, we have $\dot{\mathbf{W}}^{S}\mathbf{1}_{G_{k}^{\ast}}=\mathbf{0}$ for $k=1,\ldots,K$. This implies that $\mathbf{1}_{G_{1}^{\ast}},\ldots,\mathbf{1}_{G_{K}^{\ast}}$ are eigenvectors of $\dot{\mathbf{W}}^{S}$ corresponding to zero eigenvalues. Since any vector in $\mathbb{R}^{n}$ can be decomposed into the sum of its projections onto $\operatorname{span}\left(\mathbf{1}_{G_{k}^{\ast}}:k\in[K]\right)$ and its orthogonal complement, denoted as $\Gamma_{K}$, to show that $\dot{\mathbf{W}}^{S}\succeq 0$, it suffices to prove that
$\mathbf{v}^{\top}\dot{\mathbf{W}}^{S}\mathbf{v}\geq 0\quad\text{for all }\mathbf{v}\in\Gamma_{K}\text{ such that }\|\mathbf{v}\|_{2}=1.$ (42)
2. Derivation of $L^{S}_{1}(\mathbf{X})$.
By the definition of $\Gamma_{k}$, $\mathbf{v}\in\Gamma_{k}$ implies $\mathbf{v}^{\top}\mathbf{1}_{G_{1}^{\ast}}=\mathbf{v}^{\top}\mathbf{1}_{G_{2}^{\ast}}=\ldots=\mathbf{v}^{\top}\mathbf{1}_{G_{K}^{\ast}}=0,$ which implies
$\sum_{j\in G_{k}^{\ast}}^{n}v_{j}=\sum_{i=1}^{n}v_{i}=0~{}\text{for all}~{}k=1,\ldots,K.$ (43)
Now we compute the quadratic form $\mathbf{v}^{\top}\dot{\mathbf{W}}^{S}\mathbf{v}$ for $\mathbf{v}\in\Gamma_{K}$ such that $\|\mathbf{v}\|_{2}=1$:
$\mathbf{v}^{\top}\dot{\mathbf{W}}^{S}\mathbf{v}$ =$\overset{(i)}{=}\lambda^{S}\mathbf{v}^{\top}\textbf{I}_{n}\mathbf{v}+\frac{1}{2}\mathbf{v}^{\top}\mathbf{1}_{n}(\boldsymbol{\alpha}^{S})^{\top}\mathbf{v}+\frac{1}{2}\mathbf{v}^{\top}\boldsymbol{\alpha}^{S}\mathbf{1}_{n}^{\top}\mathbf{v}-\mathbf{v}^{\top}\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}\mathbf{v}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}$
=$\overset{(ii)}{=}\lambda^{S}+(\mathbf{v}^{\top}\boldsymbol{\alpha}_{S})\sum_{i=1}^{n}v_{i}-\|\mathbf{X}_{S,\cdot}\mathbf{v}\|_{2}^{2}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}$
=$\overset{(ii)}{=}\lambda^{S}-\|\mathbf{X}_{S,\cdot}\mathbf{v}\|_{2}^{2}-\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v},$
where step $(i)$ uses the construction $\dot{\mathbf{W}}^{S}=\lambda^{S}\textbf{I}_{n}+\frac{1}{2}\mathbf{1}_{n}(\dot{\boldsymbol{\alpha}}^{S})^{\top}+\frac{1}{2}\dot{\boldsymbol{\alpha}}^{S}\mathbf{1}_{n}^{\top}-\mathbf{X}_{S,\cdot}^{\top}\mathbf{X}_{S,\cdot}-\dot{\mathbf{B}}^{S}$, step $(ii)$ uses $\mathbf{v}^{\top}\textbf{I}_{n}\mathbf{v}=\|\mathbf{v}\|^{2}=1$, and step $(ii)$ uses (43). Using our model and notation in Definition 12,
$\mathbf{X}_{S,\cdot}\mathbf{v}=\sum_{i=1}^{n}v_{i}\mathbf{X}_{S,i}=\sum_{k=1}^{K}\biggl{(}\sum_{i\in G_{k}^{*}}v_{i}(\boldsymbol{\mu}_{k})_{S}+\sum_{i=1}^{n}v_{i}\mathcal{E}_{S,i}\biggr{)}\overset{(i)}{=}\sum_{i=1}^{n}v_{i}\mathcal{E}_{S,i}=\mathcal{E}_{S,\cdot}\mathbf{v},$
where step $(i)$ uses (43). Therefore, a sufficient condition for (42) is:
$\lambda^{S}\geq\underbrace{\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}}_{L^{S}_{1}(\mathbf{X})}+\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}.$ (44)
This completes the derivation of $L^{S}_{1}(\mathbf{X})$.
3. Derivation of $t_{lk}^{S}(\lambda^{S})$.
Now we start to derive $L_{2}^{S}(\mathbf{X},\lambda^{S})$. Recall from Lemma 15 that given $\lambda^{S}$, we construct $\dot{\mathbf{B}}$ as a block-diagonal matrix, where the diagonal blocks are zero matrices and for $1\leq k\neq l\leq K$, the $(k,l)$th off-diagonal block is defined as:
$\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}=\frac{1}{\mathbf{1}_{|G_{l}^{\ast}|}^{\top}\mathbf{r}^{(l,k,S)}}\,\mathbf{r}^{(k,l,S)}\,(\mathbf{r}^{(l,k,S)})^{\top},$
with $\mathbf{r}^{(k,l,S)}\in\mathbb{R}^{|G_{k}^{\ast}|}$ defined componentwise by
$r_{i}^{(k,l,S)}=-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{k}^{\ast}|}\lambda^{S}+\frac{|G_{l}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)}.$
Using this definition, we compute $\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}$ in (44):
$\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}\overset{(i)}{=}\sum_{1\leq k\neq l\leq K}\mathbf{v}_{G_{k}^{\ast}}^{\top}\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}\mathbf{v}_{G_{l}^{\ast}}\overset{(ii)}{=}\sum_{1\leq k\neq l\leq K}\sum_{i\in G_{k}^{\ast}}\sum_{j\in G_{l}^{\ast}}\frac{(\mathbf{v}_{G_{k}^{\ast}}^{\top}\mathbf{r}^{(k,l,S)})(\mathbf{v}_{G_{l}^{\ast}}^{\top}\mathbf{r}^{(l,k,S)})}{\underbrace{\mathbf{1}_{|G_{l}^{\ast}|}^{\top}\mathbf{r}^{(l,k,S)}}_{=t_{lk}^{S}(\lambda^{S})}},$ (45)
where step $(i)$ uses $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{k}^{\ast}}=\mathbf{0}$ for $k=1,\ldots,K$, and step $(ii)$ uses our construction of $\dot{\mathbf{B}}^{S}_{G_{k}^{\ast},G_{l}^{\ast}}$.
4. Derivation of $T^{S}_{1,kl}$, $T^{S}_{2,kl}$, and $T^{S}_{3,kl}$
From (45), we analyze one term in the numerator:
$\mathbf{v}_{G_{k}^{\ast}}^{\top}\mathbf{r}^{(k,l,S)}$
$\quad=\sum_{i\in G_{k}^{\ast}}v_{i}r_{i}^{(k,l,S)}$
$\quad=\sum_{i\in G_{k}^{\ast}}v_{i}\left\{-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{k}^{\ast}|}\lambda^{S}+\frac{|G_{l}^{\ast}|}{2}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)}\right\}$
=$\quad\stackrel{{\scriptstyle(i)}}{{=}}\frac{|G_{l}^{\ast}|}{2}\sum_{i\in G_{k}^{\ast}}v_{i}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}-\mathbf{X}_{S,i}\|_{2}^{2}-\|\overline{\mathbf{X}}^{\ast}_{S,k}-\mathbf{X}_{S,i}\|_{2}^{2}\bigr{)}$
$\quad=\frac{|G_{l}^{\ast}|}{2}\sum_{i\in G_{k}^{\ast}}v_{i}\bigl{(}\|\overline{\mathbf{X}}^{\ast}_{S,l}\|_{2}^{2}+\|\mathbf{X}_{S,i}\|_{2}^{2}-2\langle\overline{\mathbf{X}}^{\ast}_{S,l},\mathbf{X}_{S,i}\rangle-\|\overline{\mathbf{X}}^{\ast}_{S,k}\|_{2}^{2}-\|\mathbf{X}_{S,i}\|_{2}^{2}+2\langle\overline{\mathbf{X}}^{\ast}_{S,k},\mathbf{X}_{S,i}\rangle\bigr{)}$
=$\quad\overset{(ii)}{=}|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle\overline{\mathbf{X}}^{\ast}_{S,k}-\overline{\mathbf{X}}^{\ast}_{S,l},\;\mathbf{X}_{S,i}\rangle$
=$\quad\overset{(iii)}{=}|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S}+\overline{\mathcal{E}}_{S,k}^{\ast}-\overline{\mathcal{E}}_{S,l}^{\ast},\;{\mathcal{E}}_{S,i}\rangle$
$\quad=|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\;{\mathcal{E}}_{S\cap S_{0},i}\rangle+|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle\overline{\mathcal{E}}_{S,k}^{\ast}-\overline{\mathcal{E}}_{S,l}^{\ast},\;{\mathcal{E}}_{S,i}\rangle$
$\quad=I^{S}_{kl}(\mathbf{v},\mathbf{X})+N_{kl}^{S}(\mathbf{v},\mathbf{X})$
where step $(i)$ and $(ii)$ use $\sum_{i\in G_{k}^{\ast}}v_{i}=0$ for $k=1,\ldots,K$, and step $(iii)$ comes from our model and notation in Definition 12. Similarly, we have
$\mathbf{v}_{G_{l}^{\ast}}^{\top}\mathbf{r}^{(l,k,S)}=I^{S}_{lk}(\mathbf{v},\mathbf{X})+N_{lk}^{S}(\mathbf{v},\mathbf{X}).$
Therefore, we have
$\mathbf{v}^{\top}\dot{\mathbf{B}}^{S}\mathbf{v}$ $=\sum_{1\leq k\neq l\leq K}\sum_{i\in G_{k}^{\ast}}\sum_{j\in G_{l}^{\ast}}\frac{1}{t_{lk}^{S}(\lambda^{S})}\bigl{(}I^{S}_{kl}(\mathbf{v},\mathbf{X})+N_{kl}^{S}(\mathbf{v},\mathbf{X})\bigr{)}\bigl{(}I^{S}_{lk}(\mathbf{v},\mathbf{X})+N_{lk}^{S}(\mathbf{v},\mathbf{X})\bigr{)}$
$=\sum_{1\leq k\neq l\leq K}\sum_{i\in G_{k}^{\ast}}\sum_{j\in G_{l}^{\ast}}\frac{1}{t_{lk}^{S}(\lambda^{S})}\bigl{(}T^{S}_{1,kl}(\mathbf{v})+T^{S}_{2,kl}(\mathbf{v})+T^{S}_{3,kl}(\mathbf{v})\bigr{)}.$ (46)
Plugging this into (44) completes the proof of Lemma 17.
C.4.5 Proof of Lemma 19
For notational simplicity, we assume $\sigma^{2}=1$ here. The proof proceeds in the following steps:
  1. 1.
    Decomposition of $D^{S}_{kli}(\mathbf{X})$ into Chi-squares, Gaussian and inner product of Gaussians,
  2. 2.
    Bounding the Gaussian and Chi-squares using the tail bounds,
  3. 3.
    Bounding the Gaussian inner product using conditioning, Gaussian tail bounds and separation condition,
  4. 4.
    Defining the uniform bound event,
  5. 5.
    Bounding the probability of the uniform bound event.
Proof.
We proceed by following the steps outlined above in sequence.
1. Decomposition of $D^{S}_{kli}(\mathbf{X})$ into Chi-squares, Gaussian and inner product of Gaussians.
Fix indices $1\leq k\neq l\leq K$ and choose $i\in G_{k}^{\ast}$. Based on our model and notation in Definition 12, we consider the following decomposition of $D^{S}_{kli}(\mathbf{X})$:
$D^{S}_{kli}(\mathbf{X})~{}=~{}$ $\|\mathbf{X}_{S,i}-\overline{\mathbf{X}}_{S,l}^{\ast}\|_{2}^{2}-\|\mathbf{X}_{S,i}-\overline{\mathbf{X}}_{S,k}^{\ast}\|_{2}^{2}$
[…]
=$\overset{(i)}{=}~{}$ $\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\|\mathcal{E}_{S,i}\|_{2}^{2}+\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}+2\left\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S},\mathcal{E}_{S,i}\right\rangle$
$-2\left\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\overline{\mathcal{E}}_{S\cap S_{0},l}^{\ast}\right\rangle-2\left\langle\overline{\mathcal{E}}_{S,l}^{\ast},\mathcal{E}_{S,i}\right\rangle$
$-\left(\frac{|G_{k}^{\ast}|-1}{|G_{k}^{\ast}|}\right)^{2}\left(\|\mathcal{E}_{S,i}\|_{2}^{2}+\|\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\|_{2}^{2}-2\left\langle\mathcal{E}_{S,i},\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\right\rangle\right)$
[…]
$-\left(\frac{|G_{k}^{\ast}|-1}{|G_{k}^{\ast}|}\right)^{2}\|\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\|_{2}^{2}-2\left\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\overline{\mathcal{E}}_{S\cap S_{0},l}^{\ast}\right\rangle$
$+2\left\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S}-\overline{\mathcal{E}}_{S,l}^{\ast}+\left(\frac{|G_{k}^{\ast}|-1}{|G_{k}^{\ast}|}\right)^{2}\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast},\mathcal{E}_{S,i}\right\rangle,$ (47)
where step $(i)$ uses the sparsity assumption on cluster center difference.
The six terms in this decomposition are classified into the following categories:
  • Signal: $\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}$
  • Chi-square noise:
    $\operatorname{CN}_{l}^{S}$ :=¯$:=\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2},$
    $\operatorname{CN}_{i}^{S}$ $:=\frac{2|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|^{2}}\|\mathcal{E}_{S,i}\|_{2}^{2},~{}\text{and}$
    $\operatorname{CN}_{ki}^{S}$ $:=-\left(\frac{|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|}\right)^{2}\|\overline{\mathcal{E}}_{S,k\setminus\{i\}}\|_{2}^{2}$
  • Gaussian signal-noise interaction:
    $\operatorname{GSN}_{kl}^{S}:=-2\langle(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}},\overline{\mathcal{E}}_{S\cap S_{0},l}^{\ast}\rangle$
  • Inner product of Gaussians:
    $\mathrm{IP}_{kli}^{S}:=2\left\langle\bigl{(}\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast}\bigr{)}_{S}-\overline{\mathcal{E}}_{S,l}^{\ast}+\left(\frac{|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|}\right)^{2}\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast},\;\mathcal{E}_{S,i}\right\rangle,$
In summary, we have:
$D^{S}_{kli}(\mathbf{X})~{}=~{}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+\operatorname{CN}_{l}^{S}+\operatorname{CN}_{i}^{S}+\operatorname{CN}_{ki}^{S}+\operatorname{GSN}_{kl}^{S}+\mathrm{IP}_{kli}^{S}.$ (48)
2. Bounding the Gaussian and Chi-squares using the tail inequalities.
Chi-square noise and Gaussian signal-noise interaction terms $(\operatorname{CN}_{l}^{S},\operatorname{CN}_{i}^{S},\operatorname{CN}_{ki}^{S}$, and $\operatorname{GSN}_{kl}^{S})$ in (48) are straightforward to bound from below. Define the following lower bound events:
$\mathcal{B}^{(i,1)}_{kl,S}$ :=$:=\left\{\operatorname{CN}_{l}^{S}\geq\frac{|S|-2\sqrt{|S|\zeta^{S}}}{|G_{l}^{\ast}|}\right\},$ (49)
$\mathcal{B}^{(i,2)}_{kl,S}$ :=$:=\left\{\operatorname{CN}_{i}^{S}\geq\frac{2|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|^{2}}\bigl{(}|S|-2\sqrt{|S|\zeta^{S}}\bigr{)}\right\},$ (50)
$\mathcal{B}^{(i,3)}_{kl,S}$ :=$:=\left\{\operatorname{CN}_{ki}^{S}\geq\frac{1-|G_{k}^{\ast}|}{|G_{k}^{\ast}|^{2}}\bigl{(}|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S}\bigr{)}\right\},$ (51)
$\mathcal{B}^{(i,4)}_{kl,S}$ :=$:=\left\{\operatorname{GSN}_{kl}^{S}\geq-2\sqrt{\frac{2\zeta^{S}}{|G_{l}^{\ast}|}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\right\},$ (52)
By setting $\zeta^{S}=\log(n^{2}|S|^{2}{\binom{p}{|S|}})$, Lemma 3 implies that each event has a probability of at least $1-1/(n^{2}K^{2}|S|^{2}{\binom{p}{|S|}})$, for any $S\in\mathcal{S}$.
3. Bounding the Gaussian inner product using conditioning, Gaussian tail bounds and separation condition.
Here, we use the separation condition to bound the term. If we condition on ¯$\overline{\mathcal{E}}_{S,l}^{\ast}$ and ¯$\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}$, then $-\mathrm{IP}_{kli}^{S}$ has a univariate Gaussian distribution with mean zero and variance as the following:
$\mathrm{Var}(-\mathrm{IP}_{kli}^{S}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast},\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast})$ $=\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S}-\overline{\mathcal{E}}_{S,l}^{\ast}+(1-|G_{k}^{\ast}|^{-1})^{2}\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\right\|_{2}^{2}$
=$\overset{(i)}{=}\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{2}+\underbrace{\left\|\overline{\mathcal{E}}_{S,l}^{\ast}-(1-|G_{k}^{\ast}|^{-1})^{2}\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\right\|_{2}^{2}}_{:=\operatorname{CN}_{kli}^{S}}$
$\quad+\underbrace{2\bigl{\langle}(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\;(1-|G_{k}^{\ast}|^{-1})^{2}\overline{\mathcal{E}}_{S\cap S_{0},k\setminus\{i\}}^{\ast}-\overline{\mathcal{E}}_{S\cap S_{0},l}^{\ast}\bigr{\rangle}}_{:=\operatorname{GN}_{kli}^{S}},$
where step $(i)$ uses our sparsity assumption on cluster center difference.
Next, we define events that bound this variance from above:
$\mathcal{B}^{(i,5)}_{kl,S}$ :=$:=\left\{\operatorname{CN}_{kli}^{S}\leq(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})(|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S})\right\},$
$\mathcal{B}^{(i,6)}_{kl,S}$ :=$:=\left\{\operatorname{GN}_{kli}^{S}\leq 2\sqrt{2(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})\zeta^{S}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\right\}.$
By setting $\zeta^{S}=\log(n^{2}|S|^{2}{\binom{p}{|S|}})$, Lemma 3 implies that each event has a probability of at least $1-1/(n^{2}|S|^{2}{\binom{p}{|S|}})$, uniformly overall all $S\in\mathcal{S}$.
Thus, conditioned on events $\mathcal{B}^{(i,5)}_{kl,S}$ and $\mathcal{B}^{(i,6)}_{kl,S}$, the variance can be bouneded as
$\mathrm{Var}(-\mathrm{IP}_{kli}^{S}\;|\ \;\overline{\mathcal{E}}_{S,l}^{\ast},\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast})$ $\leq\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{2}+(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})(|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S})$
$\quad+2\sigma^{2}\sqrt{2(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})\zeta^{S}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}.$
To further bound this variance, it is convenient to remove the first-order signal term $\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}$. To this end, we collect some conditions. First, recall from Theorem 1 that
$m=2\min_{1\leq k\neq l\leq K}\bigl{(}|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1}\bigr{)}^{-1}.$
From (6), recall that for $S\in\mathcal{S}$, we have
$\min_{1\leq k\neq l\leq K}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}\gtrsim\sigma^{2}\left(\log n+\frac{|S|\log p}{m}+\sqrt{\frac{|S|\log p}{m}}\right),$ (53)
and $|S|\leq\sqrt{p}$. Under these conditions, by treating $K$ as a costant, we have
$\zeta^{S}=\log(n^{2}|S|^{2}{\binom{p}{|S|}})\lesssim|S|\log p+\log n,$
which implies
$2\sqrt{2(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})\zeta^{S}}\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}\leq 2\sqrt{\frac{\zeta^{S}}{m}}\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}\lesssim\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{2}.$
Therefore, we can further bound the variance as
$\mathrm{Var}(-\mathrm{IP}_{kli}^{S})$ $\lesssim\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{2}+(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})(|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S}).$
Therefore, by the Gaussian tail inequality, we have
$\mathrm{CP}_{kli}^{S}$ $:=\mathbb{P}\left(-\mathrm{IP}_{kli}^{S}>\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|^{2}_{2}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast},\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast}\right)$
$\leq\exp\left(\frac{-\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{4}}{2\mathrm{Var}(-\mathrm{IP}_{kli}^{S}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast},\overline{\mathcal{E}}_{S,k\setminus\{i\}}^{\ast})}\right)$
$\lesssim\exp\left(\frac{-\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{4}}{\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{2}+2(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})(|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S})}\right),$ (54)
where the last step uses the variance bound. We want to bound the last term by $C/n^{2}$, with $C$ a constant. To this end, it suffices to show
$\frac{\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{4}}{\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|^{2}+2(|G_{l}^{\ast}|^{-1}+|G_{k}^{\ast}|^{-1})(|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S})}\gtrsim\log n,$
which is indeed satisfied under our separation condition in (53).
4. Defining the uniform bound event.
We collect the bounds in (49), (50), (51), (52), and (54). From (48), recalled below,
$D^{S}_{kli}(\mathbf{X})~{}=~{}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+\operatorname{CN}_{l}^{S}+\operatorname{CN}_{i}^{S}+\operatorname{CN}_{ki}^{S}+\operatorname{GSN}_{kl}^{S}+\mathrm{IP}_{kli}^{S}.$
we can say
$D^{S}_{kli}(\mathbf{X})~{}\gtrsim$ $~{}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\frac{|S|-2\sqrt{|S|\zeta^{S}}}{|G_{l}^{\ast}|}+\frac{2|G^{\ast}_{k}|-1}{|G^{\ast}_{k}|^{2}}\bigl{(}|S|-2\sqrt{|S|\zeta^{S}}\bigr{)}$
$~{}+~{}\frac{1-|G_{k}^{\ast}|}{|G_{k}^{\ast}|^{2}}\bigl{(}|S|+2\sqrt{|S|\zeta^{S}}+2\zeta^{S}\bigr{)}-2\sqrt{\frac{2\zeta^{S}}{|G_{l}^{\ast}|}}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}$
$~{}+~{}\left\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\right\|_{2}^{2},$
[…]
$~{}-~{}\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}|S|-\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}\sqrt{|S|\zeta^{S}}-\frac{1}{|G_{k}^{\ast}|}\zeta^{S}$
[…]
where
$R^{S}_{kl}=\sqrt{\frac{\zeta^{S}}{|G_{l}^{\ast}|}}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}+\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}\sqrt{|S|\zeta^{S}}+\frac{1}{|G_{k}^{\ast}|}\zeta^{S}.$
As a result, we define an event
$\mathcal{A}^{S}_{kli}:=\Bigg{\{}D^{S}_{kli}(\mathbf{X})\gtrsim\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}|S|-R^{S}_{kl}\Bigg{\}},$ (55)
with the index $(k,l,i)$ ranging $1\leq k\neq l\leq K$ and all $i\in G^{*}_{k}$, and $S$ ranging over $\mathcal{S}$. The proof is complete once we show that the intersection event
$\bigcap_{1\leq k\neq l\leq K}\bigcap_{i\in G_{k}^{\ast}}\bigcap_{S\in\mathcal{S}}\mathcal{A}^{S}_{kli}$
holds with high probability.
5. Bounding the probability of the uniform bound event.
To prove the lemma, it suffices to show that
$\mathbb{P}\left(\bigcup_{1\leq k\neq l\leq K}\bigcup_{i\in G_{k}^{\ast}}\bigcup_{S\in\mathcal{S}}{\bigl{(}\mathcal{A}^{S}_{kli}\bigr{)}}^{C}\right)\lesssim\frac{1}{n}.$
To show this, first denote
$\mathcal{B}^{(i)}_{kl,S}:=\bigcap_{\ell=1}^{6}\mathcal{B}^{(i,\ell)}_{kl,S}.$
Then, we have
$\mathbb{P}\biggl{(}\bigcup_{1\leq k\neq l\leq K}$ $\bigcup_{i\in G_{k}^{\ast}}\bigcup_{S\in\mathcal{S}}{(\mathcal{A}^{S}_{kli})^{C}}\biggr{)}$
$\leq\sum_{1\leq k\neq l\leq K}\sum_{i\in G_{k}^{\ast}}\sum_{S\in\mathcal{S}}\mathbb{P}{\bigl{(}(\mathcal{A}^{S}_{kli})^{C}\bigr{)}}$
$=\sum_{1\leq k\neq l\leq K}\sum_{i\in G_{k}^{\ast}}\sum_{S\in\mathcal{S}}\mathbb{E}\bigl{[}\mathbb{P}\bigl{(}(\mathcal{A}^{S}_{kli})^{C}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast},\mathcal{E}_{S,i},\overline{\mathcal{E}}_{S,k\setminus\{i\}}\bigr{)}\bigr{]}$
$\leq\sum_{1\leq k\neq l\leq K}\sum_{i\in G_{k}^{\ast}}\sum_{S\in\mathcal{S}}\mathbb{E}\bigl{[}\mathbb{P}\bigl{(}(\mathcal{A}^{S}_{kli})^{C}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast},\mathcal{E}_{S,i},\overline{\mathcal{E}}_{S,k\setminus\{i\}}\bigr{)}\mathbbm{1}(\mathcal{B}^{(i)}_{kl,S})+\mathbbm{1}(\mathcal{B}^{(i)}_{kl,S})^{C}\bigr{]}$
$\overset{(i)}{\leq}\left(\sum_{1\leq k\neq l\leq K}\sum_{i\in G_{k}^{\ast}}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{S\in\mathcal{S},|S|=q}\mathbb{E}\bigl{[}\mathrm{CP}_{kli}^{S}\mathbbm{1}(\mathcal{B}^{(i)}_{kl,S})\bigr{]}\right)+\frac{\pi^{2}}{n}$
$\overset{(ii)}{\lesssim}\frac{K^{2}\pi^{2}}{6n}+\frac{K^{2}\pi^{2}}{n},$
where step $(i)$ uses $\sum_{q=1}^{\infty}(1/q^{2})=\pi^{2}/6$, and step $(ii)$ uses (54). Treating $K$ as a constant, we conclude that
$D^{S}_{kli}(\mathbf{X})\gtrsim\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}|S|-R^{S}_{kl}.$
uniformly over $S\in\mathcal{S}$, with probability at least $1-C/n$, where $C>0$ is a constant. This completes the proof of Lemma 19.
C.4.6 Proof of Lemma 20
Proof.
Given a support set $S\subset[p]$ with $|S|\leq\sqrt{p}$, we have
$L_{1}^{S}(\mathbf{X})=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}\leq\sup_{\mathbf{v}\in\mathbb{R}^{n},\|\mathbf{v}\|_{2}=1}\|\mathcal{E}_{S,\cdot}\mathbf{v}\|_{2}^{2}=\|\mathcal{E}_{S,\cdot}\|_{\mathrm{op}}^{2}.$
Here, $\mathcal{E}_{S,\cdot}$ is a $|S|\times n$ random matrix with i.i.d. Gaussian entries distributed as $\mathcal{N}(0,\sigma^{2})$. Therefore, by Lemma 7, if we define an event
$\mathcal{C}_{L}^{S}:=\bigl{\{}L_{1}^{S}(\mathbf{X})\leq\sigma^{2}(n+|S|+\xi^{S})\bigr{\}},$ (56)
where $\xi^{S}=\log(n|S|^{2}{\binom{p}{|S|}})$, then we have $\mathbb{P}(\mathcal{C}_{L}^{S})\geq 1-1/(n|S|^{2}{\binom{p}{|S|}}).$ Now we show that the event $\mathcal{C}_{L}^{S}$ holds uniformly across all $S\subset[p]$ such that $|S|\leq\sqrt{p}$, with probability at least $1-C_{8}/n$, where $C_{8}>0$ is a constant.
$\mathbb{P}\biggl{(}\bigcup_{S\subset[p],|S|\leq\sqrt{p}}(\mathcal{C}_{L}^{S})^{C}\biggr{)}\leq\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{P}((\mathcal{C}_{L}^{S})^{C})\leq\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\frac{1}{n|S|^{2}{\binom{p}{|S|}}}=\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\frac{1}{nq^{2}}\leq\frac{\pi^{2}}{6n},$
where the last step uses $\sum_{q=1}^{\infty}=\pi^{2}/6.$ Since $|S|\leq\sqrt{p}$, $|S|\log p$ dominates $\log|S|$. Therefore, we conclude that uniformly across $S\in[p]$ such that $|S|\leq\sqrt{p}$, with probability at least $1-C_{8}/n$, it holds that
$L_{1}^{S}(\mathbf{X})\lesssim\sigma^{2}(n+\log n+|S|\log p).$
This completes the proof of Lemma 20. ∎
C.4.7 Proof of Lemma 21
The proof proceeds in the following steps:
  1. 1.
    The first inequality $L_{1}^{S}(\mathbf{X}).\lesssim\dot{\lambda}^{S}$.
  2. 2.
    The second inequality $\dot{\lambda}^{S}\lesssim U^{S}(\mathbf{X})$.
Proof.
For simplicity, we assume $\sigma^{2}=1$ here.
1. The first inequality.
We aim to show $L_{1}^{S}(\mathbf{X})\lesssim\dot{\lambda}^{S}$. From (6), recall that for $S\in\mathcal{S}$, we have
$\Delta_{S\cap S_{0}}^{2}=\min_{1\leq k\neq l\leq K}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}\gtrsim\sigma^{2}\left(\log n+\frac{|S|\log p}{m}+\sqrt{\frac{|S|\log p}{m}}\right),$ (57)
where
$m=2\min_{1\leq k\neq l\leq K}\left\{\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\right\}.$
Therefore, with probability at least $1-C_{8}/n$ uniformly across all $S\in\mathcal{S}$, where $C_{8}>0$ is a constant from Lemma 20, it holds that
$L_{1}^{S}(\mathbf{X})$ $\lesssim(n+\log n+|S|\log p)$
$\overset{(i)}{\lesssim}|S|+\frac{1}{4}\bigl{(}m\log n+|S|\log p+\sqrt{m|S|\log p}\bigr{)}$
$\lesssim|S|+\frac{m}{4}\Delta_{S\cap S_{0}}^{2}$
$=\dot{\lambda}^{S},$
where step $(i)$ uses Lemma 20. This completes the derivation of the first inequality.
2. The second inequality.
We aim to show $\dot{\lambda}^{S}\lesssim U^{S}(\mathbf{X})$. Recall from (19) that
$U^{S}(\mathbf{X})=\min_{1\leq k\neq l\leq K}\biggl{\{}\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\min_{i\in G_{k}^{\ast}}D^{S}_{kli}(\mathbf{X})\biggr{\}}.$
By Lemma 19, with probability at least $1-C_{7}/n$ uniformly across $S\in\mathcal{S}$ and uniformly over $(k,l,i)$ where $1\leq k\neq l\leq K$ and $i\in G_{k}^{\ast}$, we have
$\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}D^{S}_{kli}(\mathbf{X})\gtrsim\underbrace{\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+|S|}_{\text{Part A}}\underbrace{-\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}R^{S}_{kl}}_{\text{Part B}},$
where
$R^{S}_{kl}=\sqrt{\frac{\xi^{S}}{|G_{l}^{\ast}|}}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}+\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}\sqrt{|S|\xi^{S}}+\frac{1}{|G_{k}^{\ast}|}\xi^{S},$
and $\xi^{S}=\log(n^{2}K^{2}|S|^{2}{\binom{p}{|S|}})$. Under the separation condition for $\mathcal{S}$, recalled in (57), and noting that $|S|\leq\sqrt{p}$ for all $S\in\mathcal{S}$, Part A dominates Part B in the inequality above. Therefore, we have
$\min_{1\leq k\neq l\leq K}\left\{\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}D^{S}_{kli}(\mathbf{X})\right\}$ $\gtrsim\min_{1\leq k\neq l\leq K}\biggl{(}\frac{1}{|G_{k}^{\ast}|}+\frac{1}{|G_{l}^{\ast}|}\biggr{)}^{-1}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}+|S|$
$\gtrsim|S|+\frac{m}{4}\Delta_{S\cap S_{0}}^{2}$
$=\dot{\lambda}^{S}.$
This concludes the derivation of the second inequality and thereby completes the proof. ∎
C.4.8 Proof of Lemma 22
Proof.
For notational simplicity, let $\sigma^{2}=1$. Fix $S\in\mathcal{S}$ and $(l,k)$ with $1\leq l\neq k\leq K$. Condition on the high-probability event in Lemma 19. Then we have
$t^{S}_{lk}(\dot{\lambda}^{S})$ =$\overset{(i)}{=}\sum_{j\in G_{l}^{\ast}}\biggl{\{}\frac{|G_{k}^{\ast}|}{2}D_{lkj}(\mathbf{X},S)-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\dot{\lambda}^{S}\biggr{\}}$
=$\overset{(ii)}{=}\sum_{j\in G_{l}^{\ast}}\biggl{\{}\frac{|G_{k}^{\ast}|}{2}D_{lkj}(\mathbf{X},S)-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\bigl{(}|S|+\frac{m}{4}\Delta_{S\cap S_{0}}^{2}\bigr{)}\biggr{\}}$
$\overset{(iii)}{\gtrsim}\sum_{j\in G_{l}^{\ast}}\biggl{\{}\frac{|G_{k}^{\ast}|}{2}\bigl{(}|S|+\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}\bigr{)}-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\bigl{(}|S|+\frac{m}{4}\Delta_{S\cap S_{0}}^{2}\bigr{)}\biggr{\}}$
$\overset{(iii)}{\geq}\sum_{j\in G_{l}^{\ast}}\biggl{(}\frac{|G_{k}^{\ast}|}{2}|S|+\frac{|G_{k}^{\ast}|}{2}\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}$
$\hskip 45.00006pt-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}|S|-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\frac{m}{4}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}\biggr{)}$
$\overset{(iv)}{\geq}\sum_{j\in G_{l}^{\ast}}\biggl{\{}\biggl{(}\frac{|G_{k}^{\ast}|}{2}-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2|G_{l}^{\ast}|}\biggr{)}|S|+\frac{|G_{k}^{\ast}|}{4}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}\biggr{\}}$
$=\frac{|G_{l}^{\ast}||G_{k}^{\ast}|}{4}\|\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast}\|_{2}^{2}+\frac{|G_{l}^{\ast}||G_{k}^{\ast}|}{2}|S|-\frac{|G_{k}^{\ast}|+|G_{l}^{\ast}|}{2}|S|.$
$\overset{(v)}{\gtrsim}|G_{l}^{\ast}||G_{k}^{\ast}|\bigl{(}\|\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast}\|_{2}^{2}+|S|\bigr{)}$
where step $(i)$ follows from the definition of $t^{S}_{lk}(\lambda^{S})$ in (26), step $(ii)$ follows from the definition of $\dot{\lambda}_{S}$ in Lemma 21, step $(iii)$ uses $-\Delta_{S\cap S_{0}}^{2}>-\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}^{2}$, step $(iv)$ uses
$-m\geq-2\left\{\frac{|G_{k}^{\ast}|\cdot|G_{l}^{\ast}|}{|G_{k}^{\ast}|+|G_{l}^{\ast}|}\right\},$
and step $(v)$ uses $|G_{k}^{\ast}|\geq 2$ and $|G_{l}^{\ast}|\geq 2$. This completes the proof of Lemma 22. ∎
C.4.9 Proof of Lemma 23
the proof proceeds in the following steps.
  1. 1.
    Individual bound by Chi-square tail bound,
  2. 2.
    Uniform bound.
Proof.
For notational simplicity, let $\sigma^{2}=1$.
1. Individual bound by Chi-square tail bound.
Fix $(k,l)$ such that $1\leq k\neq l\leq K$ and $S\in\mathcal{S}$. Then we have
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}I^{S}_{kl}(\mathbf{v},\mathbf{X})$ $=\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}|G_{l}^{\ast}|\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\;\mathcal{E}_{S\cap S_{0},i}\rangle$
$\leq\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}|G_{l}^{\ast}|\biggl{|}\sum_{i\in G_{k}^{\ast}}v_{i}\langle(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\;\mathcal{E}_{S\cap S_{0},i}\rangle\biggr{|}$
$\overset{(i)}{\leq}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}|G_{l}^{\ast}|\biggl{(}\sum_{i\in G_{k}^{\ast}}v_{i}^{2}\biggr{)}^{1/2}\biggl{(}\sum_{i\in G_{k}^{\ast}}\bigl{\langle}(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\mathcal{E}_{S\cap S_{0},i}\bigr{\rangle}^{2}\biggr{)}^{1/2}$
$\leq|G_{l}^{\ast}|\biggl{(}\sum_{i\in G_{k}^{\ast}}\bigl{\langle}(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\mathcal{E}_{S\cap S_{0},i}\bigr{\rangle}^{2}\biggr{)}^{1/2},$
where step $(i)$ uses the Cauchy-Schwarz inequality.
Now we derive an inequality for $\sum_{i\in G_{k}^{\ast}}\bigl{\langle}(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}},\mathcal{E}_{S\cap S_{0},i}\bigr{\rangle}^{2}$. Since
[…]
the square of this quantity follows a chi-square distribution with $|G_{k}^{\ast}|$ degrees of freedom, provided that $S$ is nonempty. Applying the chi-square tail inequality (Lemma 3), if we define
$\mathcal{C}_{Ikl}^{S}:=\biggl{\{}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}I^{S}_{kl}(\mathbf{v},\mathbf{X})\leq\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\bigl{(}|G_{k}^{\ast}|+2\sqrt{|G_{k}^{\ast}|\zeta_{I}^{S}}+2\zeta_{I}^{S}\bigr{)}^{1/2}\biggr{\}},$ (58)
where $\zeta_{I}^{S}=\log(nK^{2}|S|^{2}{\binom{p}{|S|}})$, then we have $\mathbb{P}\bigl{(}\mathcal{C}_{Ikl}^{S}\bigr{)}\geq 1-1/(n^{2}K^{2}|S|^{2}{\binom{p}{|S|}})$.
2. Uniform bound.
Now we show that the event $\mathcal{C}_{Ikl}^{S}$ holds uniformly across all $S\subset\mathcal{S}$ and $1\leq k\neq l\leq K$, with probability at least $1-C_{10}/n$, where $C_{10}>0$ is a constant.
$\mathbb{P}\bigl{(}\bigcup_{1\leq k\neq l\leq K}\bigcup_{S\in\mathcal{S}}(\mathcal{C}_{Ikl}^{S})^{C}\bigr{)}\leq\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{P}((\mathcal{C}_{Ikl}^{S})^{C})\overset{(i)}{\leq}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\frac{1}{n|S|^{2}{\binom{p}{|S|}}}=\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\frac{1}{nq^{2}}\leq\frac{\pi^{2}}{6n},$
where the last step uses $\sum_{q=1}^{\infty}=\pi^{2}/6.$ Since $|S|\leq\sqrt{p}$, $|S|\log p$ dominates $\log|S|$. Therefore, we conclude that uniformly across $S\in\mathcal{S}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{10}/n$, it holds that
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}I^{S}_{kl}(\mathbf{v},\mathbf{X})$
$\quad\quad\lesssim|G_{l}^{\ast}|\|(\boldsymbol{\mu}_{k}^{\ast}-\boldsymbol{\mu}_{l}^{\ast})_{S\cap S_{0}}\|_{2}\biggl{(}|G_{k}^{\ast}|+\sqrt{|G_{k}^{\ast}|\log n}+\sqrt{|G_{k}^{\ast}||S|\log p}+\log n+|S|\log p\biggr{)}^{1/2}.$
This completes the proof of Lemma 23. ∎
C.4.10 Proof of Lemma 24
Proof.
For notational simplicity, let $\sigma^{2}=1$ and for $k=1,\ldots,K$, denote
$\mathrm{IB}_{k}:=\biggl{(}|G_{k}^{\ast}|+\sqrt{|G_{k}^{\ast}|\log n}+\sqrt{|G_{k}^{\ast}||S|\log p}+\log n+|S|\log p\biggr{)}^{1/2}.$
Then we have
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\sum_{1\leq k\neq l\leq K}\frac{T^{S}_{1,kl}(\mathbf{v})}{t^{S}_{lk}(\dot{\lambda}^{S})}$ $\overset{(i)}{\lesssim}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\sum_{1\leq k\neq l\leq K}\frac{I^{S}_{kl}(\mathbf{v},\mathbf{X})\times I^{S}_{lk}(\mathbf{v},\mathbf{X})}{|G_{l}^{\ast}|\cdot|G_{k}^{\ast}|\cdot\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|}$
${\leq}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\sum_{1\leq k\neq l\leq K}\frac{I^{S}_{kl}(\mathbf{v},\mathbf{X})\times I^{S}_{lk}(\mathbf{v},\mathbf{X})}{|G_{l}^{\ast}|\cdot|G_{k}^{\ast}|\cdot\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}}$
$\overset{(ii)}{\lesssim}\sum_{1\leq k\neq l\leq K}\mathrm{IB_{k}^{S}}\times\mathrm{IB_{l}^{S}}$
$\leq\sum_{k=1}^{K}\mathrm{IB_{k}^{S}}\;\sum_{l=1}^{K}\mathrm{IB_{l}^{S}}$
$\overset{(iii)}{\lesssim}n+\sqrt{n\log n}+\sqrt{n|S|\log p}+\log n+|S|\log p$
$\overset{(iv)}{\lesssim}|S|+m\log n+\sqrt{m|S|\log p}+|S|\log p$
$\overset{(v)}{\lesssim}|S|+\frac{m}{4}\Delta_{S\cap S_{0}}^{2}=\dot{\lambda}^{S},$
where step $(i)$ uses Lemma 22 and definition $T^{S}_{1,kl}(\mathbf{v})=I^{S}_{kl}(\mathbf{v},\mathbf{X})\times I^{S}_{lk}(\mathbf{v},\mathbf{X})$ presented in (23), step $(ii)$ uses Lemma 23, step $(iii)$ uses the identity $\sum_{k=1}^{K}|G_{k}^{\ast}|=n$ and inequality $\sum_{k=1}^{K}\sqrt{|G_{k}^{\ast}|}\leq\sqrt{K\sum_{k=1}^{K}|G_{k}^{\ast}|}\lesssim\sqrt{n}$, step $(iv)$ uses the condition $m\gtrsim n/\log n$ of Theorem 1, and step $(v)$ uses the separation condition of $\mathcal{S}$, presented in (6). This completes the proof of Lemma 24. ∎
C.4.11 Proof of Lemma 25
With appropriate algebraic manipulation, the problem reduces to an application of the concentration inequality for an unbounded empirical process developed by Adamczak, (2008) (Lemma 8). This requires estimating the expectation, variance, and sub-exponential norm of the empirical process. Accordingly, the proof proceeds in the following steps.
  1. 1.
    Reformulation into a quadratic form,
  2. 2.
    Bounding the operator norm dependent terms,
  3. 3.
    Bounding the expectation of the supremum,
  4. 4.
    Individual bound using Hanson-wright inequality,
  5. 5.
    Uniform bound.
Proof.
For notational simplicity, we set $\sigma^{2}=1$ and assume, without loss of generality, that $G_{k}^{\ast}={1,\dots,|G_{k}^{\ast}|}$ throughout the proof.
1. Transformation into a quadratic form.
Fix $(k,l)$ such that $1\leq k\neq l\leq K$ and $S\subset[p]$ such that $|S|\leq\sqrt{p}$. We can express $G^{S}_{3,kl}(\mathbf{v})$ as a scaled quadratic form:
$G^{S}_{1,kl}(\mathbf{v})$ $=\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}v_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle$ (59)
$=\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}\frac{1}{2}(v_{i}+v_{j})\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle$
$=\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\underbrace{\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})^{\top}(\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|})\;\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})}_{\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|},\mathbf{X})},$
Here, the $(i,j)$th element of the matrix $\mathbf{A}^{\mathbf{v}}\in\mathbb{R}^{|G_{k}^{\ast}|\times|G_{k}^{\ast}|}$ is defined as $A^{\mathbf{v}}_{i,j}:=(v_{i}+v_{j})/2$, and $\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})$ denotes the $|S|\cdot|G_{k}^{\ast}|$-dimensional vector obtained by stacking the columns of $\mathcal{E}_{S,G_{k}^{\ast}}$ in order, with the first column at the top and the last column at the bottom, and $\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|}$ denotes the tensor product:
$\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|}=\begin{bmatrix}A^{\mathbf{v}}_{1,1}\mathbf{I}_{|S|}&\cdots&A^{\mathbf{v}}_{1,|G_{k}^{\ast}|}\mathbf{I}_{|S|}\\ \vdots&\ddots&\vdots\\ A^{\mathbf{v}}_{|G_{k}^{\ast}|,1}\mathbf{I}_{|S|}&\cdots&A^{\mathbf{v}}_{|G_{k}^{\ast}|,|G_{k}^{\ast}|}\mathbf{I}_{|S|}\end{bmatrix}.$
The quadratic form $\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|},\mathbf{X})$ is centered because
$\mathbb{E}\bigl{[}\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|},\mathbf{X})\bigr{]}=\mathbb{E}\sum_{\{(i,j)\in G_{k}^{\ast}\}}v_{i}\mathbb{E}\bigl{[}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle\bigr{]}=\sum_{i\in G_{k}^{\ast}}v_{i}\mathbb{E}\bigl{[}\|\mathcal{E}_{S,i}\|_{2}^{2}\bigr{]}\overset{(i)}{=}|S|\sum_{i\in G_{k}^{\ast}}v_{i}=0,$
where step $(i)$ uses $\sum_{i=1}^{n}v_{i}=0$, shown in (43), and indexed by
$\mathcal{M}^{S}_{kl}=\{\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|}\mid\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1\}.$
To derive the uniform bound, we apply the uniform Hanson-Wright inequality (Lemma 11) to obtain the concentration inequality for
$\sup_{\mathbf{M}\in\mathcal{M}^{S}_{kl}}\mathrm{QF}_{kl}^{S}(\mathbf{M},\mathbf{X})-\mathbb{E}\biggl{[}\sup_{\mathbf{M}\in\mathcal{M}^{S}_{kl}}\mathrm{QF}_{kl}^{S}(\mathbf{M},\mathbf{X})\biggr{]}.$
To apply Lemma 11, we need to bound the following two quantities:
$\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|\mathbf{M}\|_{\mathrm{op}}\quad\text{and}\quad\|\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{\mathcal{A}^{S}_{kl}}$ $:=\mathbb{E}\bigl{[}\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|(\mathbf{M}+\mathbf{M}^{\top})\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{2}\bigr{]},$
where bounding both terms critically depends on $\|\mathbf{A}^{\mathbf{v}}\|_{\mathrm{op}}$, and the expectation of supremum
$\mathbb{E}\biggl{[}\sup_{\mathbf{M}\in\mathcal{M}^{S}_{kl}}\mathrm{QF}_{kl}^{S}(\mathbf{M},\mathbf{X})\biggr{]}.$
We present the bounds in order.
2. Bounding the operator norm dependent terms.
We start with $\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|\mathbf{M}\|_{\mathrm{op}}$. First note that for $\mathbf{M}\in\mathcal{A}^{S}_{kl}$, its operator norm can be simplified into:
$\|\mathbf{M}\|_{\mathrm{op}}=\|\mathbf{A}^{\mathbf{v}}\otimes\mathbf{I}_{|S|}\|_{\text{op}}=\|\mathbf{A}^{\mathbf{v}}\|_{\text{op}}\|\mathbf{I}_{|S|}\|_{\text{op}}=\|\mathbf{A}^{\mathbf{v}}\|_{\text{op}}.$
Therefore it suffices to bound the supremum of $\|\mathbf{A}^{\mathbf{v}}\|_{\text{op}}$:
$\|\mathbf{A}^{\mathbf{v}}\|_{\text{op}}$ $=\max_{\|\mathbf{u}\|_{2}=1,\mathbf{u}\in\mathbb{R}^{|G_{k}^{\ast}|}}\mathbf{u}^{T}\mathbf{A}^{\mathbf{v}}\mathbf{u}$
$=\max_{\|\mathbf{u}\|_{2}=1,\mathbf{u}\in\mathbb{R}^{|G_{k}^{\ast}|}}\sum_{\{(i,j)\in G_{k}^{\ast}\}}u_{i}u_{j}\frac{v_{i}+v_{j}}{2}$
$=\max_{\|\mathbf{u}\|_{2}=1,\mathbf{u}\in\mathbb{R}^{|G_{k}^{\ast}|}}\bigl{(}\sum_{i=1}^{|G_{k}^{\ast}|}u_{i}v_{i}\bigr{)}\bigl{(}\sum_{j=1}^{|G_{k}^{\ast}|}u_{j}\bigr{)}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\max_{\|\mathbf{u}\|_{2}=1,\mathbf{u}\in\mathbb{R}^{|G_{k}^{\ast}|}}\bigl{(}\sum_{i=1}^{|G_{k}^{\ast}|}u_{i}^{2}\bigr{)}^{1/2}\bigl{(}\sum_{i=1}^{|G_{k}^{\ast}|}v_{i}^{2}\bigr{)}^{1/2}\bigl{(}\sum_{j=1}^{|G_{k}^{\ast}|}u_{j}^{2}\bigr{)}^{1/2}\sqrt{|G_{k}^{\ast}|}$
$\overset{(ii)}{\leq}\sqrt{|G_{k}^{\ast}|},$
where step $(i)$ uses Cauchy-Schwarz inequality, and step $(ii)$ uses $\|\mathbf{v}\|_{2}\leq 1$ for $\mathbf{v}\in\mathbb{V}_{k}$. Since the last term do not depend on $\mathbf{v}$, we have
$\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|\mathbf{M}\|_{\mathrm{op}}\leq\sqrt{|G_{k}^{\ast}|}.$ (60)
Next, we bound $\|\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|^{2}_{\mathcal{A}^{S}_{kl}}$:
$\|\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{\mathcal{A}^{S}_{kl}}^{2}$ $=\bigl{(}\mathbb{E}\bigl{[}\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|(\mathbf{M}+\mathbf{M}^{\top})\;\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{2}\bigr{]}\bigr{)}^{2}$
$\leq\bigl{(}\mathbb{E}\bigl{[}\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|(\mathbf{M}+\mathbf{M}^{\top})\|_{\mathrm{op}}\|\;\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{2}\bigr{]}\bigr{)}^{2}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\bigl{(}\mathbb{E}\bigl{[}2|G_{k}^{\ast}|^{1/2}\|\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{2}\bigr{]}\bigr{)}^{2}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}4|G_{k}^{\ast}|\;\mathbb{E}\bigl{[}\|\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{2}^{2}\bigr{]}$
$=4|G_{k}^{\ast}|\sum_{i\in G_{k}^{\ast}}\mathbb{E}\bigl{[}\|\mathcal{E}_{S,i}\|_{2}^{2}\bigr{]}$
$=4|G_{k}^{\ast}|^{2}|S|,$
where step $(i)$ uses (60), step $(ii)$ applies Jensen’s inequality to a convex function $x\mapsto x^{2}$.
3. Bounding the expectation of the supremum.
It is more convenient to start from (59) and decompose the quadratic form $\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}},\mathbf{X})$ into the following two terms:
$\mathrm{QF}_{kl}^{S}(\mathbf{A}^{\mathbf{v}},\mathbf{X})=\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}v_{j}\bigl{\langle}\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr{\rangle}+\sum_{i\in G_{k}^{*}}v_{i}\|\mathcal{E}_{S,i}\|_{2}^{2}.$
For these two terms, the index set $\mathcal{A}^{S}_{kl}$, which indexes the quadratic form, is more suitably represented by
$\mathbb{V}_{k}:=\left\{\mathbf{v}_{G_{k}^{\ast}}:\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1\right\}.$
We provide the bound for each of these two terms, starting with the first one. First note that given $\mathbf{v}\in\mathbb{V}_{k}$, we have
$\sup_{\mathbf{v}\in\mathbb{V}_{k}}\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}v_{j}\bigl{\langle}\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr{\rangle}$ $=\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{|}\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}v_{j}\bigl{\langle}\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr{\rangle}\bigr{|}$
$=\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{|}\sum_{j\in G_{k}^{\ast}}v_{j}\bigl{\langle}\mathcal{E}_{S,j},\sum_{i\in G_{k}^{\ast},i\neq j}\mathcal{E}_{S,i}\bigr{\rangle}\bigr{|}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{(}\sum_{j\in G_{k}^{\ast}}v_{j}^{2}\bigr{)}^{1/2}\bigl{(}\sum_{j\in G_{k}^{\ast}}\bigl{\langle}\mathcal{E}_{S,j},\sum_{i\in G_{k}^{\ast},i\neq j}\mathcal{E}_{S,i}\bigr{\rangle}^{2}\bigr{)}^{1/2}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\bigl{(}\sum_{j\in G_{k}^{\ast}}\bigl{\langle}\mathcal{E}_{S,j},\sum_{i\in G_{k}^{\ast},i\neq j}\mathcal{E}_{S,i}\bigr{\rangle}^{2}\bigr{)}^{1/2},$
where step $(i)$ uses Cauchy-Schwarz inequality and step $(ii)$ uses $\|\mathbf{v}\|_{2}\leq 1$ for $\mathbf{v}\in\mathbb{V}_{k}$. Using this bound, we can upper bound the expectation as follows:
$\mathbb{E}\biggl{[}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\sum_{\{(i,j)\in G_{k}^{\ast}:i\neq j\}}v_{j}\bigl{\langle}\mathcal{E}_{S,i},\mathcal{E}_{S,j}\bigr{\rangle}\biggr{]}$ $\leq\mathbb{E}\biggl{[}\bigl{(}\sum_{j\in G_{k}^{\ast}}\bigl{\langle}\mathcal{E}_{S,j},\sum_{i\in G_{k}^{\ast},i\neq j}\mathcal{E}_{S,i}\bigr{\rangle}^{2}\bigr{)}^{1/2}\biggr{]}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\biggl{(}\sum_{j\in G_{k}^{\ast}}\mathbb{E}\bigl{[}\bigl{\langle}(\boldsymbol{\varepsilon}_{j})_{S},\sum_{i\neq j,i\in[n_{k}]}\mathcal{E}_{S,i}\bigr{\rangle}^{2}\bigr{]}\biggr{)}^{1/2}$
=$\stackrel{{\scriptstyle(ii)}}{{=}}\bigl{(}\sum_{j\in G_{k}^{\ast}}\mathbb{E}\|\hskip-6.99997pt\sum_{i\in G_{k}^{\ast},i\neq j}\hskip-6.99997pt\mathcal{E}_{S,i}\|_{2}^{2}\;\bigr{)}^{1/2}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}\bigl{(}|G_{k}^{\ast}|(|G_{k}^{\ast}|-1)|S|\bigr{)}^{1/2}$
$\leq|G_{k}^{\ast}|\sqrt{|S|},$ (61)
where step $(i)$ uses Jensen’s inequality to a concave function $x\mapsto x^{1/2}$, step $(ii)$ the independence and zero-mean assumption, and step $(iii)$ uses isotropic Gaussian assumption.
Next, we bound the expectation of the supremum of $\sum_{i\in G_{k}^{*}}v_{i}\|\mathcal{E}_{S,i}\|_{2}^{2}$.
$\mathbb{E}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{|}\sum_{i\in G_{k}^{*}}v_{i}\|\mathcal{E}_{S,i}\|_{2}^{2}\bigr{|}$ =$\overset{(i)}{=}\mathbb{E}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{|}\sum_{i\in G_{k}^{*}}v_{i}\bigr{(}\|\mathcal{E}_{S,i}\|_{2}^{2}-|S|\bigl{)}\bigr{|}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\mathbb{E}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\bigl{(}\sum_{i\in G_{k}^{*}}v_{i}^{2}\bigr{)}^{1/2}\biggl{\{}\sum_{i\in G_{k}^{*}}\left(\|\mathcal{E}_{S,i}\|_{2}^{2}-|S|\right)^{2}\biggr{\}}^{1/2}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}\mathbb{E}\biggl{\{}\sum_{i\in G_{k}^{*}}\left(\|\mathcal{E}_{S,i}\|_{2}^{2}-|S|\right)^{2}\biggr{\}}^{1/2}$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}\biggl{\{}\sum_{i\in G_{k}^{*}}\mathbb{E}\left(\|\mathcal{E}_{S,i}\|_{2}^{2}-|S|\right)^{2}\biggr{\}}^{1/2}$
$=\bigl{(}\sum_{i\in G_{k}^{*}}\mathrm{Var}\|\mathcal{E}_{S,i}\|_{2}^{2}\bigr{)}^{1/2}$
=$\overset{(v)}{=}\sqrt{2|G_{k}^{\ast}||S|},$ (62)
where step $(i)$ uses $\sum_{i\in G_{k}^{\ast}}v_{i}=0$, which follows from the fact that $\Gamma_{k}$ is orthogonal to $\mathbbm{1}_{G_{k}^{\ast}}$, step $(ii)$ uses Cauchy-Schwarz inequality, step $(iii)$ uses $\|\mathbf{v}\|_{2}\leq 1$, step $(iv)$ uses Jensen’s inequality to a concave function $x\mapsto x^{1/2}$, and step $(v)$ uses our Gaussian noise assumption and the variance formula for Chi-square random variables.
Combining (61) and (62), we conclude that
$\mathbb{E}\biggl{[}\sup_{\mathbf{M}\in\mathcal{M}^{S}_{kl}}\mathrm{QF}_{kl}^{S}(\mathbf{M},\mathbf{X})\biggr{]}\lesssim|G_{k}^{\ast}|\sqrt{|S|}.$
4. Individual bound using Hanson-wright inequality.
Applying Lemma 11 with the bounds established in the previous steps and using the formulation (59) of $G^{S}_{1,kl}(\mathbf{v})$, we define an event
$\mathcal{C}_{N,kl}^{S}:=\biggl{\{}\sup_{\mathbf{v}\in\mathbb{V}_{k}}\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\sum_{\{(i,j)\in G_{k}^{\ast}\}}v_{i}\left\langle\mathcal{E}_{S,i},\mathcal{E}_{S,j}\right\rangle\lesssim\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\underbrace{|G_{k}^{\ast}|\sqrt{|S|}}_{\mathbb{E}\bigl{[}\sup_{\mathbf{M}\in\mathcal{M}^{S}_{kl}}\mathrm{QF}_{kl}^{S}(\mathbf{M},\mathbf{X})\bigr{]}\lesssim}~{}+~{}\frac{|G_{l}^{\ast}|}{|G_{k}^{\ast}|}\zeta^{S}_{N,kl}\biggr{\}},$
where $\zeta^{S}_{N,kl}:=|G_{k}^{\ast}|\sqrt{|S|(\log n+2\log|S|+\log{{\binom{p}{|S|}}})}$, then we have
$\mathbb{P}\biggl{(}(\mathcal{C}_{N,kl}^{S})^{C}\biggr{)}$ $\lesssim\exp\biggl{[}-\min\biggl{(}\underbrace{\frac{(\zeta^{S}_{N,kl})^{2}}{|G_{k}^{\ast}|^{2}|S|}}_{\|\mathrm{vec}(\mathcal{E}_{S,G_{k}^{\ast}})\|_{\mathcal{A}^{S}_{kl}}^{2}\lesssim},\quad\underbrace{\frac{\zeta^{S}_{N,kl}}{\sqrt{|G_{k}^{\ast}|}}}_{\sup_{\mathbf{M}\in\mathcal{A}^{S}_{kl}}\|\mathbf{M}\|_{\mathrm{op}}\lesssim}\biggr{)}\biggr{]}$
$\lesssim\frac{1}{n|S|^{2}{\binom{p}{|S|}}}.$
The proof is complete once we show that the intersection event
$\bigcap_{S\in[p],|S|\leq\sqrt{p}}\bigcap_{1\leq k\neq l\leq K}\mathcal{C}_{N,kl}^{S}$
holds with high probability.
5. Uniform bound.
Now we show that the event $\mathcal{C}_{N,kl}^{S}$ holds uniformly across all $S\in[p]$ such that $|S|\leq\sqrt{p}$ and $1\leq k\neq l\leq K$, with probability at least $1-C_{11}/n$, where $C_{11}>0$ is a constant.
$\mathbb{P}\bigl{(}\bigcup_{S\in[p],|S|\leq\sqrt{p}}\bigcup_{1\leq k\neq l\leq K}(\mathcal{C}_{N,kl}^{S})^{C}\bigr{)}$ $\leq\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{P}((\mathcal{C}_{N,kl}^{S})^{C})$
$\lesssim\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\frac{K^{2}}{n|S|^{2}{\binom{p}{|S|}}}$
$=\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\frac{K^{2}}{nq^{2}}$
$\leq\frac{K^{2}\pi^{2}}{6n},$
where the last step uses $\sum_{q=1}^{\infty}=\pi^{2}/6.$ Since $|S|\leq\sqrt{p}$, $|S|\log p$ dominates $\log|S|$. Also, we treat $K$ as a constant. Therefore, we conclude we conclude that uniformly across $S\in\mathcal{S}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{11}/n$, it holds that
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}G^{S}_{1,kl}(\mathbf{v})\lesssim|G_{l}^{\ast}|\sqrt{|S|(\log n+|S|\log p)}$
This completes the proof of Lemma 25.
C.4.12 Proof of Lemma 26
The proof proceeds in the following steps.
  1. 1.
    Conditional analysis using a concentration inequality for the supremum of a Gaussian process,
  2. 2.
    Individual bound by Chi-square tail bound,
  3. 3.
    Uniform bound
Proof.
For notational simplicity, let $\sigma^{2}=1$.
1. Conditional analysis.
Since the sum of inner products of Gaussian random vectors is challenging to analyze directly, we first condition on ¯$\overline{\mathcal{E}}_{S,l}^{\ast}$ and then apply the concentration inequality for the supremum of a Gaussian process (Lemma 9). Fix $(k,l)$ such that $1\leq k\neq l\leq K$ and $S\subset[p]$ such that $|S|\leq\sqrt{p}$. Conditioned on ¯$\overline{\mathcal{E}}_{S,l}^{\ast}$, $G^{S}_{2,kl}(\mathbf{v})$ can be viewed as a canonical Gaussian process $\langle G,\mathbf{w}\rangle$, where $G\sim\mathcal{N}(0,\mathbf{I}_{|G_{k}^{\ast}|})$ and $\mathbf{w}\in\mathbb{W}_{k}$, with
$\mathbb{W}_{k}:=\left\{\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}\mathbf{v}_{\upharpoonright G_{k}^{\ast}}:\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1\right\},$
where $\mathbf{v}_{\upharpoonright G_{k}^{*}}$ is the restriction of $\mathbf{v}\in\Gamma_{K}$ onto $G_{k}^{\ast}$, and the canonical metric on $\mathbb{W}_{k}$ is the $\ell_{2}$-metric.
To apply Lemma 9, we have to bound
$\mathrm{EC}_{kl}^{S}(\overline{\mathcal{E}}_{S,l}^{\ast})$ $:=\mathbb{E}\bigl{[}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\sum_{i\in G_{k}^{*}}v_{i}\bigl{\langle}\overline{\mathcal{E}}_{S,l}^{\ast},\mathcal{E}_{S,i}\bigr{\rangle}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast}\bigr{]}=\mathbb{E}\bigl{[}\sup_{\mathbf{w}\in\mathbb{W}_{k}}\langle G,\mathbf{w}\rangle\;|\;\overline{\mathcal{E}}_{S,l}^{\ast}\bigr{]}~{}\text{, and}$
$\mathrm{SVC}_{kl}^{S}(\overline{\mathcal{E}}_{S,l}^{\ast})$ $:=\sup_{\mathbf{v}\in\Gamma_{k},\|\mathbf{v}\|_{2}=1}\mathrm{Var}\bigl{(}\sum_{i\in G_{k}^{*}}v_{i}\bigl{\langle}\overline{\mathcal{E}}_{S,l}^{\ast},\mathcal{E}_{S,i}\bigr{\rangle}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast}\bigr{)}=\sup_{\mathbf{w}\in\mathbb{W}_{k}}\mathrm{Var}\bigl{(}\langle G,\mathbf{w}\rangle\;|\;\overline{\mathcal{E}}_{S,l}^{\ast}\bigr{)}.$
We bound $\mathrm{EC}_{kl}^{S}(\overline{\mathcal{E}}_{S,l}^{\ast})$ using Dudley’s integral inequality (Lemma 10):
[…]
where step $(i)$ uses the fact that the $\varepsilon$-covering entropy of the unit sphere in $\mathbb{R}^{|G_{k}^{\ast}|}$ is at most $|G_{k}^{\ast}|\log(1/\varepsilon)$, up to a constant factor, for any $\varepsilon\in(0,1)$.
Next, we bound $\mathrm{SVC}_{kl}^{S}(\overline{\mathcal{E}}_{S,l}^{\ast})$:
$\mathrm{SVC}_{kl}^{S}(\overline{\mathcal{E}}_{S,l}^{\ast})=\sup_{\mathbf{w}\in\mathbb{W}_{k}}\mathrm{Var}\bigl{(}\langle G,\mathbf{w}\rangle\;|\;\overline{\mathcal{E}}_{S,l}^{\ast}\bigr{)}=\sup_{\mathbf{w}\in\mathbb{W}_{k}}\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}\sum_{i\in G_{k}^{\ast}}v_{i}^{2}\leq\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2},$
where the last inequality follows from $\|\mathbf{v}\|_{2}=1$ for $\mathbf{v}\in\Gamma_{K}$. Let us denote $\xi^{S}=\log(n|S|^{2}{\binom{p}{|S|}})$. Define the following conditional probability as follows:
$\mathrm{CP}_{N,kl}^{S}:=\mathbb{P}\biggl{(}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}|G_{2,kl}^{S}(\mathbf{v})|\gtrsim|G_{l}^{\ast}|\underbrace{\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}\sqrt{|G_{k}^{\ast}|}}_{\gtrsim\mathrm{EC}_{kl}^{S}(\overline{\mathcal{E}}_{S,l}^{\ast})}+|G_{l}^{\ast}|\underbrace{\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}}_{\gtrsim\sqrt{\mathrm{SVC}_{kl}^{S}(\overline{\mathcal{E}}_{S,l}^{\ast}})}\sqrt{\xi^{S}}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast}\biggr{)},$
Then by Lemma 9, we have
$\mathrm{CP}_{N,kl}^{S}\leq\frac{1}{n|S|^{2}{\binom{p}{|S|}}}.$ (63)
2. Chi-square tail bound
Now we use the Chi-square tail inequality to further bound ¯$\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}$ and ¯$\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}$. If we define the following event:
$\mathcal{B}_{N,kl}^{S}:=\left\{\|\overline{\mathcal{E}}_{S,l}^{\ast}\|_{2}^{2}\leq\frac{1}{|G_{l}^{\ast}|}\left(|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\right)\right\}.$
By Lemma 3 , we have
$\mathbb{P}\bigl{(}(\mathcal{B}_{N,kl}^{S})^{C}\bigr{)}\leq\frac{1}{n|S|^{2}{\binom{p}{|S|}}}.$
Combining the bounds in $\mathrm{CP}_{N,kl}^{S}$ and $\mathcal{B}_{N,kl}^{S}$, we define an event:
$\mathcal{A}_{N,kl}^{S}:=\biggl{\{}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}|G_{2,kl}^{S}(\mathbf{v})|\lesssim\sqrt{|G_{l}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}(\sqrt{|G_{k}^{\ast}|+\sqrt{\xi^{S}})}\biggr{\}}.$
The proof is complete once we show that the intersection event
$\bigcap_{1\leq k\neq l\leq K}\bigcap_{S\in[p],|S|\leq\sqrt{p}}\mathcal{A}^{S}_{N,kl}$
holds with high probability.
3. Uniform bound.
We have
$\mathbb{P}\biggl{(}\bigcup_{1\leq k\neq l\leq K}$ $\bigcup_{S\in[p],|S|\leq\sqrt{p}}{(\mathcal{A}^{S}_{N,kl})^{C}}\biggr{)}$
$\leq\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{P}{\bigl{(}(\mathcal{A}^{S}_{N,kl})^{C}\bigr{)}}$
$=\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{E}\bigl{[}\mathbb{P}\bigl{(}(\mathcal{A}^{S}_{N,kl})^{C}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast}\bigr{)}\bigr{]}$
$\leq\sum_{1\leq k\neq l\leq K}\sum_{q=1}^{\lfloor\sqrt{p}\rfloor}\sum_{|S|=q}\mathbb{E}\bigl{[}\mathbb{P}\bigl{(}(\mathcal{A}^{S}_{N,kl})^{C}\;|\;\overline{\mathcal{E}}_{S,l}^{\ast}\bigr{)}\mathbbm{1}(\mathcal{B}^{S}_{N,kl})+\mathbbm{1}\bigl{(}(\mathcal{B}^{S}_{N,kl})^{C}\bigr{)}\bigr{]}$
[…]
$\overset{(ii)}{\lesssim}\frac{K^{2}\pi^{2}}{6n}+\frac{K^{2}\pi^{2}}{6n},$
where step $(i)$ uses $\sum_{q=1}^{\infty}(1/q^{2})=\pi^{2}/6$, and step $(ii)$ also uses it and additionally uses (63).
Treating $K$ as a constant, we conclude that uniformly across $S\in\mathcal{S}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{12}/n$, where $C_{12}>0$ is a constant, it holds that
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}|G_{2,kl}^{S}(\mathbf{v})|\lesssim\sqrt{|G_{l}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}\sqrt{|G_{k}^{\ast}|+\sqrt{\xi^{S}}}.$
This completes the proof of Lemma 26. ∎
C.4.13 Proof of Lemma 27
Leveraging the decomposition of $T_{2,kl}^{S}$ into $G_{1,kl}^{S}$ and $G_{2,kl}^{S}$ given in Appendix C.3.3, the proof proceeds in the following intermediate steps:
  1. 1.
    Problem decomposition
  2. 2.
    Inequality for the $G_{1}\times G_{1}$ term
  3. 3.
    Inequality for the $G_{2}\times G_{2}$ term
  4. 4.
    Conclusion
Proof.
For notational simplicity, let $\sigma^{2}=1$.
1. Problem decomposition.
Recall that we defined $\Delta_{S\cap S_{0}}^{2}=\min_{1\leq k\neq l\leq K}\|(\boldsymbol{\mu}_{l}^{\ast}-\boldsymbol{\mu}_{k}^{\ast})_{S\cap S_{0}}\|_{2}^{2}$ in Theorem 1, and $\dot{\lambda}^{S}=|S|+m\Delta_{S\cap S_{0}}^{2}/4$ in Lemma 21.
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\sum_{1\leq k\neq l\leq K}\frac{T^{S}_{2,kl}(\mathbf{v})}{t^{S}_{lk}(\dot{\lambda}^{S})}$
=$\overset{(i)}{=}\sum_{1\leq k\neq l\leq K}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\frac{N^{S}_{kl}(\mathbf{v},\mathbf{X})\;N^{S}_{lk}(\mathbf{v},\mathbf{X})}{t^{S}_{lk}(\dot{\lambda}^{S})}$
$\overset{(ii)}{\lesssim}\sum_{1\leq k\neq l\leq K}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\frac{N^{S}_{kl}(\mathbf{v},\mathbf{X})\;N^{S}_{lk}(\mathbf{v},\mathbf{X})}{|G_{l}^{\ast}||G_{k}^{\ast}|\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$=\sum_{1\leq k\neq l\leq K}\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\frac{\bigl{(}G_{1,kl}^{S}(\mathbf{v})+G_{2,kl}^{S}(\mathbf{v})\bigr{)}\bigl{(}G_{1,lk}^{S}(\mathbf{v})+G_{2,lk}^{S}(\mathbf{v})\bigr{)}}{|G_{l}^{\ast}||G_{k}^{\ast}|\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
where step $(i)$ uses the definition of $T^{S}_{2,kl}(\mathbf{v})$ stated in (24) and step $(ii)$ uses Lemma 22. Expanding the numerator yields four terms. We establish the inequality for two of them, the $G_{1}\times G_{1}$ and $G_{2}\times G_{2}$ terms, as the cross terms can be easily handled by leveraging the results from these two cases. Before proceeding, we recall from (6) that $\mathcal{S}$ is defined as
$\mathcal{S}=\left\{S:S\subset[p],|S|\leq\sqrt{p},\Delta_{S\cap S_{0}}^{2}\gtrsim\left(\log n+\frac{|S|\log p}{m}+\sqrt{\frac{|S|\log p}{m}}\right)\right\}.$
2. Inequality for the $G_{1}\times G_{1}$ term.
By Lemma 25, Uniformly across $S\in\mathcal{S}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{11}/n$, it holds that
$\sum_{1\leq k\neq l\leq K}$ $\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\frac{G_{1,kl}^{S}(\mathbf{v})G_{1,lk}^{S}(\mathbf{v})}{|G_{l}^{\ast}||G_{k}^{\ast}|\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$\overset{(i)}{\lesssim}\sum_{1\leq k\neq l\leq K}\frac{|S|(\log n+|S|\log p)}{\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$\lesssim\frac{|S|\log n+|S|^{2}\log p}{\bigl{(}\Delta_{S\cap S_{0}}^{2}+|S|\bigr{)}}.$
Therefore it suffices to show that
$|S|\log n+|S|^{2}\log p\lesssim\bigl{(}|S|+m\Delta_{S\cap S_{0}}^{2}\bigr{)}\bigl{(}\Delta_{S\cap S_{0}}^{2}+|S|\bigr{)},$
which is true since $|S|\Delta_{S\cap S_{0}}^{2}\gtrsim|S|\log n$ and $m|S|\Delta_{S\cap S_{0}}^{2}\gtrsim|S|^{2}\log p$ under the separtion condition of $\mathcal{S}$.
3. Inequality for the $G_{2}\times G_{2}$ term.
By Lemma 26, Uniformly across $S\in\mathcal{S}$ and $(k,l)$ such that $1\leq k\neq l\leq K$, with probability at least $1-C_{12}/n$, it holds that
$\sum_{1\leq k\neq l\leq K}$ $\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\frac{G_{2,kl}^{S}(\mathbf{v})G_{2,lk}^{S}(\mathbf{v})}{|G_{l}^{\ast}||G_{k}^{\ast}|\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$\overset{(i)}{\lesssim}\sum_{1\leq k\neq l\leq K}\biggl{\{}\frac{\sqrt{|G_{l}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}}{|G_{l}^{\ast}||G_{k}^{\ast}|\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}\sqrt{|G_{k}^{\ast}|+\sqrt{\xi^{S}}}}$
×$\hskip 60.00009pt\times\sqrt{|G_{k}^{\ast}|}\sqrt{|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}}\sqrt{|G_{l}^{\ast}|+\sqrt{\xi^{S}}}\biggr{\}}$
$=\sum_{1\leq k\neq l\leq K}\frac{\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}\sqrt{|G_{k}^{\ast}||G_{l}^{\ast}|+\sqrt{\xi^{S}}\bigl{(}|G_{k}^{\ast}|+|G_{l}^{\ast}|\bigr{)}+\xi^{S}}}{\sqrt{|G_{l}^{\ast}||G_{k}^{\ast}|}\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$=\sum_{1\leq k\neq l\leq K}\frac{\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}\bigl{(}\sqrt{|G_{k}^{\ast}||G_{l}^{\ast}|}+\sqrt{\sqrt{\xi^{S}}\bigl{(}|G_{k}^{\ast}|+|G_{l}^{\ast}|\bigr{)}}+\sqrt{\xi^{S}}\bigr{)}}{\sqrt{|G_{l}^{\ast}||G_{k}^{\ast}|}\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}},$
where $\xi^{S}=\log(n|S|^{2}{\binom{p}{|S|}})$. The last term can be decomposed into three parts, and we establish an inequality for each one in turn, starting with the first term:
$\sum_{1\leq k\neq l\leq K}$ $\frac{\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}\bigl{(}\sqrt{|G_{k}^{\ast}||G_{l}^{\ast}|}\bigr{)}}{\sqrt{|G_{l}^{\ast}||G_{k}^{\ast}|}\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$=\sum_{1\leq k\neq l\leq K}\frac{\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}}{\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$\lesssim\frac{\bigl{(}|S|+|S|\sqrt{\log p}+\sqrt{|S|\log n}+\log n+|S|\log p\bigr{)}}{\Delta_{S\cap S_{0}}^{2}+|S|}$
$\lesssim|S|+m\Delta_{S\cap S_{0}}^{2},$
where the last inequality holds due to the sepration condition of $\mathcal{S}$.
Now, for the second term, we first recall from Theroem 1 that
$m=2\min_{1\leq k\neq l\leq K}\left\{\frac{|G_{k}^{\ast}||G_{l}^{\ast}|}{|G_{k}^{\ast}|+|G_{l}^{\ast}|}\right\}.$
Then we have:
$\sum_{1\leq k\neq l\leq K}$ $\frac{\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}\bigl{(}\sqrt{\sqrt{\xi^{S}}\bigl{(}|G_{k}^{\ast}|+|G_{l}^{\ast}|\bigr{)}}\bigr{)}}{\sqrt{|G_{l}^{\ast}||G_{k}^{\ast}|}\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}},$
$\lesssim\frac{\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}(\xi^{S})^{1/4}}{\sqrt{m}\bigl{(}\Delta_{S\cap S_{0}}^{2}+|S|\bigr{)}}$
$\lesssim|S|+m\Delta_{S\cap S_{0}}^{2},$
where the last inequality holds because under the separation condition of $\mathcal{S}$, we have $\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}\lesssim|S|+m\Delta_{S\cap S_{0}}^{2}$ and $(\xi^{S})^{1/4}\lesssim\sqrt{m}\bigl{(}\Delta_{S\cap S_{0}}^{2}+|S|\bigr{)}$.
Finally, for the third term, we have
$\sum_{1\leq k\neq l\leq K}$ $\frac{\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}\sqrt{\xi^{S}}}{\sqrt{|G_{l}^{\ast}||G_{k}^{\ast}|}\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$\lesssim\sum_{1\leq k\neq l\leq K}\frac{\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}\sqrt{\xi^{S}}}{\sqrt{m}\bigl{(}\|(\boldsymbol{\mu}^{\ast}_{k}-\boldsymbol{\mu}^{\ast}_{l})_{S\cap S_{0}}\|_{2}^{2}+|S|\bigr{)}}$
$\lesssim|S|+m\Delta_{S\cap S_{0}}^{2},$
where the last inequality holds because under the separation condition of $\mathcal{S}$, we have $\bigl{(}|S|+2\sqrt{|S|\xi^{S}}+2\xi^{S}\bigr{)}\lesssim|S|+m\Delta_{S\cap S_{0}}^{2}$ and $(\xi^{S})^{1/2}\lesssim\sqrt{m}\bigl{(}\Delta_{S\cap S_{0}}^{2}+|S|\bigr{)}$.
4. Conclusion.
The inequalities for the cross terms can be verified in a straightforward manner. Combining all the inequalities established above, we conclude that, uniformly over $S\in\mathcal{S}$ and all $(k,l)$ with $1\leq k\neq l\leq K$, the following holds with probability at least $1-C_{13}/n$, where $C_{13}>0$ is a constant, it holds that:
$\sup_{\mathbf{v}\in\Gamma_{K},\|\mathbf{v}\|_{2}=1}\sum_{1\leq k\neq l\leq K}\frac{T^{S}_{2,kl}(\mathbf{v})}{t^{S}_{lk}(\dot{\lambda}^{S})}.$
This completes the proof of Lemma 27. ∎
Appendix D Proof of Theorem2
The proof proceeds in the following intermediate steps:
  1. 1.
    Reduction to symmetric two-cluster problem (64),
  2. 2.
    Reduction to average case risk (65),
  3. 3.
    Reduction to Bayes risk (66),
  4. 4.
    Bounding the Bayes risk (68).
Proof.
We proceed by following the steps outlined above in sequence.
1. Reduction to symmetric two-cluster problem.
We denote the separation by $\rho>0$ and leave its value unspecified for now. First, we define a reduced class of distributions consisting of two symmetric, sparse cluster centers.
$\tilde{\Theta}(s,\rho,2):=\tilde{\mathcal{M}}(s,\rho,2)\times\mathcal{H}(2)$
where
~$\tilde{\mathcal{M}}(s,\rho,2):=\bigl{\{}(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}):$ $\boldsymbol{\mu}_{1}=\boldsymbol{\mu},\boldsymbol{\mu}_{2}=-\boldsymbol{\mu},\boldsymbol{\mu}\in\mathbb{R}^{p},\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}^{2}\geq\rho^{2},$
$\mathrm{supp}(\boldsymbol{\mu})=S_{0}\subset[p],|S_{0}|\leq s\bigr{\}},$
Then we have (by augmenting the matrices in $\mathcal{H}(2)$ with zero columns to match the dimensions):
$\tilde{\Theta}(s,\rho,2)\subset\Theta(s,\rho,K).$
This inclusion immediately implies that
$\inf_{\hat{\mathbf{H}}}\sup_{\boldsymbol{\theta}\in\Theta(s,\rho,K)}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H})\geq\inf_{\hat{\mathbf{H}}}\sup_{\boldsymbol{\theta}\in\tilde{\Theta}(s,\rho,2)}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H}),$ (64)
since the supremum over a smaller set is less than or equal to that over a larger set.
2. Reduction to average case risk.
Next, we lower bound the worst-case risk in (64) by the average risk.
$\inf_{\hat{\mathbf{H}}}\sup_{\boldsymbol{\theta}\in\tilde{\Theta}(s,\rho,2)}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H})$ $=\inf_{\hat{\mathbf{H}}}\sup_{(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})\in\tilde{\mathcal{M}}(s,\rho,2)}\sup_{\mathbf{H}\in\mathcal{H}(2)}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H})$
$\geq\inf_{\hat{\mathbf{H}}}\sup_{(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})\in\tilde{\mathcal{M}}(s,\rho,2)}\mathbb{E}_{\mathbf{H}(\mathbf{X})\sim\pi}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}\neq\mathbf{H}),$ (65)
where $\pi$ is a uniform prior on $\mathcal{H}(2)$, formally given by, for $i=1,\ldots,n$,
[…]
3. Reduction to Bayes risk.
The risk (65) is equivalent to the minimax risk under zero-one loss, where the data is generated from a uniform hierarchical Gaussian mixture model defined as follows:
$H_{i,1}$ $\sim\mathrm{Ber}(1/2),$
$\mathbf{X}_{i}\mid H_{i,1}$ $\sim\mathcal{N}\left((-1)^{H_{i,1}}\boldsymbol{\mu}\;,\mathbf{I}_{p}\right).$
In this setting, the algorithm that minimizes $\mathbb{E}_{\mathbf{H}\sim\pi}\mathbb{P}_{\boldsymbol{\theta}}\bigl{(}\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H}\bigr{)}$ is the Bayes classifier $\mathbf{H}^{\boldsymbol{\mu}}$, which assumes knowledge of the true cluster centers. It operates by thresholding the Gaussian likelihood ratio, and under symmetric cluster centers with isotropic covariance, this yields the decision rule $H^{\boldsymbol{\mu}}_{i,1}=\mathbbm{1}\left(\boldsymbol{\mu}^{\top}\mathbf{X}_{i}>0\right)$ for $i=1,\ldots,n$. Its error probability for one observation, known as the Bayes rate, is explicitly given by (Cai and Zhang,, 2019):
$\mathcal{R}^{\ast}_{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}}:=\Phi\biggl{(}-\frac{\|\boldsymbol{\mu}_{1}-\boldsymbol{\mu}_{2}\|_{2}}{2\sigma}\biggr{)},$ (66)
where $\Phi$ denotes the cumulative distribution function of standard normal distribution. Therefore we can bound (65) further from below:
$\inf_{\hat{\mathbf{H}}}\sup_{(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})\in\tilde{\mathcal{M}}(s,\rho,2)}\mathbb{E}_{\mathbf{H}(\mathbf{X})\sim\pi}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}\neq\mathbf{H})$ $~{}\overset{(i)}{\geq}\sup_{(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})\in\tilde{\mathcal{M}}(s,\rho,2)}\inf_{\hat{\mathbf{H}}}\;\mathbb{E}_{\mathbf{H}(\mathbf{X})\sim\pi}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}\neq\mathbf{H})$
=$~{}\overset{(ii)}{=}\sup_{(\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2})\in\tilde{\mathcal{M}}(s,\rho,2)}1-(1-\mathcal{R}^{\ast}_{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}})^{n}$
=$~{}\overset{(iii)}{=}1-\bigl{(}1-\Phi(-\rho/\sigma)\bigr{)}^{n},$ (67)
where step $(i)$ applies the max–min inequality (see, for example, Section 5.4.1 of Boyd and Vandenberghe, 2004), step $(ii)$ follows from the optimality of Bayes classifier, and step $(iii)$ uses (66).
4. Bounding the Bayes risk.
If we set $\rho^{2}=C_{4}\sigma^{2}\log n$, we have
$\Phi(-\rho/\sigma)=\Phi\bigl{(}-\sqrt{C_{4}\log n}\bigr{)}=1-\mathbb{P}(Z>-\sqrt{C_{4}\log n})\overset{(i)}{\geq}1-\exp(-\frac{C_{4}}{2}\log n)=1-n^{-C_{4}/2},$
where $Z$ denotes a standard normal random variable and step $(i)$ uses Gaussian tail bound. Leveraging this, for any $n\geq 2$, we further bound (67) as:
$1-\bigl{\{}1-\Phi\bigl{(}-\sqrt{C_{4}\log n}\bigr{)}\bigr{\}}^{n}\geq 1-n^{-C_{4}n/2}>C_{5},$ (68)
where $C_{5}>0$ depends on $C_{4}$. Collecting the reductions (64), (65), (66), and the bound (68), we conclude that
$\inf_{\hat{\mathbf{H}}}\sup_{\boldsymbol{\theta}\in\Theta(s,\sigma^{2}\log n,K)}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}(\mathbf{X})\neq\mathbf{H})>C_{5}.$
This completes the proof of Theorem 2. ∎
Remark 29.
The reduction in step $(i)$ of (67) is not tight in some regime. A sharper analysis can be obtained by decomposing the average risk $\mathbb{E}_{\mathbf{H}(\mathbf{X})\sim\pi}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}\neq\mathbf{H})$ into the Bayes risk $\mathcal{R}^{\ast}_{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}}$ and the excess risk $\mathbb{E}_{\mathbf{H}(\mathbf{X})\sim\pi}\mathbb{P}_{\boldsymbol{\theta}}(\hat{\mathbf{H}}\neq\mathbf{H})-\mathcal{R}^{\ast}_{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2}}$. Cai et al., (2019) shows that the excess risk is of order $(s\log p)/n$ in a setting slightly different from ours. Rather than modifying their analysis, we assume a standard high-dimensional sparse regime $s\log p/n=o(1)$, so that the Bayes risk term $\log n$ dominates. In this regime, this lower bound matches the separation bound established in Theorem 1.
Appendix E Details for the SDP objective function
This section presents supplementary information for the original SDP objective (Appendix E.1) and the reduced SDP objective in Algorithm 1 (Appendix E.2).
E.1 Derivation of the SDP objective function
This section presents the derivation of the SDP objective (1), which was omitted from Section 2.1 for brevity. The derivation follows Sections 2 and 3 of Zhuang et al., (2023) and Section 1 of Chen and Yang, (2021), and is included here for completeness. We proceed in the following order:
  1. 1.
    From Gaussian mixture likelihood to K-means objective,
  2. 2.
    From K-means objective to SDP objective.
1. From Gaussian mixture likelihood to K-means objective.
We assume that the common covariance $\boldsymbol{\Sigma}^{\ast}$ is known. The log-likelihood for the parameters $G_{1},\ldots,G_{K}$ and $\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{K}$ is
$\ell(G_{1},\ldots,G_{K},\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{K})=-\frac{np}{2}\log(2\pi|\boldsymbol{\Sigma}^{\ast}|)-\frac{1}{2}\sum_{k=1}^{K}\sum_{i\in G_{k}}(\mathbf{X}_{i}-\boldsymbol{\mu}_{k})^{\top}\boldsymbol{\Omega}^{\ast}(\mathbf{X}_{i}-\boldsymbol{\mu}_{k}),$
where $|\boldsymbol{\Sigma}^{\ast}|$ denotes the determinant of $\boldsymbol{\Sigma}^{\ast}$, $G_{1}\cup G_{2}\cup\ldots\cup G_{K}=[n]$ and $G_{k}\cap G_{l}=\emptyset$ for all $1\leq k\neq l\leq K$. Since the cluster center $\boldsymbol{\mu}_{k}$’s are not of direct interest, we substitute each $\boldsymbol{\mu}_{k}$ with its maximum likelihood estimate given the clustering $G_{1},\ldots,G_{K}$:
$\bar{\mathbf{X}}^{k}:=\frac{1}{|G_{k}|}\sum_{j\in G_{k}}\mathbf{X}_{j}.$
By further omitting constants that are irrelevant to the clustering assignment $G_{1},\ldots,G_{K}$, and multiplying the remaining term by 2, we define the scaled profile log-likelihood for the clustering as follows:
$\ell_{\mathrm{lik}}(G_{1},\ldots,G_{K};\mathbf{X}):=-\sum_{k=1}^{K}\sum_{i\in G_{k}}(\mathbf{X}_{i}-\bar{\mathbf{X}}^{k})^{\top}\boldsymbol{\Omega}^{\ast}(\mathbf{X}_{i}-\bar{\mathbf{X}}^{k}).$ (69)
The following is a well-known identity, which can be found in Equation 12.18 of James et al., (2013):
$\ell_{\mathrm{lik}}(G_{1},\ldots,G_{K};\mathbf{X})$ $=-\sum_{k=1}^{K}\sum_{i\in G_{k}}(\mathbf{X}_{i}-\bar{\mathbf{X}}^{k})^{\top}\boldsymbol{\Omega}^{\ast}(\mathbf{X}_{i}-\bar{\mathbf{X}}^{k})$
$=-\sum_{k=1}^{K}\sum_{i\in G_{k}}(\mathbf{X}_{i}-\frac{1}{|G_{k}|}\sum_{j\in G_{k}}\mathbf{X}_{j})^{\top}\boldsymbol{\Omega}^{\ast}(\mathbf{X}_{i}-\frac{1}{|G_{k}|}\sum_{j\in G_{k}}\mathbf{X}_{j})$
$=-\frac{1}{2}\sum_{k=1}^{K}\frac{1}{|G_{k}|}\sum_{i,j\in G_{k}}(\mathbf{X}_{i}-\mathbf{X}_{j})^{\top}\boldsymbol{\Omega}^{\ast}(\mathbf{X}_{i}-\mathbf{X}_{j}).$ (70)
If $\boldsymbol{\Sigma}^{\ast}=\mathbf{I}_{p}$, the maximization of (70) is formulated as the following K-means problem, by removing the constant factor $-1/2$:
$\min_{G_{1},\ldots,G_{K}}$ $~{}\sum_{k=1}^{K}\frac{1}{|G_{k}|}\sum_{i,j\in G_{k}}\|\mathbf{X}_{i}-\mathbf{X}_{j}\|_{2}^{2}$ (71)
$s.t.$ $~{}G_{1}\cup G_{2}\cup\ldots\cup G_{K}=\{1,\ldots,n\},~{}\text{and}$
$~{}G_{k}\cap G_{l}=\emptyset~{}\text{for all}~{}1\leq k\neq\leq K,$
where within-cluster dissimilarity is quantified by the sum of squared Euclidean distances.
2. From K-means objective to SDP objective.
The likelihood objective (70) can be further simplified as
$\ell_{\mathrm{lik}}(G_{1},\ldots,G_{K};\mathbf{X})$ $=-\frac{1}{2}\sum_{k=1}^{K}\frac{1}{|G_{k}|}\sum_{i,j\in G_{k}}(\mathbf{X}_{i}-\mathbf{X}_{j})^{\top}\boldsymbol{\Omega}^{\ast}(\mathbf{X}_{i}-\mathbf{X}_{j})$
$=\frac{1}{2}\sum_{k=1}^{K}\frac{1}{|G_{k}|}\sum_{i,j\in G_{k}}\bigl{(}2\mathbf{X}_{i}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}_{j}-\mathbf{X}_{i}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}_{i}-\mathbf{X}_{j}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}_{j}\bigr{)}$
$=\sum_{k=1}^{K}\biggl{(}\frac{1}{|G_{k}|}\sum_{i,j\in G_{k}}\mathbf{X}_{i}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}_{j}-\frac{1}{|G_{k}|}\sum_{i\in G_{k}}|G_{k}|(\mathbf{X}_{i}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}_{i})\biggr{)}$
$=\ell_{\operatorname{sdp}}(G_{1},\ldots,G_{k})-\sum_{i=1}^{n}\mathbf{X}_{i}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}_{i},$ (72)
where we recall that in (1) and (2), the SDP objective function is defined as follows:
$\ell_{\operatorname{sdp}}(G_{1},\ldots,G_{k};\mathbf{X}):=\sum_{k=1}^{K}\frac{1}{|G_{k}|}\sum_{i,j\in G_{k}}\mathbf{X}_{i}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}_{i}.$ (73)
From (72), we have the following relationship between the likelihood objective (69) and the SDP objective (73):
$\ell_{\operatorname{sdp}}(G_{1},\ldots,G_{k};\mathbf{X})$ $=\ell_{\operatorname{lik}}(G_{1},\ldots,G_{k};\mathbf{X})+\sum_{i=1}^{n}\mathbf{X}_{i}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}_{i}.$
The second term of the SDP objective does not depend on the cluster structure.
E.2 Intuition for the SDP objective in Algorithm 1
This section provides intuition on how the objective $\langle\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}^{\top}\boldsymbol{\Sigma}^{\ast}_{\hat{S}^{t},\hat{S}^{t}}\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot},\mathbf{Z}\rangle$ in Line 1 of Algorithm 1 arises from the original SDP objective $\langle\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X},\mathbf{Z}\rangle$ in (2), when restricted to a support set $S$, while capturing the key sparse clustering signal $\boldsymbol{\beta}^{\ast}=\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$.
We temporarily assume that $\boldsymbol{\Omega}^{\ast}$ is known, and denote the transformed data matrix as $\tilde{\mathbf{X}}:=\boldsymbol{\Omega}^{\ast}\mathbf{X}\in\mathbb{R}^{p\times n}$. Under this transformation, the similarity matrix in the original SDP objective (2) becomes
$\mathbf{A}:=\mathbf{X}^{\top}\boldsymbol{\Omega}^{\ast}\mathbf{X}=\tilde{\mathbf{X}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{X}}.$
We use the matrix-based notation introduced in Definition 12. Here, $\mathbf{M}^{\ast}$ denotes the deterministic mean matrix, where each column is either $\boldsymbol{\mu}_{1}^{\ast}$ or $\boldsymbol{\mu}_{2}^{\ast}$, and, the data matrix and its transformed version can be written as:
$\mathbf{X}=\mathbf{M}^{\ast}+\mathcal{E}~{}\text{and}~{}\tilde{\mathbf{X}}=\tilde{\mathbf{M}}^{\ast}+\tilde{\mathcal{E}},~{}\text{where}~{}\tilde{\mathbf{M}}^{\ast}:=\boldsymbol{\Omega}^{\ast}\tilde{\mathbf{M}}~{}\text{and}~{}\tilde{\mathcal{E}}:=\boldsymbol{\Omega}^{\ast}\mathcal{E},$
and the columns of $\mathcal{E}=(\mathcal{E}_{1},\ldots,\mathcal{E}_{n})$ are i.i.d. random vectors distributed as $\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}^{\ast})$. Using this notation, the similarity matrix can be decomposed into:
$\mathbf{A}=(\tilde{\mathbf{M}}^{\ast})^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{M}}^{\ast}+(\tilde{\mathbf{M}}^{\ast})^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathcal{E}}+\tilde{\mathcal{E}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{M}}^{\ast}+\tilde{\mathcal{E}}^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathcal{E}}.$
In the above decomposition, both the second and third terms are zero-mean, and the fourth term is independent of the cluster structure. Thus we focus on $(\tilde{\mathbf{M}}^{\ast})^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{M}}^{\ast}$.
Recall that $S_{0}\subset[p]$ denotes the index set where the corresponding entries of $\boldsymbol{\mu}^{\ast}_{1}$ and $\boldsymbol{\mu}^{\ast}_{2}$ differ (signal). Let $N:=[p]\setminus S_{0}$ denote the index set where the corresponding entries of $\boldsymbol{\mu}^{\ast}_{1}$ and $\boldsymbol{\mu}^{\ast}_{2}$ are identical (noise). Since $\boldsymbol{\beta}^{\ast}=\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$ is sparse, the row indices of $\tilde{\mathbf{M}}^{\ast}$ can also be decomposed into the signal set $S_{0}$ and the noise set $N$. Then $(\tilde{\mathbf{M}}^{\ast})^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{M}}^{\ast}$ is decomposed into
$(\tilde{\mathbf{M}}^{\ast})^{\top}\boldsymbol{\Sigma}^{\ast}\tilde{\mathbf{M}}^{\ast}$ $=(\tilde{\mathbf{M}}^{\ast}_{S_{0},\cdot})^{\top}\boldsymbol{\Sigma}_{S_{0},S_{0}}^{\ast}\tilde{\mathbf{M}}^{\ast}_{S_{0},\cdot}+(\tilde{\mathbf{M}}^{\ast}_{S_{0},\cdot})^{\top}\boldsymbol{\Sigma}_{S_{0},N}^{\ast}\tilde{\mathbf{M}}^{\ast}_{N,\cdot}$
$\quad\quad+(\tilde{\mathbf{M}}^{\ast}_{N,\cdot})^{\top}\boldsymbol{\Sigma}_{N,S_{0}}^{\ast}\tilde{\mathbf{M}}^{\ast}_{S_{0},\cdot}+(\tilde{\mathbf{M}}^{\ast}_{N,\cdot})^{\top}\boldsymbol{\Sigma}_{N,N}^{\ast}\tilde{\mathbf{M}}^{\ast}_{N,\cdot}.$
In the above decomposition, the first term, $(\mathbf{M}^{\ast}_{S_{0},\cdot})^{\top}\boldsymbol{\Sigma}_{S_{0},S_{0}}^{\ast}\tilde{\mathbf{M}}^{\ast}_{S_{0},\cdot}$ contains all the essential information for clustering, by the following reason. By expressing $\mathbf{Z}$ as $\mathbf{HBH}$ and tracing back the derivation in Appendix E.1, we have:
$\langle$ $(\mathbf{M}^{\ast}_{S_{0},\cdot})^{\top}\boldsymbol{\Sigma}_{S_{0},S_{0}}^{\ast}\tilde{\mathbf{M}}^{\ast}_{S_{0},\cdot}\mathbf{Z}\rangle$
$=\sum_{k=1}^{K}\frac{1}{|G_{k}|}\sum_{i,j\in G_{k}}{\mathbf{M}^{\ast}_{S_{0},i}}^{\top}\boldsymbol{\Sigma}_{S_{0},S_{0}}^{\ast}\mathbf{M}^{\ast}_{S_{0},i}$
$=-\frac{1}{2}\sum_{k=1}^{K}\frac{1}{|G_{k}|}\sum_{i,j\in G_{k}}(\mathbf{M}^{\ast}_{S_{0},i}-\mathbf{M}^{\ast}_{S_{0},j})^{\top}\boldsymbol{\Sigma}_{S_{0},S_{0}}^{\ast}(\mathbf{M}^{\ast}_{S_{0},i}-\mathbf{M}^{\ast}_{S_{0},j})+\sum_{i=1}^{n}\mathbf{M}^{\ast}_{S_{0},i}\boldsymbol{\Sigma}_{S_{0},S_{0}}^{\ast}\mathbf{M}^{\ast}_{S_{0},i}.$
The first term contains $\boldsymbol{\beta}^{\ast}$. Therefore, we propose running the SDP-relaxed K-means (2) using a noisy version of $(\mathbf{M}^{\ast}_{S_{0},\cdot})^{\top}\boldsymbol{\Sigma}_{S_{0},S_{0}}^{\ast}\tilde{\mathbf{M}}^{\ast}_{S_{0},\cdot}$ , given by the matrix $\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}^{\top}{\boldsymbol{\Sigma}}^{\ast}_{\hat{S}^{t},\hat{S}^{t}}\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}.$
Appendix F Details for Iterative Algorithms
This section supplements the iterative algorithms presented in Section 3 by providing omitted details. Appendix F.1 derives the uniform bound presented in (8), which is used in the variable selection step of Algorithm 2. Appendix F.2 provides additional details for the ISEE subroutine, which plays a central role in Algorithm 3.
F.1 Derivation of the uniform bound in the variable selection step of Algorithm 2
This section derives the uniform bound presented in equation (8) in Section 3.1. Since $\mathbf{X}_{i}\sim\mathcal{N}(0,\boldsymbol{\Sigma}^{\ast})$, it follows that $\tilde{\mathbf{X}}_{i}=\mathbf{X}_{i}\boldsymbol{\Omega}^{\ast}\sim\mathcal{N}(0,\boldsymbol{\Omega}^{\ast}\boldsymbol{\Sigma}^{\ast}\boldsymbol{\Omega}^{\ast}).$ Therefore, for any $j\notin S_{0}$, we have $\tilde{X}_{j,i}\sim\mathcal{N}(0,\omega_{jj})$, where $\omega_{jj}$ is the $j$th diagonal entry of $\boldsymbol{\Omega}^{\ast}$.
Given cluster assignments $G_{1}$ and $G_{2}$, the empirical transformed cluster means are defined as
$\bar{\tilde{\mathbf{X}}}_{G_{1}}:=\frac{1}{|G_{1}|}\sum_{i\in G_{1}}\tilde{\mathbf{X}}_{i},\quad\text{and}\quad\bar{\tilde{\mathbf{X}}}_{G_{2}}:=\frac{1}{|G_{2}|}\sum_{i\in G_{2}}\tilde{\mathbf{X}}_{i}.$
Regardless of the accuracy of the cluster assignments, for $j\notin S_{0}$, the $j$th entry of $\bar{\tilde{\mathbf{X}}}_{G_{1}}$ and $\bar{\tilde{\mathbf{X}}}_{G_{2}}$ follows the distributions
$\mathcal{N}\left(0,\frac{\omega_{jj}}{|G_{1}|}\right)\quad\text{and}\quad\mathcal{N}\left(0,\frac{\omega_{jj}}{|G_{2}|}\right),$
respectively. Let $\hat{\boldsymbol{\beta}}(G_{1},G_{2}):=\bar{\tilde{\mathbf{X}}}_{G}-\bar{\tilde{\mathbf{X}}}_{G_{2}}$. Then, for each $j\notin S_{0}$, the scaled $j$th element satisfies:
$\frac{1}{\sqrt{\omega_{jj}}}\hat{\beta}_{j}(G_{1},G_{2})\sim\mathcal{N}\left(0,\frac{n}{|G_{1}|\cdot|G_{2}|}\right),$
because $|G_{1}|+|G_{2}|=n$. Although the entries of $\hat{\boldsymbol{\beta}}(G_{1},G_{2})$ are not independent, the Gaussian maximal inequality still applies. Consequently, for any $\delta>0$, with probability at least $1-\delta$, we have:
$\max_{G_{1},G_{2}}\max_{j\in S\setminus S_{0}}\left|\sqrt{\frac{|G_{1}|\cdot|G_{2}|}{\omega_{jj}}}\hat{\beta}_{j}(G_{1},G_{2})\right|\leq\sqrt{2n\log\left(2(p-|S_{0}|)/\delta\right)}.$
This completes the derivation of the bound in equation (8).
F.2 Supplememtary details for ISEE subroutine
This section provides the omitted details for the ISEE subroutine. We proceed in the following steps:
  1. 1.
    Regression relationship on the data sub-vectors (Appendix F.2.1),
  2. 2.
    Key identities on transformed cluster center and noise (Appendix F.2.2),
  3. 3.
    Estimation error bound for the ISEE subroutine (Appendix F.2.3),
Without loss of generality, we focus on the case where $i\in G_{1}^{\ast}$. The derivations build on those in Fan and Lv, (2016), with the key distinction that they only consider the special case $\boldsymbol{\mu}_{1}^{\ast}=\boldsymbol{0}$. At the end of this section, Algorithm 4 formulates the ISEE subroutine presented in Section 3.2.1 as an algorithm.
F.2.1 Regression relationship on the data sub-vectors
Using the conditional distribution property of multivariate Gaussians and the Schur complement, we derive the following regression relationship between sub-vectors of $\mathbf{X}_{i}\sim\mathcal{N}(\boldsymbol{\mu}_{1}^{\ast},\boldsymbol{\Sigma}^{\ast})$ for $i\in G_{1}^{\ast}$, as presented in equation (10) in Section 3.2.1. Given a small index set $A\subset[p]$, we recall the following relationship:
[…]
where $\mathbf{E}_{A,i}\sim\mathcal{N}(0,(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1})$. Let us partition $\mathbf{X}_{i}$, $\boldsymbol{\mu}_{1}^{\ast}$, and $\boldsymbol{\Omega}^{\ast}$ according to :
$\mathbf{X}_{i}=\begin{bmatrix}\mathbf{X}_{A,i}\\ \mathbf{X}_{A^{c},i}\end{bmatrix},\quad\boldsymbol{\mu}_{1}^{\ast}=\begin{bmatrix}(\boldsymbol{\mu}_{1}^{\ast})_{A}\\ (\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}\end{bmatrix},\quad\boldsymbol{\Sigma}^{\ast}=\begin{bmatrix}\boldsymbol{\Sigma}^{\ast}_{A,A}&\boldsymbol{\Sigma}^{\ast}_{A,A^{c}}\\ \boldsymbol{\Sigma}^{\ast}_{A^{c},A}&\boldsymbol{\Sigma}^{\ast}_{A^{c},A^{c}}\end{bmatrix}.$
Then the conditional distribution $\mathbf{X}_{A,i}\mid\mathbf{X}_{A^{c},i}$ follows $\mathcal{N}(\bar{\boldsymbol{\mu}},\bar{\boldsymbol{\Sigma}})$, where
$\bar{\boldsymbol{\mu}}:=(\boldsymbol{\mu}_{1}^{\ast})_{A}+\boldsymbol{\Sigma}^{\ast}_{A,A^{c}}(\boldsymbol{\Sigma}^{\ast}_{A^{c}A^{c}})^{-1}\bigl{(}\tilde{\mathbf{X}}_{A^{c},i}-(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}\bigr{)},$ (74)
and
$\bar{\boldsymbol{\Sigma}}:=\boldsymbol{\Sigma}^{\ast}_{A,A}-\boldsymbol{\Sigma}^{\ast}_{A,A^{c}}\left(\boldsymbol{\Sigma}^{\ast}_{A^{c},A^{c}}\right)^{-1}\boldsymbol{\Sigma}^{\ast}_{A^{c},A}.$ (75)
Next, to apply the Schur complement, we first introduce the following simplfied notation for the matrix partition:
$\boldsymbol{\Sigma}^{\ast}=\begin{bmatrix}\boldsymbol{\Sigma}^{\ast}_{A,A}&\boldsymbol{\Sigma}^{\ast}_{A,A^{c}}\\ \boldsymbol{\Sigma}^{\ast}_{A^{c},A}&\boldsymbol{\Sigma}^{\ast}_{A^{c},A^{c}}\end{bmatrix}=\begin{bmatrix}\mathbf{A}&\mathbf{B}\\ \mathbf{C}&\mathbf{D}\end{bmatrix},$ (76)
Then by the Schur complement, we have:
$\boldsymbol{\Omega}^{\ast}=(\boldsymbol{\Sigma}^{\ast})^{-1}=\begin{bmatrix}\boldsymbol{\Omega}^{\ast}_{A,A}&\boldsymbol{\Omega}^{\ast}_{A,A^{c}}\\ \boldsymbol{\Omega}^{\ast}_{A^{c},A}&\boldsymbol{\Omega}^{\ast}_{A^{c},A^{c}}\end{bmatrix}=\begin{bmatrix}\mathbf{W}&-\mathbf{W}\mathbf{B}\mathbf{D}^{-1}\\ -\mathbf{D}^{-1}\mathbf{C}\mathbf{W}&\mathbf{D}^{-1}+\mathbf{D}^{-1}\mathbf{C}\mathbf{W}\mathbf{B}\mathbf{D}^{-1}\end{bmatrix},$ (77)
where
$\mathbf{W}:=(\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{C})^{-1}.$ (78)
Leveraging this representation of $\boldsymbol{\Omega}^{\ast}$, we can derive:
$-\boldsymbol{\Sigma}_{A,A^{c}}^{\ast}(\boldsymbol{\Sigma}_{A^{c},A^{c}}^{\ast})^{-1}\overset{(i)}{=}-\mathbf{B}\mathbf{D}^{-1}\overset{(ii)}{=}\mathbf{W}^{-1}(-\mathbf{W}\mathbf{B}\mathbf{D}^{-1})\overset{(iii)}{=}(\boldsymbol{\Omega}_{A,A}^{\ast})^{-1}\boldsymbol{\Omega}_{A,A^{c}}^{\ast},$
where step $(i)$ uses (76), step $(ii)$ uses (78), and step $(iii)$ uses (77). This lets us rewrite $\bar{\boldsymbol{\mu}}$ in (74) as follows:
$\bar{\boldsymbol{\mu}}$ $=(\boldsymbol{\mu}_{1}^{\ast})_{A}+\boldsymbol{\Sigma}^{\ast}_{A,A^{c}}(\boldsymbol{\Sigma}^{\ast}_{A^{c}A^{c}})^{-1}\bigl{(}\tilde{\mathbf{X}}_{A^{c},i}-(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}\bigr{)}$
$=(\boldsymbol{\mu}_{1}^{\ast})_{A}-(\boldsymbol{\Omega}_{AA}^{\ast})^{-1}\boldsymbol{\Omega}_{AA^{c}}^{\ast}\bigl{(}\tilde{\mathbf{X}}_{A^{c},i}-(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}\bigr{)}$
$=\underbrace{(\boldsymbol{\mu}_{1}^{\ast})_{A}+(\boldsymbol{\Omega}_{A,A}^{\ast})^{-1}\boldsymbol{\Omega}_{A,A^{c}}^{\ast}(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}}_{(\boldsymbol{\alpha}_{1})_{A}}-(\boldsymbol{\Omega}_{A,A}^{\ast})^{-1}\boldsymbol{\Omega}_{A,A^{c}}^{\ast}\tilde{\mathbf{X}}_{A^{c},i},$
Likewise, we can also simplify $\bar{\boldsymbol{\Sigma}}$ in (75) as follows:
$\bar{\boldsymbol{\Sigma}}=\boldsymbol{\Sigma}^{\ast}_{A,A}-\boldsymbol{\Sigma}^{\ast}_{A,A^{c}}\left(\boldsymbol{\Sigma}^{\ast}_{A^{c},A^{c}}\right)^{-1}\boldsymbol{\Sigma}^{\ast}_{A^{c},A}\overset{(i)}{=}\mathbf{A}-\mathbf{B}\mathbf{D}^{-1}\mathbf{C}\overset{(ii)}{=}\mathbf{W}^{-1}=(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1},$
where step $(i)$ uses (76), step $(ii)$ uses (78), and step $(iii)$ uses (77). Substituting these expressions into the conditional distribution $\mathbf{X}_{A,i}\mid\mathbf{X}_{A^{c},i}\sim\mathcal{N}(\bar{\boldsymbol{\mu}},\bar{\boldsymbol{\Sigma}})$, we obtain the regression relationship (10).
F.2.2 Key identities on transformed cluster center and noise
Next, we turn to the two key identities presented in (11):
$(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A}=\boldsymbol{\Omega}^{\ast}_{A,A}(\boldsymbol{\alpha}_{1})_{A},$
and
(~$(\tilde{\mathcal{E}}_{i})_{A}=\boldsymbol{\Omega}_{A,A}\mathbf{E}_{A,i},$
which play a central role in the ISEE subroutine.
We start with the first identity. By pre-multiplying $\boldsymbol{\Omega}_{A,A}^{\ast}$ to $(\boldsymbol{\alpha}_{1})_{A}$, we have:
$\boldsymbol{\Omega}_{A,A}^{\ast}(\boldsymbol{\alpha}_{1})_{A}=\boldsymbol{\Omega}_{A,A}^{\ast}(\boldsymbol{\mu}_{1}^{\ast})_{A}+\boldsymbol{\Omega}_{A,A^{c}}^{\ast}(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}.$
Therefore, we can express $(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A}$ as follows:
$(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A}=\boldsymbol{\Omega}^{\ast}_{A,\cdot}\;\boldsymbol{\mu}_{1}^{\ast}=\boldsymbol{\Omega}^{\ast}_{A,A}(\boldsymbol{\mu}_{1}^{\ast})_{A}+\boldsymbol{\Omega}^{\ast}_{A,A^{c}}(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}=\boldsymbol{\Omega}^{\ast}_{A,A}(\boldsymbol{\alpha}_{1})_{A}.$
This completes the derivation of the first identity.
Next, we derive the second identity:
~$\tilde{\mathcal{E}}_{A,i}$ $=\boldsymbol{\Omega}^{\ast}_{A,\cdot}\;\mathbf{X}_{i}-\boldsymbol{\Omega}^{\ast}_{A,\cdot}\;\boldsymbol{\mu}^{\ast}_{1}$
$=\boldsymbol{\Omega}^{\ast}_{A,A}\mathbf{X}_{A,i}+\boldsymbol{\Omega}^{\ast}_{A,A^{c}}\mathbf{X}_{A^{c},i}-\boldsymbol{\Omega}^{\ast}_{A,A}(\boldsymbol{\mu}_{1}^{\ast})_{A}+\boldsymbol{\Omega}^{\ast}_{A,A^{c}}(\boldsymbol{\mu}_{1}^{\ast})_{A^{c}}$
$=\boldsymbol{\Omega}^{\ast}_{A,A}\biggl{(}\mathbf{X}_{A,i}+(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1}\boldsymbol{\Omega}^{\ast}_{A,A^{c}}\mathbf{X}_{A^{c},i}-(\boldsymbol{\mu}^{\ast}_{1})_{A}+(\boldsymbol{\Omega}^{\ast}_{A,A})^{-1}\boldsymbol{\Omega}^{\ast}_{A,A^{c}}(\boldsymbol{\mu}^{\ast}_{1})_{A^{c}}\biggr{)}$
$=\boldsymbol{\Omega}^{\ast}_{A,A}\mathbf{E}_{A,i}.$
where the last equality uses the regression relationship (10). This completes the derivation of the second identity.
F.2.3 Estimation error bound for the ISEE subroutine
We derive the estimation error bound presented in (12) in Section 3.2.2, recalled below:
$\bigl{\|}\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-$ $(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\bigr{)}-\bigl{(}({\tilde{\boldsymbol{\mu}}}_{1}^{\ast})_{A_{\ell}}-({\tilde{\boldsymbol{\mu}}}_{2}^{\ast})_{A_{\ell}}\bigr{)}\bigr{\|}_{2}\lesssim C_{6}\max\{J\log p/n,\sqrt{\log p/n}\},$
where $J$ is the maximum number of nonzero off-diagonal entries per row of $\boldsymbol{\Omega}^{\ast}$, and $C_{6}=\|\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}\|_{2}$.
We start by using the triangle inequality to obtain the following bound:
$\bigl{\|}\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\bigr{)}-\bigl{(}({\tilde{\boldsymbol{\mu}}}_{1}^{\ast})_{A_{\ell}}-({\tilde{\boldsymbol{\mu}}}_{2}^{\ast})_{A_{\ell}}\bigr{)}\bigr{\|}_{2}~{}\leq~{}\|(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A_{\ell}}\|_{2}+\|(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}-(\tilde{\boldsymbol{\mu}}_{2}^{\ast})_{A_{\ell}}\|_{2}.$
We can bound the $\|(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A_{\ell}}\|_{2}$ term as follows:
$\|(\widehat{\tilde{\boldsymbol{\mu}}}_{1})$ ${}_{A_{\ell}}-(\tilde{\boldsymbol{\mu}}_{1}^{\ast})_{A_{\ell}}\|_{2}$
=$\;\overset{(i)}{=}\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\boldsymbol{\alpha}_{1})_{A_{\ell}}\|_{2}$
=$\;\overset{(ii)}{=}\bigl{\|}\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}~{}+~{}\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\boldsymbol{\alpha}_{1})_{A_{\ell}}\bigr{\|}_{2}$
$\;\;=\bigl{\|}\bigl{(}\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}_{A_{\ell},A_{\ell}}^{\ast}\bigr{)}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}~{}+~{}\boldsymbol{\Omega}_{A_{\ell},A_{\ell}}^{\ast}\bigl{(}(\boldsymbol{\alpha}_{1})_{A_{\ell}}-(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\bigr{)}\bigr{\|}_{2}$
$\;\;\leq\bigl{\|}\bigl{(}\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}_{A_{\ell},A_{\ell}}^{\ast}\bigr{)}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\bigr{\|}_{2}~{}+~{}\bigl{\|}\boldsymbol{\Omega}_{A_{\ell},A_{\ell}}^{\ast}\bigl{(}(\boldsymbol{\alpha}_{1})_{A_{\ell}}-(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\bigr{)}\bigr{\|}_{2}$
$\;\overset{(iii)}{\leq}\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}\;\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2}~{}+~{}\|\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}\;\|(\boldsymbol{\alpha}_{1})_{A_{\ell}}-(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2},$
where step $(i)$ uses the key identity from equation (11) and the fact that the ISEE subroutine employs a plug-in estimator, step $(ii)$ adds and subtracts the term $\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}$, and step $(iii)$ applies the operator norm bound $\|\mathbf{A}\mathbf{x}\|_{2}\leq\|\mathbf{A}\|_{2}\|\mathbf{x}\|_{2}$. We can also derive a similar bound for $\|(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}-(\tilde{\boldsymbol{\mu}}_{2}^{\ast})_{A_{\ell}}\|_{2}$. Therefore we have:
$\bigl{\|}\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-$ $(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\bigr{)}-\bigl{(}({\tilde{\boldsymbol{\mu}}}_{1}^{\ast})_{A_{\ell}}-({\tilde{\boldsymbol{\mu}}}_{2}^{\ast})_{A_{\ell}}\bigr{)}\bigr{\|}_{2}$
$\leq\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}\;\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2}~{}+~{}\|\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}\;\|(\boldsymbol{\alpha}_{1})_{A_{\ell}}-(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2}$
$\quad+\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}\;\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}\|_{2}~{}+~{}\|\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}\;\|(\boldsymbol{\alpha}_{2})_{A_{\ell}}-(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}\|_{2}.$
Under some regularity conditions, Fan and Lv, (2016) proves that $\|\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}-\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}$, $\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}-(\boldsymbol{\alpha}_{1})_{A_{\ell}}\|_{2}$, and $\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}-(\boldsymbol{\alpha}_{2})_{A_{\ell}}\|_{2}$ all have have convergence rates of order
$\max\{J\log p/n,\sqrt{\log p/n}\}.$
Therefore, we can conclude that
$\bigl{\|}\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}-(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\bigr{)}$ $-\bigl{(}({\tilde{\boldsymbol{\mu}}}_{1}^{\ast})_{A_{\ell}}-({\tilde{\boldsymbol{\mu}}}_{2}^{\ast})_{A_{\ell}}\bigr{)}\bigr{\|}_{2}$
$\lesssim\bigl{(}\|(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}}\|_{2}+\|(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}\|_{2}+2\|\boldsymbol{\Omega}^{\ast}_{A_{\ell},A_{\ell}}\|_{2}\bigr{)}\max\biggl{\{}\frac{J\log p}{n},\sqrt{\frac{\log p}{n}}\biggr{\}}$
$\lesssim C_{6}\max\biggl{\{}\frac{J\log p}{n},\sqrt{\frac{\log p}{n}}\biggr{\}}.$
This completes the estimation error bound presented in (12).
Algorithm 4 ISEE subroutine: ISEE
1:Data matrix $\mathbf{X}$, cluster estimate $\hat{G}_{1},\hat{G}_{2}$
2:$A_{1},\ldots,A_{L}\leftarrow$ non-overlapping subsets of $[p]$ of size 2 or 3
3:for  $\ell=1,2,\ldots L$ do
4:     $\mathcal{D}_{1}^{\ell}\leftarrow\{(\mathbf{X}_{A_{\ell}^{c},i},\mathbf{X}_{A_{\ell},i}):i\in\hat{G}_{1}\}$
5:     $(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}},\{\hat{\mathbf{E}}_{A_{\ell},i}\}_{i\in\hat{G}_{1}}\leftarrow$ slope and residuals obtained from running lasso on $\mathcal{D}_{1}^{\ell}$
6:     $\mathcal{D}_{2}^{\ell}\leftarrow\{(\mathbf{X}_{A_{\ell}^{c},i},\mathbf{X}_{A_{\ell},i}):i\in\hat{G}_{2}\}$
7:     $(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}},\{\hat{\mathbf{E}}_{A_{\ell},i}\}_{i\in\hat{G}_{2}}\leftarrow$ slope and residuals obtained from running lasso on $\mathcal{D}_{2}^{\ell}$
8:     $\hat{\mathbf{E}}_{A_{\ell},\cdot}\leftarrow(\hat{\mathbf{E}}_{A_{\ell},1},\ldots,\hat{\mathbf{E}}_{A_{\ell},n})$triangleright Residual block
9:     $\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}\leftarrow(\frac{1}{n}\hat{\mathbf{E}}_{A_{\ell},\cdot}\hat{\mathbf{E}}_{A_{\ell},\cdot}^{\top})^{-1},$triangleright Precision matrix block
10:     $(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}\leftarrow\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{1})_{A_{\ell}},$triangleright Transformed cluster 1 center block
11:     $(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}\leftarrow\hat{\boldsymbol{\Omega}}_{A_{\ell},A_{\ell}}(\hat{\boldsymbol{\alpha}}_{2})_{A_{\ell}}$triangleright Transformed cluster 2 center block
12:     $\hat{\tilde{\mathbf{M}}}_{A_{\ell},\cdot}\leftarrow$ an $|A_{\ell}|\times n$ matrix where the $i$th column is $(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{\ell}}$ if $i\in\hat{G}_{1}$, and $(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{\ell}}$ if $i\in\hat{G}_{2}$
13:     $\hat{\tilde{\mathbf{X}}}_{A_{\ell},\cdot}\leftarrow\hat{\tilde{\mathbf{M}}}_{A_{\ell},\cdot}+\boldsymbol{\Omega}_{A_{\ell},A_{\ell}}\hat{\mathbf{E}}_{A_{\ell},\cdot}$triangleright Transformed data block
14:$\hat{\tilde{\boldsymbol{\mu}}}_{1}\leftarrow\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{1}}^{\top},\ldots,(\hat{\tilde{\boldsymbol{\mu}}}_{1})_{A_{L}}^{\top}\bigr{)}^{\top},$triangleright Aggregate cluster 1 center blocks
15:$\hat{\tilde{\boldsymbol{\mu}}}_{2}\leftarrow\bigl{(}(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{1}}^{\top},\ldots,(\hat{\tilde{\boldsymbol{\mu}}}_{2})_{A_{L}}^{\top}\bigr{)}^{\top}$triangleright Aggregate cluster 2 center blocks
[…]
17:Output $\hat{\tilde{\boldsymbol{\mu}}}_{1},\hat{\tilde{\boldsymbol{\mu}}}_{2},\hat{\tilde{\mathbf{X}}}$
Appendix G Supplementary details for numerical study
This section provides additional details on the numerical studies, omitted from Section 4. Appendix G.1 provides details on parameter tuning, early stopping criterion, implementation details, and computing resources. Appendix G.2 describes sparsity-aware high-dimensional clustering methods used for comparison. Appendix G.3 presents two additional simulation results: 1) numerical evidence supporting the convergence of our algorithm under the proposed stopping criterion, and 2) evaluation of Algorithm 2 under known non-isotropic covariance.
G.1 Supplementary details for the simulations
This section provides details on hyperparameter tuning, software implementation, early stopping criteria, and computing resources related to the simulation study presented in Section 4.
Parameter tuning for lasso.
For regularization parameter tuning in the ISEE subroutine (Algorithm 4) at each iteration of Algorithm 3, we use the lasso path algorithm to efficiently explore a range of regularization values and select the one that minimizes the Akaike Information Criterion (AIC). This procedure is implemented using functions from the MATLAB package penalized (McIlhagga,, 2016).
Early stopping.
For efficiency, we employ a data-driven stopping criterion to terminate the algorithm before reaching the $T$th iteration. At each iteration $t$, the stopping condition is met if either of the following conditions is met: (1) the relative change in the objective value falls below $\pi\%$; or (2) after a warm-up period of $w$ iterations, the best objective value observed shows less than $\pi\%$ improvement over the most recent $\eta$ iterations. The algorithm terminates only when this condition holds for both the SDP objective in SDP K-means subroutine (line 1 of Algorithm 1) and the original K-means objective. Unless stated otherwise, we set $w=10$, $\eta=5$, and $\pi=1\%$ in our experiments. Appendix G.3 provides a numerical evaluation of this stopping criterion.
Software implementation.
Our algorithms are implemented in MATLAB, and all simulations are run using version R2023b. To solve the SDP problems, we use the ADMM-based solver SDPNAL++ (Sun et al.,, 2020), which scales efficiently for sample sizes up to $n<5000$. In the final step of the SDP-relaxed K-means, we extract the leading eigenvector (an $n$-dimensional vector) from the output matrix $Z$ and apply the kmeans function from MATLAB Statistics and Machine Learning Toolbox. For the ISEE subroutine, we utilize the MATLAB Parallel Computing Toolbox to simulataeneously estimate the cluster centers and precision matrix blocks. For graphical lasso in Algorithm 5, we use the implementation provided by Leung et al., (2016). In Scenario 1, hierarchical clustering is performed using the linkage function with default options, also from MATLAB Statistics and Machine Learning Toolbox. When comparing against baseline methods, we use their original implementations, some in MATLAB and others in R, executed within their respective environments. All implementations and simulation codes, including those for the baseline methods written by their original authors, are available at our GitHub repository.
Computing resources.
All simulations are conducted on a Linux machine equipped with an AMD EPYC 7542 processor (32 cores, 2.90 GHz) and 16 GB of RAM. A single run of Algorithm 3, from initialization to the final iteration, completes in under one minute on this system.
G.2 Sparsity-aware clustering methods used for benchmarking
This section provides details of the baseline sparsity-aware methods that were compared to our algorithms in Section 4.
IF-PCA.
Influential Feature PCA (IF-PCA; Jin and Wang, 2016) is a two step method. In the variable selection step, features with the highest Kolmogorov-Smirnov (KS) scores are selected, with the selection threshold determined in a data-driven manner using the Higher Criticism principle (Donoho and Jin,, 2004). In the clustering step, spectral clustering is applied: the top $K-1$ left singular vectors are extracted from the normalized data matrix after feature selection and used as inputs to the classical $K$-means algorithm to estimate cluster labels. For simulations, we use the authors’ MATLAB implementation.
SAS.
Sparse Alternate Similarity clustering (SAS; Arias-Castro and Pu, 2017) is an iterative method that uses the fact that the within-cluster dissimilarity of K-means, measured by the squared $\ell_{2}$ distances, can be decomposed coordinate-wise. The algorithm alternates between clustering and support estimation: given a support set, it applies K-means to estimate the clusters; given the cluster assignments, it selects the top $\hat{|S_{0}|}$ coordinates with the largest average within-cluster dissimilarity. The tuning parameter $\hat{|S_{0}|}$, which denotes the number of signal features, is selected using the following permutation-based approach: For each $\hat{|S_{0}|}=1,\ldots,p$, run the SAS algorithm to obtain corresponding cluster and support estimates. For each run, compare the within-cluster dissimilarity of the selected features under the estimated cluster assignment to that under a randomly permuted cluster assignment. The discrepancy between the two is quantified using the gap statistic (Tibshirani et al.,, 2001). Then choose the value of $\hat{|S_{0}|}$ that yields the largest gap statistic. We use the authors’ R implementation.
Sparse K-means.
Witten and Tibshirani, (2010) proposes Sparse K-means, an iterative method which exploits the fact that the within-cluster dissimilarity in K-means, measured by squared $\ell_{2}$ distances, can be decomposed coordinate-wise. The method assigns a feature weight $w_{j}$ to each variable $j=1,\ldots,p$, and constructs a feature-weighted version of the K-means objective. It then optimizes this objective over both feature weights and cluster assignments, subject to a combined lasso and ridge penalty. The proposed algorithm alternates between maximizing the objective with respect to the cluster assignments (given fixed feature weights) and with respect to the feature weights (given fixed cluster assignments). For simulations, we use the author’s R package sparcl.
CHIME.
CHIME (Cai et al.,, 2019) is an EM-type iterative method, where the E-step updates the cluster assignments based on current Gaussian mixture model parameters, and the M-step updates the model parameters with cluster labels fixed. Notably, CHIME avoids explicit covariance estimation by employing an $\ell_{1}$-regularized estimation of the sparse discriminative direction $\boldsymbol{\beta}^{\ast}=\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$. After a suitable number of iterations, the estimated $\boldsymbol{\beta}^{\ast}$ is plugged into the Fisher discriminant rule to produce the final cluster assignments. For simulations, we use the authors’ MATLAB implementation.
Algorithm 5 Iterative SDP K-means with unknown covariance using graphical lasso
1:Data matrix $\mathbf{X}$, maximum number of iteration $T$
2:Let $\hat{G}_{1}^{0},\hat{G}_{2}^{0}$ be cluster estimates obtained by running spectral clustering on $\mathbf{X}$
3:for  t = 1, 2, …$T$ do
4:     $\hat{\bar{\mathbf{X}}}_{1}^{t}\leftarrow|\hat{G}^{t-1}_{1}|^{-1}\sum_{i\in\hat{G}^{t-1}_{1}}\mathbf{X}_{i}$
5:     $\hat{\bar{\mathbf{X}}}_{2}^{t}\leftarrow|\hat{G}^{t-1}_{2}|^{-1}\sum_{i\in\hat{G}^{t-1}_{2}}\mathbf{X}_{i}$
6:     $\dot{\mathbf{X}}^{t}\leftarrow\bigl{(}\mathbf{X}_{\cdot,\hat{G}^{t-1}_{1}}-\hat{\bar{\mathbf{X}}}_{1}^{t}\mathbf{1}_{\hat{G}^{t-1}_{1}}^{\top},\;\mathbf{X}_{\cdot,\hat{G}^{t-1}_{2}}-\hat{\bar{\mathbf{X}}}_{2}^{t}\mathbf{1}_{\hat{G}^{t-1}_{2}}^{\top}\bigr{)}$triangleright Cluster-wise demeaned data matrix
7:     $\hat{\boldsymbol{\Omega}}^{t}\leftarrow\texttt{GraphicalLasso}\bigl{(}\dot{\mathbf{X}}^{t}\bigr{)}$
8:     $\hat{\omega}_{11}^{t},\ldots\hat{\omega}_{pp}^{t}\leftarrow\mathrm{diag}(\hat{\boldsymbol{\Omega}}^{t})$
9:     $\hat{\boldsymbol{\beta}}^{t}\leftarrow\hat{\boldsymbol{\Omega}}^{t}\bigl{(}\hat{\bar{\mathbf{X}}}_{1}^{t}-\hat{\bar{\mathbf{X}}}_{2}^{t}\bigr{)}$
10:     $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\hat{\omega}_{jj}^{t}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\right\}$triangleright Variable selection
11:     $\hat{\tilde{\mathbf{X}}}^{t}\leftarrow\hat{\boldsymbol{\Omega}}^{t}\mathbf{X}$
12:     $\hat{\boldsymbol{\Sigma}}^{t}_{\hat{S}^{t},\hat{S}^{t}}\leftarrow\bigl{(}(\hat{\boldsymbol{\Omega}}^{t})^{-1}\bigr{)}_{\hat{S}^{t},\hat{S}^{t}}$
13:     $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{SDPkmeans}(\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot},\hat{\boldsymbol{\Sigma}}^{t}_{\hat{S}^{t},\hat{S}^{t}})$triangleright Clustering (Algorithm 1)
14:Output $\hat{G}^{T}_{1},\hat{G}^{T}_{2}$
Next, to assess the efficacy of Algorithm 3, which avoids full precision matrix estimation and uses a cutoff based on the convergence rate of ISEE, we introduce two modified variants.
Algorithm 5: Algorithm 3 with graphical lasso.
At each iteration $t$, Algorithm 5 fully estimates the precision matrix using graphical lasso (Friedman et al.,, 2008), and uses it to compute the necessary quantities: $\tilde{\mathbf{X}}$, $\boldsymbol{\beta}^{\ast}$, and $\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}$. These estimates are then used in variable selection and clustering steps.
Algorithm 6: Algorithm 3 with threshold based on Gaussian maximal inequality.
The estimated submatrix blocks of the precision matrix, computed in line 8 of Algorithm 4, also provide precision matrix diagonal estimates, denoted $\hat{\omega}_{11},\ldots,\hat{\omega}_{pp}$. Algorithm 6 leverages these diagonal estimates to compute plug-in estimates for the quantities required by Algorithm 3, instead of using $\hat{\tilde{\boldsymbol{\mu}}}_{1}^{t}$ and $\hat{\tilde{\boldsymbol{\mu}}}_{2}^{t}$.
Algorithm 6 Iterative SDP K-means with unknown covariance using threshold based on Gaussian maximal inequality
1:Data matrix $\mathbf{X}$, maximum number of iteration $T$
2:Let $\hat{G}_{1}^{0},\hat{G}_{2}^{0}$ be cluster estimates obtained by running spectral clustering on $\mathbf{X}$
3:for  t = 1, 2, …$T$ do
4:     $\hat{\tilde{\mathbf{X}}}^{t},\omega_{11}^{t},\ldots,\omega_{pp}^{t}\leftarrow\texttt{ISEE}(\mathbf{X},\hat{G}_{1}^{t-1},\hat{G}_{2}^{t-1})$$\triangleright$ Algorithm 4 modified to output $\hat{\omega}_{11}^{t},\ldots,\hat{\omega}_{pp}^{t}$
5:     $\hat{\boldsymbol{\beta}}^{t}\leftarrow|\hat{G}^{t-1}_{1}|^{-1}\sum_{i\in\hat{G}^{t-1}_{1}}\hat{\tilde{\mathbf{X}}}^{t}_{i}-|\hat{G}^{t-1}_{2}|^{-1}\sum_{i\in\hat{G}^{t-1}_{2}}\hat{\tilde{\mathbf{X}}}^{t}_{i}.$
6:     $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\hat{\omega}_{jj}^{t}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\right\}$triangleright Variable selection
7:     $\hat{\boldsymbol{\Sigma}}^{t}_{\hat{S}^{t},\hat{S}^{t}}\leftarrow$ pooled covariance of covariances computed from $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{1}}$ and $\mathbf{X}_{\hat{S}^{t},\hat{G}^{t-1}_{2}}$
8:     $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{SDPkmeans}(\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot},\;\hat{\boldsymbol{\Sigma}}^{t}_{\hat{S}^{t},\hat{S}^{t}})$triangleright Clustering (Algorithm 1)
9:Output $\hat{G}_{1}^{T},\hat{G}_{2}^{T}$
G.3 Additional simulation results
This section presents two additional simulation results.
G.3.1 Numerical Evidence Supporting the Convergence of Our Algorithm
We provide numerical evidence supporting the convergence of Algorithm 3 under the stopping criterion presented in Appendix G.1. We re-use the setup from Scenario 2 of Section 4, speicifically the scenario $(p,\Delta_{\boldsymbol{\Sigma}^{\ast}},\rho)=(400,4,0.45)$. Line 1 of Table 3 presents the details of the setting. The stopping criteria are applied using the hyperparameters $T=100$, $\pi=0.5\%$, $w=40$, and $\eta=20$. We conduct 200 independent replications to record when the iterative algorithm 3 terminates. At each stopping point, we record the mis-clustering rate, the number of true positives (signal entries retained), and false positives (noise entries retained). The results in Figure 3 reveal two main insights about the behavior of our stopping criteria and the performance of Algorithm 3: First, the stopping criteria are effective. Figure 3-(a) shows that the algorithms consistently stop before reaching the maximum iteration limit, providing empirical evidence of convergence. Across different estimation-cutoff strategies, the stopping points are similarly concentrated. Second, true discoveries contribute more to clustering performance than the avoidance of false ones. As shown in Figure 3-(b), the original Algorithm 3 demonstrates more stable clustering performance compared to its variant, which often yields mis-clustering rates near 0.5. Figures 3-(c) and 3-(d) shows that the original algorithm makes slightly more true discoveries and incurs many more false discoveries, respectively.
Figure 3: Histograms from 200 independent runs comparing Algorithm 3 and its variants shown in Figure 2. The data setting follows line 1 of Table 3, and the stopping criteria are applied using the hyperparameters $T=100$, $\pi=0.5\%$, $w=40$, and $\eta=20$.
Table 3: Summary of Simulation Settings. $i,j=1,\ldots,p$.
Methods Evaluated Covariance structure $p$ $n$ Separation
Early stopping of Algorithm 3 unknown, $\Omega^{*}_{i,i+1}=\Omega^{*}_{i+1,i}=\rho$, $\Omega^{*}_{i,i}=1$, $\Omega^{*}_{i,j}=0$ o.w., $\rho=0.45$ 400 500 $\Delta_{\boldsymbol{\Sigma}^{\ast}}^{2}=4^{2}$
Algorithm 2 (for known $\boldsymbol{\Sigma}^{\ast}$) known, $\Sigma^{\ast}_{i,j}=\rho^{|i-j|}$, $\rho\in\{0.2,0.5,0.8\}$ 50-5000 200 $\Delta_{\boldsymbol{\Sigma}^{\ast}}^{2}=3^{2},4^{2}$
G.3.2 Evaluation of Algorithm 2 under known non-isotropic covariance
Under the known AR(1) covariance setting, we numerically validate that Algorithm 2 effectively adapts to sparsity of $\boldsymbol{\beta}^{\ast}$ and examine how varying correlation levels impact clustering. The data generation setting is the same as in Section 4, with details specified in line 2 of Table 3. The result presented in Figure 4 highlights two key findings. First, the iterative algorithm successfully adapts to sparsity of discriminative direction: its performance remains stable (especially in $\Delta_{\boldsymbol{\Sigma}^{\ast}}=4$) across increasing ambient dimensions $p$. Second, when the covariance structure is known, higher correlation improves clustering. As $\rho$ increases, mis-clustering rates degrade more slowly with $p$ due to the redundancy in noise.
Figure 4: Mis-clustering rates of Algorithm 2 with spectral initialization under the AR(1) correlation structures for $\rho=0.2,0.5,0.8$. The data generation setting is the same as in Section 4, with details specified in line 2 of Table 3. The dotted line of the same color represents the mis-clustering rate of spectral initialization under the same covariance structure.